### v1

# Deliberative XAI: How Explanations Impact Understanding and Decision-Making of AI Novices in Collective and Individual Settings

Abstract.
XAI research often focuses on settings where people learn about and assess algorithmic systems individually. However, as more public AI systems are deployed, it becomes essential for XAI to facilitate collective understanding and deliberation. We conducted a task-based interview study involving 8 focus groups and 12 individual interviews to explore how explanations can support AI novices in understanding and forming opinions about AI systems. Participants received a collection of explanations organized into four information categories to solve tasks and decide about a system’s deployment. These explanations improved or calibrated participants’ self-reported understanding and decision confidence and facilitated group discussions. Participants valued both technical and contextual information and the self-directed and modular explanation structure. Our contributions include an explanation approach that facilitates both individual and collaborative interaction and explanation design recommendations, including active and controllable exploration, different levels of information detail and breadth, and adaptations to the needs of decision subjects.
1. Introduction
A growing number of algorithmic decision-making systems (ADM systems) are deployed and used in the public sector to decide about critical issues, such as employment, migration, and criminal justice (Scott et al., 2022; Bansak et al., 2018; Züger and Asghari, 2023; Chouldechova, 2017). These systems can result in discrimination and damages for decision subjects (Brown et al., 2019; Raji et al., 2022), especially when decision-making is opaque and uncontestable (Ananny and Crawford, 2018; Alfrink et al., 2023; de Fine Licht and de Fine Licht, 2020). The field of explainable AI (XAI) aims to reduce these risks by developing methods to make ADM systems more understandable and facilitate their assessment and control (Langer et al., 2021). But while much of XAI research is focused on individuals, research has shown that learning about and understanding complex topics benefits from collaboration in group settings (Smith et al., 2009; Moshman and Geil, 1998; Nokes-Malach et al., 2015; Navajas et al., 2018). Work on public AI further argues that multi-stakeholder deliberation is an essential step in ensuring public participation in deployment decisions of these systems (Kawakami et al., 2024; Naiseh et al., 2024; Brown et al., 2019), but how explanations can fulfill these demands is an open question in XAI.
Crucially, many people affected by decisions of ADM systems have no technical background in developing or analyzing these systems. That is, a majority of decision subjects, domain experts, and other stakeholders of the public are lay people (Lima et al., 2023; DeVito et al., 2018; Shulner-Tal et al., 2022) or AI novices111We use the term in reference to ”visualization novices” (Burns et al., 2023) and synonymously with ”lay people” (Szymanski et al., 2021; Shen et al., 2020; DeVito et al., 2018).. Explanations for AI novices naturally need to fulfill different requirements than for AI practitioners, as they have to be adapted to their background (Ehsan et al., 2024), interests (Langer et al., 2021), and prior knowledge (Szymanski et al., 2021; Cheng et al., 2019; Schmude et al., 2024), which is why personalization (Shulner-Tal et al., 2022; Conati et al., 2021; Naiseh et al., 2020) and interactivity (Cheng et al., 2019; Bertrand et al., 2023; Guesmi et al., 2023) are central concepts in explanation design. Further, while explanation formats (e.g., interactive, visual, textual) are known to impact AI novices’ understanding of algorithmic systems (Cheng et al., 2019; Bove et al., 2022; Szymanski et al., 2021), the precise effect is inconsistent between studies (Bove et al., 2022) due to contextual factors such as participants’ perceptions of the use case domain. Previous qualitative research identified the information needs of AI novices (Schmude et al., 2024), emphasized user-preference for social and conversational formats (Bertrand et al., 2023), and employed participatory methods to design explanations in collaboration with end-users (Weitz et al., 2024; Lee et al., 2019). This work connects to these insights and draws from disciplines such as CSCW (Convertino et al., 2008; Fiesler et al., 2019), cognitive science (Keil, 2006, 2003), and educational psychology (Smith et al., 2009; Nokes-Malach et al., 2015) to examine how explanations can support understanding and decision-making of AI novices in individual and collaborative settings (Naiseh et al., 2024; Weitz et al., 2024).
Developing this design approach to improve understanding and facilitate deliberation meets multiple challenges. Group composition and dynamics place special demands on explanation design (Naiseh et al., 2021), as explanations must cover a diverse set of information needs (Schmude et al., 2024; Bove et al., 2024) and format preferences (Bertrand et al., 2023). In addition, they must support collaborative interaction (Long et al., 2021), such as sharing and simultaneous processing of information. That is, explanations need to provide comprehensive information but remain clear and navigable, and should foster group interaction but still be suitable for individual settings. We address these challenges by proposing a modular explanation design that spans four information categories from which users can select information according to their needs. Additionally, when validating explanation approaches qualitatively, it is crucial to contact the relevant stakeholder groups, which can be especially challenging with decision subjects in minority groups or difficult situations, who arguably have the highest risk of incurring negative consequences from ADM. To address this, we conducted two focus groups with decision subjects to include their perspectives and voices in developing explanations for AI systems in the public sector (Brown et al., 2019).
This paper explores how explanations can be used in individual and collaborative settings to support AI novices’ learning, interaction, and decision-making regarding the deployment of ADM systems. We present the findings of a task-based interview study with 43 AI novices involving 8 focus groups and 12 single interviews. We propose an explanation design that addresses AI novices’ information needs and applies current best practices in XAI (Byrne, 2023). The design consists of a collection of 36 explanations in the form of question-answer pairs organized into the four categories data, system details, usage, and context, which are further subdivided into subtopics and levels of detail (Figure 1). Participants used these explanations to solve the study tasks and decide about deploying an ADM system (Figure 3). The study’s use case is an algorithm for the prediction of job-seekers’ employability, connecting to previous work on ADM systems in employment (Scott et al., 2022; Niklas et al., 2015; Lopez, 2019). We analyze participants’ self-reported understanding, decision confidence, and perceptions of key information, and conduct a thematic analysis of group interactions with explanations. We discuss explanation design suggestions and the benefits and challenges of using explanations in collaborative settings. In the analysis of our findings, we put particular focus on the following research questions:
[RQ1]
Explanations: How does a question-driven, modular explanation design support AI novices’ understanding in individual and group settings?
[RQ2]
Deliberation: How do the explanations support group interaction with information and deliberation processes?
The contributions of our work include i) the validation of an explanation design building on current XAI principles, including question-driven and modular design, different levels of completeness and soundness, and suitability for both individual and collective settings; ii) an analysis of the kind and amount of information that participants requested most often and perceived as most important; iii) an analysis of how our explanations support group collaboration in solving tasks and group deliberation in forming opinions about ADM systems; and iv) recommendations for the design and use of these explanations and implications for the incorporation of group settings in XAI.
2. Background and Related Work
In this section, we briefly describe how our work connects to human-centered AI, research on explanation design for AI novices, and literature on understanding, interaction, and decision-making in groups.
2.1. Human-centered explainable AI
Explainability is often described as a cornerstone of responsible AI systems222While not every ADM system includes AI, we use both terms as we focus on decision-making systems with machine learning components. (Thiebes et al., 2021), as explanations can enable people to understand (Langer et al., 2021) and contest AI decisions (Alfrink et al., 2023). Similar concepts are embedded in the notion of human-centered AI (Capel and Brereton, 2023), which proposes to 1) build systems based on user observation and stakeholder engagement and 2) empower rather than replace people by making these systems controllable and autonomy-preserving (Shneiderman, 2022; Shin, 2023; Xu, 2019). These measures become especially important in high-risk settings (European Commission, 2021), such as employment (Scott et al., 2022; Flügge, 2021), immigration (Bansak et al., 2018), or criminal justice (Chouldechova, 2017), where erroneous and non-transparent algorithmic decisions can cause severe harm to those affected (Raji et al., 2022). To better understand how explanations can contribute to these principles, human-centered explainable AI places emphasis on the sociotechnical dimensions of explaining AI systems (Ehsan et al., 2023b) in that it assumes that transparency alone is not enough to make AI systems explainable (Ananny and Crawford, 2018), but that explanations need to consider the system’s social context (Wenzelburger et al., 2022), its lifecycle (Dhanorkar et al., 2021) and its different stakeholder groups (Ehsan et al., 2023b). In the context of this work, human-centered explainability is realized by testing and validating a design approach that has been developed based on qualitative exploration of AI novices’ information needs (Schmude et al., 2024). Further, our work aims to examine suitable formats for explanations that can accompany the deployment of AI systems in public institutions (de Fine Licht and de Fine Licht, 2020) and inform approaches of practicing responsible AI in high-risk application settings (Ehsan et al., 2023b).
2.2. Explanation design for AI novices
Explanations for non-technical audiences have different requirements than those for AI practitioners. Explanation methods like LIME (Ribeiro et al., 2016), SHAP (Lundberg and Lee, 2017), and surrogate models (Molnar, 2022) are better suited to experts due to their reliance on previous knowledge and technical literacy, while effective methods for AI novices are still being developed. As transparency does not equal understanding (Ananny and Crawford, 2018), simply making all information about a system available is no valid explanation strategy. Instead, explanations must be selected and designed to be suitable to the recipient’s stakeholder role (Langer et al., 2021), prior knowledge (Schmude et al., 2024), beliefs (Miller, 2019), and explanatory stance (Byrne, 2023; Keil, 2006). Various studies examined how such explanations for AI novices might be realized: Cheng et al. (2019) showed that ”white-box” explanations (transparent models) can improve ”objective”333We use the term in reference to Cheng et al. (2019) and Bove et al. (2022), it means to describe factual or testable understanding. understanding but may overwhelm users, reducing perceived understanding. However, Bove et al. (2022) used similar explanations and found that they had the opposite effect on participant understanding, which they attribute to a difference in the studies’ use case domains (student admission vs. finance). Szymanski et al. (2021) examined how expert and lay users rated explanations of different formats and found that while lay users favored visual explanations, they performed better with textual ones. Other studies underscore this variability, indicating that comprehension varies with personal and demographic factors and domain knowledge (Wang and Yin, 2021; Shulner-Tal et al., 2022; Ehsan et al., 2024) and that personalized explanations could be effective in addressing this (Conati et al., 2021; Shulner-Tal et al., 2022). Further aspects to be considered include the explanation’s purpose (Freiesleben and König, 2023) and the recipient’s familiarity with AI (Kramer et al., 2018). Regarding the information needs, previous qualitative research outlined that AI novices value information about the context and intention of a system’s deployment (Schmude et al., 2024; Kawakami et al., 2024) as well about the responsible institution (Brown et al., 2019). Notably, Lee et al. (2019) and Weitz et al. (2024) conducted participatory workshops to design explanations with end-users in the public sector, finding that participants valued the collaborative settings and their improved understanding of algorithmic systems. Drawing from these approaches, this work focuses on the design and qualitative validation of explanations such that they address the information needs of AI novices (Schmude et al., 2024) and support their understanding (Langer et al., 2021) and decision-making (Bertrand et al., 2023) in individual and collective settings.
Our explanation design draws from various approaches in previous work. The information base used for the explanations was compiled from different sources documenting the development and planned deployment of employment prediction algorithms (Scott et al., 2022; Allhutter et al., 2020), producing an extensive collection of ”scavenged” (Wieringa, 2023) material. To structure this collection, we drew from Lim and Dey (2009)’s work on intelligibility types, Liao et al. (2020)’s concept of question-driven explanation design, and Schmude et al. (2024)’s separation of information categories into data, system details, usage, and context. We further apply Kulesza et al. (2013)’s principles on explanation soundness (fidelity, complexity) and completeness (coverage, density) and draw from works that apply these principles in explanation design (Chatti et al., 2022; Guesmi et al., 2023; Kulesza et al., 2015), by introducing a structure of sub-topics and a hierarchy of explanation levels. This combination of question-driven explanations, different levels of information detail and breadth, and user-controlled selection of information aims to introduce modularity and interactivity (Cheng et al., 2019; Schmude et al., 2023; Guesmi et al., 2023) while allowing for adjusting of explanations to users’ needs (Conati et al., 2021; Shulner-Tal et al., 2022).
2.3. Understanding and decision-making in collaborative settings
Improving understanding is a central goal of explanations (Langer et al., 2021) and enables people to pursue their goals, such as assessing or contesting a system’s decisions (Hoffman et al., 2023). However, understanding can be defined in numerous ways (Grimm, 2019; Baumberger et al., 2017; Zagzebski, 2019; Keil, 2006). This work draws from literature in the learning sciences, cognitive sciences, and XAI to define understanding as i) connecting and applying information, whereas ”knowledge” only stores information (Grimm, 2019; Baumberger et al., 2017), ii) being the attempt to grasp the underlying structure of a phenomenon by way of simplification (Zagzebski, 2019), iii) consisting of several ”facets” that include both the analytical and the emotional connection to information (Wiggins and McTighe, 2005), and iv) rarely being exhaustive, but rather a ”working” mental model that is attained by recognizing and filling gaps until the learner deems it sufficient (Keil, 2006). Due to the challenge of defining and measuring understanding (Sato et al., 2019), recent research has proposed an ”abilities-based” approach (Speith et al., 2024) which defines understanding through the skills the learner possesses, connecting to comparable operationalizations by the learning sciences (Wiggins and McTighe, 2005).
While individual understanding has been the subject of many studies in XAI (Chen et al., 2023; Cheng et al., 2019; Wang and Yin, 2021; Schmude et al., 2023), understanding in group settings has been less explored. However, different disciplines have investigated collaborative understanding in detail: In the cognitive sciences, distributed cognition (Keil, 2003) and outsourced understanding (Keil, 2006) are components for the analysis of explanatory understanding, whereas educational psychology produced insight on the effects of group tutoring (Nokes-Malach et al., 2015; Smith et al., 2009; Moshman and Geil, 1998; Navajas et al., 2018) and comparisons to one-on-one tutoring (Bloom, 1984). Further, research in the field of computer-supported collaborative work provides approaches for the analysis of collaborative settings (Fiesler et al., 2019), group composition (Convertino et al., 2008), and group interactions (Sutcliffe, 2005). Recent research on public AI (Züger and Asghari, 2023) has emphasized the importance of including stakeholders by introducing deliberation early in the design process (Kawakami et al., 2024; Yurrita et al., 2022; Zhang et al., 2023). Further, recent work in XAI has begun to consider how explanations for group interactions could be approached, describing that ”many-to-one” interactions (multiple people interacting with an explanation) will likely differ enormously from ”one-to-one” interactions due to ”complexities in group dynamics, cognitive bias amplification, trust issues within the group, and group-centric evaluation” (Naiseh et al., 2024). Drawing from these conceptualizations, this work aims to explore empirically how explanations can support group deliberation by improving understanding and informing decision-making processes.
3. Methods
In this section, we describe our methods and study procedure. We conducted a task-based semi-structured interview study with 43 participants recruited from civil society organizations, a job agency, and the author’s extended network (Section 3.5), structured into 8 focus groups (3–5 participants) and 12 single interviews. Participants were presented with the study’s employment prediction use case (Section 3.1) and received a collection of explanations about this system (Section 3.2) before solving four tasks and making decisions about the system’s deployment (Section 3.3). The study closed with an interview, lasting 90–120 minutes for focus groups and 60 minutes for single interviews. We analyzed the explanations’ effect, participants’ prioritization of information, and the individual and collective interaction with the explanations (Section 3.4). The university’s research ethics committee approved this study.
3.1. Use case: The employment prediction algorithm
The employment prediction algorithm444More detailed information on the use case has been omitted to adhere to the anonymization policy but will be re-inserted for the final version. is a system meant to support counselors of a public employment agency in assessing job-seekers’ employability. It uses a logistic regression model trained on historical data to assess how demographic features (e.g., age, education, nationality) influence employment chances, producing short-term and long-term employment scores as recommendations. Counselors use these recommendations to decide about adequate support measures for the job-seeker (courses, stabilizing measures, etc.) but can override the system’s recommendations with justification. The system is comparable to algorithmic profiling tools that are used in various countries to assist in job-seeker assessment and resource allocation (Scott et al., 2022; Lopez, 2019; Niklas et al., 2015) and which repeatedly highlight the limits of automation in interpersonal interaction (Bundesagentur für Arbeit, 2021; Niklas et al., 2015; Desiere and Struyven, 2021; Scott et al., 2022). Due to its documentation and exemplification of typical ADM systems’ challenges in public employment, the employment prediction algorithm was thus chosen as the use case for this study.
3.2. Explanation design
Our explanation design was composed of 36 question-answer pairs about the employment prediction algorithm. Each question belonged to one of four categories, data (format, content, limitations), system details (features, model, examples), usage (operation by and interaction with users), and context (intention of deployment, target group, responsible actors), which were informed by work about the information needs of AI novices (Schmude et al., 2024). Each category was further divided into subtopics with three levels of increasing detail (base level, level 2, level 3). Every explanation contained a question (e.g., ”Who operates the system?”) answered with a brief text or image printed on an A5 paper sheet (Figure 2). Participants first received an explanation overview (Figure 1) and the four base explanations and could request levels 2 and 3 at any time during the explanation phase (as depicted in Figure 3).
The structure of the explanations was intended to i) facilitate uptake by providing information in a modular format and ii) provide different levels of both soundness (information levels) and completeness (subtopics) of information (Kulesza et al., 2013, 2015; Chatti et al., 2022). The selection of subtopics and information levels was guided by work on ADM systems in public employment (Scott et al., 2022; Niklas et al., 2015; Allhutter et al., 2020) and research on the prioritization of information by AI novices (Schmude et al., 2024; Weitz et al., 2024). The self-directed provision of information allowed participants to select explanations according to their interests (Conati et al., 2021) and information needs. The question-answer style was motivated by work on question-driven explanations (Liao et al., 2020) and was intended to encourage engagement. The explanations made use of different techniques that include feature importance, local and global explanations, examples, counterfactuals, and argumentative approaches (Figure 2). Lastly, the explanations were presented in an analog paper format to encourage interactions such as handing around paper sheets, pointing, and reading aloud.
3.3. Study setup and procedure
The study procedure consisted of several steps that are depicted in Figure 3. This section describes the introduction of the use case, the explanation phase comprising exploration, tasks, and the group decision, and the mechanisms for participants’ self-reports on understanding, confidence, and other aspects.
Use case introduction. Participants received initial information about the employment prediction algorithm in the form of a mock newspaper article. The article provided key information and featured the perspectives of employers and employee associations.
Exploration phase. After receiving the explanation overview and base level explanations of each category, participants had 15 minutes (focus groups) or 20 minutes (interviews) to become familiar with the structure and mark relevant explanations. Participants could freely decide which explanations to read and request and were encouraged to discuss them.
Tasks. Participants then received a fictional job-seeker case example describing Mr. Harald G.555Details on the inspiration for this case example where omitted to adhere to anonymization policy, but will be re-inserted for the final version. , a fictional job-seeker with a brief backstory and a list of features that would be used to calculate his employment chances. Participants solved four tasks formulated as questions covering i) the job-seeker’s ability to change information stored about him (such as education, physical impairment, etc.), ii) his employability scoring by the employment prediction algorithm, iii) the support measures he would receive, and iv) his ability to contest the decision. All questions had multiple answers with one correct choice and could be answered with information present in the explanations in different categories and detail levels. Whereas tasks 1, 3, and 4 required mostly information retrieval, task 2 could be solved in two ways: by either giving an estimate based on the rough weightings in the system details base explanations or by calculating the precise employment score. Participants had 15 (focus groups) or 20 (single interviews) minutes to solve the tasks. During that time, they could access and request all explanations and discuss possible solutions.
Decision task. Lastly, participants in focus groups had 10 minutes to discuss the system’s deployment and make a collective yes / no decision. If no consensual decision was reached in time, participants either decided themselves how to resolve the situation (e.g., majority vote) or received the option of stating conditions for the system’s deployment.
Self-reports. Participants reported their understanding, decision confidence, and inclusion and voice in the group with 5-point Likert scales, their decision on the system’s deployment with yes / no voting ballots, and their choice of the explanation categories most ”helpful for their understanding” and ”influential to their decisions” by selecting the categories that applied. Self-reports were conducted using a color-coded system to ensure anonymity during focus groups. Each participant was assigned a color and given matching reporting material. During the second self-report, the investigator also asked interview questions about participants’ interaction with the explanations, their understanding processes, the most relevant information, and any additional situational questions. In single interviews, the collective decision and questions about inclusion and voice in the group were omitted.
3.4. Analysis
All focus groups and interviews were audio-recorded and transcribed. These transcripts, participants’ self-reports during the studies, and the investigators’ field notes, provided the data basis for analysis.
For RQ1-Explanations, we focus on how the explanations helped and impacted participants’ understanding of the ADM system and examine the differences between individual and group settings. For this, we compare participants’ self-reported understanding before and after the explanation phases, their performance in the study tasks, and which and how many explanations were requested. Further, we conduct an inductive thematic analysis (Braun and Clarke, 2006) of participants’ responses to the investigator’s interview questions, capturing articulations about the strengths and weaknesses of the explanations’ design, the impact of the time limit, understanding barriers, and information perceived as most relevant. While the quantitative items thus serve to characterize the diversity of participants’ perceptions (Maxwell, 2010) and facilitate qualitative exploration (Weiss, 1995), they are not intended to invoke ”inference […] of greater generality” (Maxwell, 2010) nor impose a mental model based on variance theory (Patton, 1990).
For RQ2-Deliberation, we first focus on how participants shared and combined information to solve the study tasks and, second, on how explanations enabled group deliberation. For the first part, we conduct an inductive and deductive thematic analysis of the groups’ interaction processes with explanations, connecting our descriptions of these interaction processes to mechanisms for collaborative success and failures as described by Nokes-Malach et al. (2015). For the second part, we compare self-reported decision confidence before and after participants received and discussed the explanations, summarize their feedback about the group setting, and conduct an inductive thematic analysis of focus group conversations in which explanations drove group deliberation and convergence, facilitating the opinion-forming of participants.
3.5. Participants
Recruitment. Table 1+2 provides an overview of the study participants. Participants were recruited through cooperation with civil society organizations, an employment agency, and the authors’ extended network. Recruitment criteria included full legal age and no technical knowledge about AI systems, i.e., AI novices. All studies were conducted in person in office or public spaces. Participants were compensated with 30€ for participation in focus groups (90–120 minutes) and 20€ for single interviews (60 minutes). Our approach for organizing, composing, and moderating focus groups was informed by Krueger (2004). Concerning the participant sample size, we are guided by research on qualitative methods, which suggests that the number of participants should be determined by code and meaning saturation (Hennink et al., 2017).
Choice of participants. Participants were selected to be representatives of one of three roles: domain experts, decision subjects, or members of the general public. We define domain experts as people who are competent in the field that the ADM system is used in, such as job counselors or advisors (groups 2, 3, 7, and 8). We define decision subjects as people who would potentially be impacted by the system’s decision, such as job-seekers and people who had previously been job-seeking (groups 5 and 6). All remaining participants are considered as members of the general public and were included to test changes in explanation effects and participants’ perceptions (groups 1 and 4). The study was conducted with separate participants in three pilot groups to test and refine the explanation design and study procedure.
| Group | ID | Age | Education | Occupation | Group | ID | Age | Education | Occupation |
| Group A | A1 | 63 | University | Retired | Group F | F1 | 48 | A-levels | Job-seeking |
| A2 | 69 | Secondary school | Retired | F2 | 35 | n/a | Job-seeking | ||
| A3 | 63 | Vocational school | Retired | F3 | 49 | A-levels | Job-seeking | ||
| A4 | 70 | Vocational school | Retired | F4 | 50 | Vocational school | Job-seeking | ||
| Group B | B1 | 46 | University | Social counselor | F5 | 48 | A-levels | Job-seeking | |
| B2 | 76 | A-levels | Retired | Group G | G1 | 37 | University | Executive staff | |
| B3 | 46 | University | Social counselor | G2 | 49 | University | DSGVO officer | ||
| B4 | 70 | A-levels | Retired | G3 | 44 | Secondary school | Training counselor | ||
| Group C | C1 | 60 | Apprenticeship | Personnel counselor | G4 | 58 | University | Executive staff | |
| C2 | 60 | University | Personnel counselor | Group H | H1 | 37 | University | Team lead | |
| C3 | 51 | Apprenticeship | Job trainer | H2 | 56 | Apprenticeship | Job trainer | ||
| Group D | D1 | 65 | University | Business consultant | H3 | 45 | University | Job trainer | |
| D2 | 53 | University | Retired | H4 | 43 | University | Job trainer | ||
| D3 | 52 | University | Business consultant | H5 | 60 | University | Administrative staff | ||
| Group E | E1 | 36 | University | Graphic designer | |||||
| E2 | 32 | Apprenticeship | Job-seeking | ||||||
| E3 | 40 | Apprenticeship | Job-seeking |
| ID | Age | Education | Occupation | ID | Age | Education | Occupation |
| S1 | 74 | University | Retired | S7 | 40 | University | Job trainer |
| S2 | 29 | A-levels | Nurse | S8 | 43 | University | Rehabilitation counselor |
| S3 | 28 | University | Social counselor | S9 | 44 | University | Social center manager |
| S4 | 29 | University | Doctoral student | S10 | 52 | University | Rehabilitation counselor |
| S5 | 37 | University | Administrative staff | S11 | 59 | University | Social center manager |
| S6 | 28 | University | Job-seeking | S12 | 39 | University | Education program manager |
4. Results
In this section, we present our results, structured according to our research questions: how question-driven, modular explanations support understanding in individual and group settings (RQ1, Section 4.1) and how groups used explanations666We describe every question and answer pair as one ”explanation” and their entirety as a ”collection of explanations” (Section 3). to solve the study tasks and form opinions about the system’s deployment (RQ2, Section 4.2). Participant labels denote the study setting (focus group: A–H / single interviews: S) and the individual participant, as listed in Tables 1+2.
4.1. RQ1-Explanations: How does a question-driven, modular explanation design support AI novices’ understanding in individual and group settings?
Participants received a collection of 36 explanations to solve the study tasks and decide about the system’s deployment individually and (in focus groups) collectively (Section 3). The following subsections describe the explanations’ effect on participants’ understanding in both settings (4.1.1), design aspects that impeded understanding (4.1.2), the time limit’s effect on prioritization of information (4.1.3), the interaction differences between single and group settings (4.1.4), and participants’ ratings of explanations’ helpfulness and influence (4.1.5).
4.1.1. Explanations had mixed effects on participants’ understanding.
| Focus gropus | ||||||||||||||
| Increased understanding | F2 | C3 | F1 | B4 | A2 | H1 | H2 | E1 | H4 | |||||
| Self-reported understanding | +++ | ++ | ++ | + | + | + | + | + | + | |||||
| Decision confidence | - | ++ | = | +++ | + | + | + | = | = | |||||
| Unchanged understanding | H3 | B3 | A3 | G1 | H5 | A1 | A4 | C2 | D3 | E2 | C1 | E3 | F4 | F5 |
| Self-reported understanding | = | = | = | = | = | = | = | = | = | = | = | = | = | = |
| Decision confidence | +++ | ++ | + | + | + | = | = | = | = | = | - | - | - | - |
| Decreased understanding | D1 | G4 | D2 | G2 | B1 | B2 | F3 | G3 | ||||||
| Self-reported understanding | - | - | - | - | - - | - - | - - - | - - - | ||||||
| Decision confidence | + | + | = | = | +++ | = | + | = | ||||||
| Single studies | S6 | S12 | S8 | S7 | S4 | S9 | S1 | S3 | S11 | S5 | S10 | S2 | ||
| Self-reported understanding | +++ | + | + | = | = | = | = | = | = | - | - | - - | ||
| Decision confidence | ++ | + | = | +++ | + | + | = | - - | - - | ++ | = | - |
| Focus groups | Single interviews | |||||||||
| ID | 1st decision | Collective decision | 2nd decision | ID | 1st decision | Collective decision | 2nd decision | ID | 1st decision | 2nd decision |
| A1 | No | Yes | F1 | Yes | Yes | S1 | Yes | Yes | ||
| A2 | No | Yes | F2 | Yes | Yes | S2 | No | No | ||
| A3 | Yes | Yes | F3 | Yes | No | S3 | No | No | ||
| A4 | No | Yes | Yes | F4 | No | No | S4 | No | No | |
| B1 | No | No | F5 | Yes | No | Yes | S5 | Yes | No | |
| B2 | Yes | No | G1 | Yes | Yes | S6 | No | No | ||
| B3 | Yes | No | G2 | No | No | S7 | No | No | ||
| B4 | Yes | No | No | G3 | Yes | No | S8 | No | No | |
| C1 | No | No | G4 | No | No | No | S9 | No | No | |
| C2 | No | No | H1 | No | No | S10 | Yes | Yes | ||
| C3 | No | No | Yes | H2 | No | No | S11 | Yes | No | |
| D1 | No | No | H3 | No | No | S12 | No | No | ||
| D2 | No | No | H4 | No | No | |||||
| D3 | No | No | Yes | H5 | No | No | No | |||
| E1 | No | No | ||||||||
| E2 | No | No | ||||||||
| E3 | No | No | Yes | |||||||
Participants reported their understanding before and after receiving and discussing the explanations. Table 3 shows that self-reported understanding decreased and increased for comparable numbers of participants, while also often remaining unchanged. Participants gave divergent feedback on how they perceived the explanations, describing their structure and design as ”nicely presented” (A2, C2), ”good to get an overview” (C3, H4), ”well-selected” regarding the questions (G2), ”active and controllable” (S8), and as a good primer (F4). Critical comments described them as ”too much” (D1, S4), ”at first overwhelming” (H5), ”confusing” (B1, D1), ”theoretical” (D2), and ”quite complex” (C2). However, this was not always seen as an issue: ”I believe that it is complex. But I also believe that it needs to be.” (D2).
Participants gave nuanced reasons why their perceived understanding decreased. Several focus group participants stated they did not fully understand the technical explanations, including numbers and a formula (B3, C3, G4). In contrast, participants in single interviews with comparable backgrounds met fewer challenges with technical explanations, often stating that they appreciated applying this knowledge when solving the tasks (S2, S8, S11). In these cases, the setting likely mediated the ease of interaction, as participants in single interviews could concentrate without distraction (further described in Section 4.1.4+ 5.2). However, irrespective of the setting, confrontation with mathematical details seemed to re-adjust participants’ benchmarks for ”high understanding” and calibrate their perceptions. Likewise, several participants stated that the different subtopics and levels of detail in the explanations helped them identify the information they were still missing to judge the whole system (B1, F3, G3). Both observations indicate that the explanations revealed participants’ blind spots and helped them to overcome the ”illusion of explanatory depth” (Rozenblit and Keil, 2002) – an overestimation of one’s understanding due to limited knowledge and misled intuition – which is exemplified by S2’s comment: ”The more I look at the explanations now, the more I realize that I simply assumed a lot of things but didn’t actually know how it worked.” In short, participants gauged their understanding relative to the amount of information available. That is, since participants only had time to read through parts of all available explanations and were aware that they had not seen all of them, they reported a decrease in perceived understanding when it, in fact, improved. The implications for measuring explanations’ effect on understanding are discussed in Section 5.
4.1.2. The explanation design was not equally suited to every participant group.
In contrast to participants whose perceived understanding was improved and calibrated by the explanations, we observed that the explanations did not fully support understanding or interaction for some groups. In these cases, the explanations’ language and textual format were listed as impeding factors. The language style was described discrepantly as ”well understandable” (P3, E3, H1) but also ”very difficult” (E2, F2) and ”too technical” (B3, B4) depending on participants’ background. Crucially, in Groups E and F (decision subjects), participants stated that the language posed challenges. E2 compared the explanations to official letters they receive: ”I get letters that I don’t understand from the court. I understand every single word, but I don’t understand the context.” In line with this, job counselors in Groups C, G, and H stated that explanations in simple language would be required if they were supposed to be useful for their job-seeking clients. Similar comments were made about the explanations’ heavy reliance on textual information, described as demanding and requiring plenty of concentration. While previous work found that textual explanations can be effective despite users’ dispreference (Szymanski et al., 2021) or are even preferred (Weitz et al., 2024), a more interactive and visual design would likely facilitate interaction: ”If there’s a program where, when you click on it, it opens as a window and you can read about it, that would be great, and you can deal with it at your own pace.” (H5) Both, simpler language options and an interactive version, are thus useful suggestions for improving the explanation design to ensure that explanations are helpful for a broader range of stakeholder groups. Further, Section 4.2.1 describes how these understanding challenges impacted group interaction.
4.1.3. Time limits led to forced prioritization of information, with different effects on participants.
All participants had 40 minutes to become familiar with the explanation structure, request new explanations, solve the given tasks, and, in focus groups, make a collective decision (Table 4). The time limit was meant to incentivize participants to prioritize information and, where possible, work together to process larger amounts. This succeeded for most focus group’s participants, who stated that ”when you have little time, the teamwork is extremely good” (B2) and that ”even a brief exchange is enough, otherwise it becomes too much” (C3). However, the time limit also led to induced stress, especially in single interviews (S2, S6, S7, S8, S9, S11). S9 commented that this reminded them of their work as a social center manager: ”Basically, the time to read about something new does not exist.” (S9). This points to the need for brief and compact synopses complementing the detailed explanations. Regarding prioritization, many participants liked choosing explanations according to their own interest, which gave the feeling of being in control (S8), provided a direction for exploration (H1), and justified skipping explanation categories in favor of others (S6, S9). Allowing for the self-directed selection of relevant information by users could thus be beneficial to encourage interaction and engagement.
4.1.4. Some single participants worked through surprising amounts of explanations.
Compared to focus groups, participants in single interviews often requested and engaged with equal or even higher amounts of explanations (as depicted in Figure 4). Often, participants in single interviews requested ”the whole context” (S3) or ”all of level 3” (S4) of one category, appearing to deal with the amount of information easily. A likely reason for this is that single interviews supported more focused and in-depth interaction with explanations compared to group settings. Participants in single interviews also performed better in the study tasks overall and, as mentioned above, often calculated the job-seeker’s precise employment chance in task 2, which required following a multi-step process. This is noteworthy, as none of the groups, aside from the pilot groups, attempted to calculate the precise chance. This cannot be explained by a difference in educational or professional background, as even focus groups with university graduates in domain expert professions did not perform these calculations, again indicating that the reason might be the difference in supported interaction. For example, the parallel processing of information might have impeded understanding of the system’s entirety in groups, as described by G3: ”You only pick out the most important things. And then you haven’t gathered information about the thing as a whole.” This highlights the benefits and trade-offs of both settings, suggesting that individual engagement with explanations might be helpful before discussing them in a group setting. The implications of this are discussed further in Section 5
4.1.5. All explanation categories were deemed helpful and influential, but system details and context were most often mentioned.
In the second self-reports, participants were asked which explanation category was most helpful for their understanding, most influential for their decisions, and whether specific explanations stood out. Regarding understanding, participants in focus groups responded with a rather even distribution over all categories, as depicted in Figure 5, and frequently mentioned that ”all of them [are relevant]… I don’t think you can leave anything out, really” (D3). Both participants in focus groups and single interviews further stated that system details were crucial, as they were perceived as ”tangible” (S6), ”concrete” (S8), and to be ”the devil in the detail” (B4). Specifically, participants found explanations about the features and weighting most important (D1, G1, H3, S5), which G4 commented with: ”That is the central point, the basis of the whole system.” Even though context was less often reported as an influential explanation category, participants requested these explanations most often (Figure 6) and frequently mentioned affected persons’ inability to contest decisions (C1, D1, E1) and the system’s political background (A1, S3) as relevant – both belonging to context. Notably, participants in single interviews found data much less helpful and less influential, stating, e.g., that GDPR would likely cover the important aspects of it (S8) or that they prioritized another category in the time available: ”I didn’t deal with it in such detail […]. You read it [the base level] and already feel informed. As is often the case with data protection, ’yeah, close enough’.” (S6). Other participants stated that the technical details were not too important for them but that they instead focused on how the system was used practically and which values it represented (S3, S10), corresponding to the categories usage and context.
4.2. RQ2-Deliberation: How do the explanations support group interaction with information and deliberation processes?
In this section, we describe how our explanations supported group collaboration in exchanging information (Section 4.2.1) and in discussing and forming opinions about the system’s deployment (Section 4.2.2).
4.2.1. Explanations support collaborative success if group dynamic is constructive.
We identify several processes that groups used to interact with information collectively. Our descriptions of these interaction processes include splitting (by category, order, priority), sharing (swapping, reading out loud, proposing solutions), analyzing (discussing, explaining), and combining (merging, locating, correcting). The degree to which groups used these processes appeared to depend on the familiarity between group members, their domain expertise, and their understanding. In the following, we analyze three exemplary conversation excerpts that showcase different forms of interaction in focus groups while also demonstrating the diversity in collective interaction. For each excerpt, we describe the interaction processes that emerged and then connect them to mechanisms that drive collaborative failure or success (marked by quotation marks) as outlined by Nokes-Malach et al. (2015).
Case 1: Information sharing leads to increased engagement and construction of common ground. Group G was composed of participants who hold executive positions in a civil society organization and are responsible for the management and training of employees. Participants in this group showed high engagement with information, frequently sharing explanations and exchanging their perspectives by tapping on their professional experience. The excerpt begins several minutes after participants received the explanations just as G4 examines an explanation comparing the algorithmic score for two job-seekers:
G4: That’s bad, the two of them. Look, I’ve got this [system details-C-base], ”What chances would the system calculate for Martin and Schifteh?” (sharing: reading out loud)
G2: Schifteh is probably worse off, isn’t she? (analyzing: discussing)
G4: Schifteh has 30% chance of employment and Martin 52%, even though Schifteh has a degree and would be working in the IT sector. And Martin has compulsory schooling and works in the cleaning sector. Martin’s chances of employment are almost twice as high as Schifteh’s. (sharing: reading out loud; analyzing: explaining) […]
G3: I think that’s a bit weird. […] Because if she can speak English very well and has the specialist knowledge that our IT sector needs… (analyzing: discussing)
G4: She gets two minuses for living in [a specific city district777Information was omitted to adhere to anonymization policy and will be re-inserted for the final version.]. People who live in [this district], or in a district that is not well regarded, have totally negative… (analyzing: explaining)
G1: Yes, and here you have it in writing, I’ll have to look at that too. (sharing: swapping)
Framed in the context of mechanisms that drive collaborative success, G4 first draws the group’s attention to an explanation they deem relevant, describing the algorithm’s treatment of two different cases of job-seekers. G4 reads the information to the group, which prompts G2 and G3 to contribute their own perspectives, ”increasing engagement” while also ”cueing their prior knowledge”. As a result, G1 is also engaged by the information and articulates interest in looking at it. In combination, the interaction has established ”common ground” by sharing information the group can use for further collaboration. Interactions like these were common in Groups G and H (job counselors and trainers), indicating that the explanations’ format and content suited these compositions.
Case 2: Discussion of solutions leads to correction. Group B consisted of both staff members and volunteers of a civil society organization. The group interaction was mostly driven by one participant who requested and shared most of the information and thus acted as a guide and informer, potentially driven by her role as deputy manager. While this led to a more one-sided dynamic, the other participants later appreciated outsourcing part of the understanding process to B3. In the excerpt below, B3 proposes a solution for task 2 (chance of employment that the system assigns Harald G.; group low, medium, or high) that is then discussed:
B3: I chose “group low” for task 2. (sharing: proposing solutions)
B4: I put him [Harald G.] in ”group medium”. (sharing: proposing solutions)
B1: Why not medium, B3? (analyzing: discussing) […]
B3: With 49 years he has one minus. And the knee problems could be a “health impairment”. (analyzing: explaining)
B2: But if he has a job where he sits, that doesn’t matter. (analyzing: discussing)
B3: But I don’t see that this is considered anywhere. And I think the fact that he hasn’t been employed much in the last 4 years has a very negative effect. What information do you have that speaks against that? (combining: correcting; analyzing: discussing)
B1: Because with 50, he would still be in ”group medium”. […] Then he’s male and from [this country888Information was omitted to adhere to anonymization policy and will be re-inserted for the final version.]. (analyzing: explaining)
B3: Male is only 0, the citizenship is also only 0. That doesn’t make up for anything. (combining: correcting)
[B1 agrees] […]
B4: Then I’ll correct it to low again. (combining: merging)
The group members in this situation ”negotiate multiple perspectives” on the correct solution, notice discrepancies in their ”common ground,” and ”re-expose” themselves to information to ”correct errors” in their knowledge. B3 led the interaction and corrected other participants repeatedly, which could have caused adverse reactions from them (e.g., due to being ”evaluated”). However, B2 and B4 in fact later articulated appreciation for this interaction, stating that ”B3 filtered that out well” (B2) and that ”it was helpful that you [B3] always read it out and corrected it straight away” (B4). Correction was therefore seen as constructive to the discussion, possibly because of B3’s leading role in the organization and because participants already knew each other.
Case 3: Attempted error-correction fails to create common ground. Group E was composed of people who were job-seeking at the time of the interview or in the past. Members of this group were not familiar with each other. Interaction was mostly fragmented and participants rarely shared or discussed information. In the excerpt below, participants were asked about their responses to task 4 (whether Harald G. is able to contest the system’s decision):
E2: I say you can always object. Even if you’re locked up, you can go to court somewhere. (sharing: proposing solutions)
E3: Yes, I think so too. (sharing: proposing solutions)
E1: Well, I’ve already read that [the explanation], and it says that you can’t appeal against it. But it says that you could if a legal basis were first created, which doesn’t exist at the moment. (combining: correcting; analyzing: explaining)
Investigator: Okay. Is there anything you [E2, E3] would like to change now, based on this information?
[E2 & E3 indicate that they do not]
Investigator: You would say that he can contest the decision?
[E2 & E3 agree]
Despite E1’s reference to an explanation and their attempt to ”correct errors” in what E2 and E3 believe to be the solution, the interaction fails to establish ”common ground” or ”negotiate multiple perspectives”, as the different viewpoints are neither elaborated nor discussed collectively. Possibly, E2 and E3 did not believe that E1 or the explanation were correct, or they felt embarrassed for being corrected. Further, E1 had no particular motivation to persuade the other participants of the correct solution, as E1 later self-described as ”introverted” and stated that they usually preferred individual over group settings. Comparable mechanisms unfolded only in Group F, where participants tended to read explanations themselves or disengage rather than interact with other participants. This might have had multiple reasons: First, both Groups E and F indicated a low interest in the topic of labor; second, participants in both groups had trouble understanding the explanations due to language barriers and the heavy reliance on text; and third, participants in both groups were unknown to each other. In combination, this might have created a group dynamic that was not constructive, i.e., neither argumentative (Mercier and Sperber, 2011), solution-oriented (Niculae and Danescu-Niculescu-Mizil, 2016), or familiar (Johnson and Johnson, 1985), which disrupted collaborative interaction with the explanations as well as discussion processes.
Summary. The conversation excerpts in cases 1 and 2 demonstrate that the explanation design supported various collaborative interaction processes. These processes led participants to create common ground and find a task solution through error correction, which are markers of ”collaborative success” (Nokes-Malach et al., 2015). In contrast, case 3 demonstrates that the explanations did not always achieve their purpose of facilitating interaction, highlighting that the interplay of explanation format and group dynamics must be considered in the design of explanations in group settings.
4.2.2. Explanations in group settings allow for informed discussion of pivotal topics and opinion formation about system deployment.
For most participants, deciding about the employment prediction algorithm’s deployment was a clear choice: 7 out of 8 groups voted ”No”, and a majority of participants also voted ”No” in their individual decisions (Table 4). Confidence in these decisions increased strongly after the explanation phase (Table 3), with most participants indicating that they now felt better informed. Comparing single interviews with focus groups, two effects were notable: First, a third of focus group participants changed their vote in the second self-report, and second, participants in single interviews stated repeatedly that they would have found a discussion helpful.
This is further underscored by participants’ feedback on the focus groups: Most felt included in the discussions (26 of 31) and were influenced in their decisions by them (18 of 31). Specifically, participants appreciated the group size, which allowed for balanced contributions and productive discussion (C1, D1, E2), and valued the exchange of perspectives: ”Everyone understands a little differently and sees a different focus, I think it’s really important to talk about that.” (H5). We see this as evidence that the explanations were suitable for the collaborative setting and supported group deliberation. This section thus aims to describe the synergy effects that emerge when combining explanations with group discussion and how this affected participants’ perceptions. In the following, we describe three conversation excerpts in which explanations lead groups to discuss their core concerns and beliefs.
Case 1: Group discusses changes of perspectives on the system’s deployment. Group B was composed of staff members and volunteers of a civil society organization. Three participants in this group changed their votes from ”Yes” in the first report to ”No” in the second report. In the excerpt below, B2 gives their perspective on the system’s deployment, and B3 responds with theirs, referring to explanations about the system’s features and weightings (System details, Base, B + Level 2, C), which fuel the discussion that follows:
B2: I’m skeptical, but I’m still in favor of introducing it. Because it could be an aid and a relief for the staff working there.
B3: I was originally in favor for these reasons, but since I’ve seen these parameters, I would be very much against it. Because I think there’s a lot of ideology in it. I think it’s no longer acceptable that men are favored over women and that duty of care only applies to women – this comes from a time that should be long gone.
B2: Those are strong arguments.
B3: The things that come out are so absurd as well. For example, the apprenticeship of our Harald [the example job-seeker] was rated positively. He can’t use the apprenticeship at all for retraining. These things are given far too little consideration. […] As much as I like the idea, I don’t like the parameters.
B1: Did you vote yes first?
B3: I ticked yes at first, but I was really shocked at what was in there. […]
B1: What I’m wondering is, what would be the real benefit of introducing the system? What could be found out with much less effort? After all, employees should be able to see which degree you have quite quickly.
B4: It’s a grid, a structure for the people who work at the agency so that they can quickly find a box.
The excerpt highlights how the provided information changed B3’s attitude toward the system and how this change, in turn, affected the group discussion. The combination of explanations and group discussion here had a major effect on participants’ perceptions, as three participants in this focus group changed their vote from ”yes” in the first self-report to ”no” in the second. Interactions like B2’s appreciation of B3’s arguments and B1’s questions about the system’s benefits depict processes of constructive group deliberation directly supported by the explanations.
Case 2: Group debates the meaning of ”unalterable” facts. Group D was composed of participants who were neither job-seekers nor domain experts and thus would be affected only indirectly by the system’s deployment. Even so, when D1 referred to an explanation describing how two different job-seekers would be judged by the system (System details, Base, C), it led to a discussion about the objectivity and adequacy of using data to represent a person:
D3: I believe that the system can form the initial basis, based on the unalterable facts, which are of course weighted, but then it has to be enriched by a human being. […]
D1: But I don’t believe that there are unalterable facts – well, not in this area. It’s all a question of representation and the lens through which you see the world.
D3: When the job-seeker says, I only have four years of elementary school, then that’s four years of elementary school… […]
D2: That doesn’t mean that he can’t still be a very educated person.
D3: But that is hard to sell to an employer, right? […]
D2: I’m skeptical about the data. You [D3] said the ”basis”, I think there are cracks in this basis. And I’m afraid of these cracks, that something will be pre-determined…
D3: But the human decision is always subjective.
D2: That has to be weighed up. On the one hand, you have the arbitrariness of the individual employee, yes, and on the other hand you have an incomplete picture of a person.
D1: Or a false image.
D2: An incomplete one, I would say.
In this excerpt, the discussion surfaces discrepancies in how participants perceived the factuality and value of data. While the discussion in this focus group did not lead to a consensual decision in the time given, we still see this exchange as an important process in the deliberation on public AI systems. The merit of this discussion was later acknowledged by D1, who found the explanations confusing but stated that these exchanges were the ”centerpiece” and the most intriguing part. This highlights the role of explanations in the surfacing and debating of divergent opinions.
Case 3: Group discusses consequences of the system’s deployment for affected people. Group H was composed of staff members of a civil society organization specializing in helping young adults to take up apprenticeships. While all participants were clearly against the system’s deployment, the group decision task still led to a discussion about why specifically they opposed the system. The excerpt begins with H2’s comment about an explanation describing that job-seekers do not have a right of contestation (Context, Level 3, B):
H2: Just the fact that you can’t contest the decision, that’s a no-go.
H1: I think that having such systems introduced by institutions that themselves have very little idea of the real work – they may even want to help and support work systems, but that completely misses the mark because they are making decisions that are again driven by economic interests.
H4: And the life of the people who are then affected by the system is not represented at all, the psychosocial situation of the people is not taken into account.
H3: It is if they are depressed…
H4: Then it’s minus, minus, exactly. It simply deprives people of the opportunities they deserve. And that’s probably also my personal experience, I’ve worked with long-term unemployed adults who also had physical or mental impairments. The problem is that the longer I’m out of the job market, the fewer opportunities I have and people are already in a bad enough situation. And then you’re totally stuck in this [algorithmic] system. I just think that’s unfair.
H2: In terms of labor market policy, it may be justified, but in terms of social policy, it’s totally wrong.
The excerpt shows that even participants with unanimous attitudes toward the system found value in discussing and sharing their reasons for disapproving its deployment. H4 in particular appeared to be incensed by the system’s representation of soft factors. Their account of working with long-term job-seekers was later appreciated by the other participants, who stated that this perspective was valuable to them. The explanations in this case served to facilitate group conversation about the meaning and consequences of the system’s deployment and encouraged participants to share their views, stances, and experiences.
Summary. While Section 4.2.1 describes that explanations were effective in supporting group collaboration, this Section shows that their role in providing information for discussions between participants is equally important. Explanations helped participants in debating topics that are pivotal to the acceptance, trust, and application of ADM systems, in effect supporting their collaborative reasoning (Moshman and Geil, 1998). In this way, the explanations support what Kawakami et al. (2024) describe as ”deliberations around whether and under what conditions to move forward with developing or deploying” a public sector AI tool. This deliberation is important, as it allows participants to express and reflect on their opinions, feel heard, and consider other perspectives when making decisions.
5. Discussion
In this section, we discuss the implications of our findings for the design of explanations for AI novices (Section 5.1) as well as suggestions for their applications in collaborative settings (Section 5.2), which are summarized in Figure 7.
5.1. Reflections on the explanation design
Explanations improved or calibrated participants’ understanding, but measurements should be reconsidered. In Section 4.1.1, we described that while participants gave discrepant responses in their understanding self-reports (Table 3), they rarely felt that their understanding truly decreased or remained the same. Instead, we argue that the explanations changed their frame of reference for ”understanding” by revealing blind spots and making participants realize that their initial understanding was worse than assumed. Drawing from definitions in the cognitive sciences, the explanations thus supported participants in developing understanding by i) making participants realize that their initial understanding was incomplete (Rozenblit and Keil, 2002) and ii) supporting them in locating and closing relevant understanding gaps (Keil, 2006). The effect of participants reporting decreased perceived understanding following the explanations can be attributed to the fact that these explanations led them to attain a ”partial understanding” (Keil, 2019), but also showed them the difference to a complete, comprehensive understanding. The difficulty of self-assessing one’s understanding is a known issue (Rozenblit and Keil, 2002) and previous studies in XAI have shown similar effects caused by white-box explanations (Cheng et al., 2019), which, due to their high information density, led to increases in objective understanding but decreases in self-perceived understanding. This is not necessarily an issue regarding participants’ understanding, and together with the substantial increases in decision confidence, it can be seen as evidence that our explanations effectively improved and calibrated participants’ understanding. However, it also demonstrates that the relativity of internal cognitive states should be accounted for when measuring understanding. For this reason, recent work in XAI has proposed measurement based on abilities (Speith et al., 2024). This approach appears promising, considering that our single interview participants often found applying their knowledge in the study tasks helpful. Still, it likely requires modification when used in collective settings, as, for example, participants in focus groups never calculated the precise employment chance in task 2. The operationalization of the abilities-based approach for both individual and group settings is thus a direction for future research.
Participants considered all four explanation categories relevant and appreciated technical and contextual information. Our explanation design covered the categories data, system details, usage, and context. In the second self-reports, participants indicated which of these categories were ”helpful for their understanding” and ”influential to their decisions” (Section 3.3). In Section 4.1.5, we describe that focus groups reported that all categories helped their understanding but that system details more often than other categories influenced their decisions. In contrast, participants in single interviews reported all categories equally often as helpful and influential except data. Possible explanations for this effect include the forced prioritization due to the time limit, which led participants in single interviews to skip information about data as it was believed to be less relevant. In comparison, some focus groups proceeded chronologically through the explanations and thus spent more time on data as the first category. Further, the findings highlight that participants welcomed non-technical information about the system’s intention, ethical values, and practical usage (Figure 6) but also assigned high importance to the model’s exact feature weights. We see this as confirmation that participants’ information needs cover the entirety of an ADM system and that explanations for AI novices should aim to provide both technical (data, system details) and contextual (usage, context) information.
Explanations for decision subjects require a co-design approach. In Section 4.1.2, we describe that our explanations were most effective for participants who did not face challenges with either the language, large amounts of text, or the explanations’ structure. While these attributes are not entirely congruent with participants’ education levels, we observe that university-educated participants faced comparatively little difficulty interacting with the explanations, as the conversation excerpts in Sections 4.2.1+4.2.2 show. The differences were most apparent in Groups E and F, where participants stated that the explanations did not help them fully grasp the system as they required a high language and reading proficiency. We see this as evidence that a co-design approach similar to Weitz et al. (2024)’s work is required to build explanations for groups composed like Groups E and F, as it might be difficult to identify which content and format of explanations are required by decision subjects. On the other hand, the findings demonstrate the explanation design’s suitability for groups of domain experts, who also stated that this format of conveying information would be useful as a training measure.
Suggestions for improving the interaction design and combining individual and collective settings. In Sections 4.1.1 and 4.1.2, we described that according to participants’ feedback, the explanation design offers comprehensive and flexible information selection and encourages self-directed and active exploration, but that it also has a high access threshold and requires adjustment to the subdivided structure. For focus group participants, this design resulted in participants splitting and swapping explanations, sharing relevant parts, and discussing surprising information, which overall resembled an associative approach that supported group interaction but also led to a more fragmented picture of the system. In single interviews, participants could select and interact with information at their own pace and in a focused manner but, on the other hand, missed exchange and external perspectives. These effects can be described in the context of collaborative success and failure (Nokes-Malach et al., 2015), which states that groups require extra cognitive effort to coordinate and manage interactions, which can lead to memory limitations, information retrieval disruption, and blocking of thought processes if these coordination costs are too high. In contrast, individuals do not incur these cognitive costs, but of course can also not benefit from collaborative success mechanisms such as increased working memory or complementary knowledge. A future iteration of the explanation design could thus follow an individual with a collective interaction phase to combine the advantages and offset limitations. Further, a digital version would facilitate overview (summaries) and navigation (menus) while allowing for simple language options and cross-references. Lastly, the essential information of each explanation category, which in our design was contained in the base levels (Section 3.2), could use a balance of textual and visual design to ensure that essential information does not rely only on textual understanding.
On a side note, we emphasize that identifying explanation subtopics and splitting them up into levels of detail is challenging. Our explanations’ structure provides different levels of soundness and completeness (Kulesza et al., 2013), but deciding how the available information is allocated into this structure requires a subjective choice. In our study, detail levels 2 and 3 were supposed to convey more background rather than essential information gradually. Still, some participants stated that the most critical information for them resided in level 3. We see avenues for future work in selecting and hierarchically structuring information to be included in explanations for ADM systems, such that this structure allows for exploration while not obscuring essential or complex information.
5.2. Benefits and challenges of using explanations in collaborative settings
In Sections 4.2.1+4.2.2, we described that focus groups’ interactions with information led to ”collaborative success” (Nokes-Malach et al., 2015) when the group dynamic was constructive, i.e., argumentative (Mercier and Sperber, 2011), familiar (Johnson and Johnson, 1985), and solution-oriented (Niculae and Danescu-Niculescu-Mizil, 2016), and that explanations can inform conversations about central topics in public AI deployment (Kawakami et al., 2024). In the following, we discuss these findings by outlining the resulting benefits and challenges for collaborative XAI.
Benefits: Explanations support collaborative reasoning, problem-solving, and deliberation. In the best cases, focus groups in our study had an open, communicative, and investigative atmosphere that facilitated the up-taking, sharing, and discussion of information. In these settings, the explanations showed their strengths by allowing for the distribution of reading tasks among group members, providing high levels of detail and breadth if needed, and offering different viewpoints that could be used as argumentative and conversational starting points. In this sense, the explanations fulfilled their aim of supporting collaboration and informed decision-making around deploying a public AI system (Kawakami et al., 2024) by creating common ground that participants used as a basis for deliberation. This deliberation process is the differentiating factor compared to ”one-to-one” (Naiseh et al., 2021) explanation settings, which might allow for more concentrated learning (Bloom, 1984), but miss the exchange of knowledge and perspectives that could become central for societal discourse about public AI (Züger and Asghari, 2023). While this collaborative atmosphere has several preconditions, such as containing cognitive biases (groupthink (Janis, 1972), equality bias (Naiseh et al., 2024)) and requiring trust between group members (Johnson and Johnson, 1985), our findings suggest that explanations can be an effective enabler in supporting these crucial group interactions. Regarding learning and understanding, XAI could further benefit from connecting to research showing that small groups can exhibit collective rationality (Moshman and Geil, 1998), outperform the wisdom of the crowds (Navajas et al., 2018), and solve tasks correctly even if no group member knows the solution beforehand (Smith et al., 2009).
Challenges: Explanations must adapt to different groups and require constructive dynamics. While our explanations were effective in supporting interaction and opinion formation in most focus groups, we also observed several challenges resulting from group compositions and trust between participants. While the explanations typically worked well in groups of domain experts, especially G (executive staff) and H (job trainers and counselors), special interaction situations occurred in Groups B and H due to two participants’ leading roles in the organizations. This could have resulted in disrupted interaction due to fear of being evaluated by someone perceived as superordinate, but, in fact, had no negative impact on collaboration. In Groups E and F (decision subjects), however, the unfamiliarity of participants and the explanation’s high access threshold led to a lack of interactions and discussion, indicating that the explanation did not fully support these compositions. In these groups, a lack of trust between participants likely disrupted interactions when participants were unknown to each other, amplifying effects such as the fear of being evaluated (Nokes-Malach et al., 2015). This underscores the importance of creating trust between group members in collaborative XAI settings (Johnson and Johnson, 1985). Intuitive measures could be the introduction of a simple task that the group solves collaboratively before engaging with explanations, such as the Wason card selection task (Wason, 1968), or creating an atmosphere that allows participants to interact in roles (e.g., proponents and opposition) to facilitate discussion without revealing personal opinions. Future work should examine how such measures can be incorporated into explanation design to support interaction in groups of comparable compositions. Lastly, and surprisingly, we did not observe many cognitive biases such as ”groupthink” (Janis, 1972) (aiming for consensus at the cost of negotiation and alternatives) or the ”equality bias” (Naiseh et al., 2024) (downplaying one’s expertise to weigh everyone’s opinion) in our study. In one example of groupthink, participants in Group A changed their vote to ”Yes” to reach a group decision, perhaps driven by their (stated) lack of being personally affected by the system’s deployment. However, this is contrasted by participants in Group D, who debated at length about the system’s deployment without reaching a consensus, despite also not being affected personally. Potential measures to avoid groupthink in discussion could thus be to encourage debate, which again could be the introduction of roles to improve the ”dialectic argumentation” (Mercier and Sperber, 2011), and to explain the system in a way that makes it more personally relevant to participants (Wiggins and McTighe, 2005), e.g., by emphasizing connections to their own experiences.
5.3. Explanation design suggestions
We summarize the implications of our findings in the form of suggestions for the design of explanations that are suited to AI novices in individual and group settings, as depicted in Figure 7.
6. Limitations
In this section, we briefly describe limitations of our study that the reader should consider. Due to the limited sample size, we did not analyze the impact of sex and/or gender on our results, limiting the results’ generalizability regarding these aspects but not their overall validity. Further, our participants were recruited from organizations and networks in the same geographical region, perhaps resulting in regional or cultural biases. The presented use case is embedded in a specific sociotechnical context (Ehsan et al., 2023a) that might affect participants’ understanding and perceptions (e.g., perceptions might differ between employability prediction and credit approval), and thus a change in the domain might also change the explanations’ effect. However, this does not limit the transferability of the explanation design, which can be seen as a template that can be adjusted to other use cases. We further note that our participant sample has a bias toward university education in the single interviews, which we addressed by comparing these participants mostly with university-educated participants in the focus groups.
7. Conclusion
In this paper, we proposed and empirically validated a question-driven, modular explanation design for individual and collaborative settings. We conducted an interview study with AI novices in 8 focus groups and 12 single interviews, analyzing the explanations’ effect on understanding and decision confidence, participants’ perceptions of key information, and the interaction processes in both settings. We found that explanations were effective at supporting and calibrating participants’ understanding and at informing individual and collective decision-making. Our findings further showed that our explanations support focused and detail-oriented interaction in individual settings while also being suited to be shared and discussed in group settings. In groups with constructive dynamics, interactions with the explanations led to ”collaborative success” (Nokes-Malach et al., 2015) and informed group discussions about pivotal topics regarding the system’s deployment. For groups in which the explanations were less effective, we suggest adjusting the group dynamic, modifying the explanation’s format, and using co-design approaches to better adapt explanations to various needs. Based on these findings, we present suggestions for the design of explanations for AI novices and outline how their application can support collaborative reasoning and discussion. With this work, we aim to contribute an explanation design approach to the field of collaborative XAI and present implications for how explanations can support stakeholder deliberation about public AI systems.
Acknowledgements.
This work has been funded by the Vienna Science and Technology Fund (WWTF) [10.47379/ICT20058] as well as [10.47379/ICT20065].References
- (1)
- Alfrink et al. (2023) Kars Alfrink, Ianus Keller, Neelke Doorn, and Gerd Kortuem. 2023. Contestable Camera Cars: A Speculative Design Exploration of Public AI That Is Open and Responsive to Dispute. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 8, 16 pages. https://doi.org/10.1145/3544548.3580984
- Allhutter et al. (2020) Doris Allhutter, Astrid Mager, Florian Cech, Fabian Fischer, and Gabriel Grill. 2020. Der AMS-Algorithmus: Eine Soziotechnische Analyse des Arbeitsmarktchancen-Assistenz-Systems (AMAS). Technical Report. Österreichische Akademie der Wissenschaften. epub.oeaw.ac.at.
- Ananny and Crawford (2018) Mike Ananny and Kate Crawford. 2018. Seeing without knowing: Limitations of the transparency ideal and its application to algorithmic accountability. New Media & Society 20, 3 (March 2018), 973–989. https://doi.org/10.1177/1461444816676645
- Bansak et al. (2018) Kirk Bansak, Jeremy Ferwerda, Jens Hainmueller, Andrea Dillon, Dominik Hangartner, Duncan Lawrence, and Jeremy Weinstein. 2018. Improving refugee integration through data-driven algorithmic assignment. Science 359, 6373 (Jan. 2018), 325–329. https://doi.org/10.1126/science.aao4408
- Baumberger et al. (2017) Christoph Baumberger, Claus Beisbart, and Georg Brun. 2017. What is Understanding? An Overview of Recent Debates in Epistemology and Philosophy of Science. In Explaining Understanding: New Perspectives from Epistemolgy and Philosophy of Science, Stephen Grimm, Christoph Baumberger, and Sabine Ammon (Eds.). Routledge, 1–34.
- Bertrand et al. (2023) Astrid Bertrand, James R. Eagan, and Winston Maxwell. 2023. Questioning the ability of feature-based explanations to empower non-experts in robo-advised financial decision-making. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’23). Association for Computing Machinery, New York, NY, USA, 943–958. https://doi.org/10.1145/3593013.3594053
- Bloom (1984) Benjamin Bloom. 1984. The 2 Sigma Problem: The Search for Methods of Group Instruction as Effective as One-to-One Tutoring. (1984).
- Bove et al. (2022) Clara Bove, Jonathan Aigrain, Marie-Jeanne Lesot, Charles Tijus, and Marcin Detyniecki. 2022. Contextualization and Exploration of Local Feature Importance Explanations to Improve Understanding and Satisfaction of Non-Expert Users. In Proceedings of the 27th International Conference on Intelligent User Interfaces (¡conf-loc¿, ¡city¿Helsinki¡/city¿, ¡country¿Finland¡/country¿, ¡/conf-loc¿) (IUI ’22). Association for Computing Machinery, New York, NY, USA, 807–819. https://doi.org/10.1145/3490099.3511139
- Bove et al. (2024) Clara Bove, Thibault Laugel, Marie-Jeanne Lesot, Charles Tijus, and Marcin Detyniecki. 2024. Why do explanations fail? A typology and discussion on failures in XAI. http://arxiv.org/abs/2405.13474
- Braun and Clarke (2006) Virginia Braun and Victoria Clarke. 2006. Using thematic analysis in psychology. Qualitative Research in Psychology 3, 2 (Jan. 2006), 77–101. https://doi.org/10.1191/1478088706qp063oa
- Brown et al. (2019) Anna Brown, Alexandra Chouldechova, Emily Putnam-Hornstein, Andrew Tobin, and Rhema Vaithianathan. 2019. Toward Algorithmic Accountability in Public Services: A Qualitative Study of Affected Community Perspectives on Algorithmic Decision-making in Child Welfare Services. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. ACM, Glasgow Scotland Uk, 1–12. https://doi.org/10.1145/3290605.3300271
- Bundesagentur für Arbeit (2021) Bundesagentur für Arbeit. 2021. Bearbeiten von Bewerberdaten durch Träger. arbeitsagentur.de/datei/dok_ba013193.pdf.
- Burns et al. (2023) Alyxander Burns, Christiana Lee, Ria Chawla, Evan Peck, and Narges Mahyar. 2023. Who Do We Mean When We Talk About Visualization Novices?. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 819, 16 pages. https://doi.org/10.1145/3544548.3581524
- Byrne (2023) Ruth M.J. Byrne. 2023. Good Explanations in Explainable Artificial Intelligence (XAI): Evidence from Human Explanatory Reasoning. In Proceedings of the Thirty-Second International Joint Conference on Artificial Intelligence. Macau, SAR China, 6536–6544. https://doi.org/10.24963/ijcai.2023/733
- Capel and Brereton (2023) Tara Capel and Margot Brereton. 2023. What is Human-Centered about Human-Centered AI? A Map of the Research Landscape. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. ACM, Hamburg Germany, 1–23. https://doi.org/10.1145/3544548.3580959
- Chatti et al. (2022) Mohamed Amine Chatti, Mouadh Guesmi, Laura Vorgerd, Thao Ngo, Shoeb Joarder, Qurat Ul Ain, and Arham Muslim. 2022. Is More Always Better? The Effects of Personal Characteristics and Level of Detail on the Perception of Explanations in a Recommender System. In Proceedings of the 30th ACM Conference on User Modeling, Adaptation and Personalization. ACM, Barcelona Spain, 254–264. https://doi.org/10.1145/3503252.3531304
- Chen et al. (2023) Chacha Chen, Shi Feng, Amit Sharma, and Chenhao Tan. 2023. Machine Explanations and Human Understanding. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency (Chicago, IL, USA) (FAccT ’23). Association for Computing Machinery, New York, NY, USA, 1. https://doi.org/10.1145/3593013.3593970
- Cheng et al. (2019) Hao-Fei Cheng, Ruotong Wang, Zheng Zhang, Fiona O’Connell, Terrance Gray, F. Maxwell Harper, and Haiyi Zhu. 2019. Explaining Decision-Making Algorithms through UI: Strategies to Help Non-Expert Stakeholders. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. ACM, Glasgow Scotland Uk, 1–12. https://doi.org/10.1145/3290605.3300789
- Chouldechova (2017) Alexandra Chouldechova. 2017. Fair Prediction with Disparate Impact: A Study of Bias in Recidivism Prediction Instruments. Big Data 5, 2 (2017), 153–163. https://doi.org/10.1089/big.2016.0047
- Conati et al. (2021) Cristina Conati, Oswald Barral, Vanessa Putnam, and Lea Rieger. 2021. Toward personalized XAI: A case study in intelligent tutoring systems. Artificial Intelligence 298 (Sept. 2021), 103503. https://doi.org/10.1016/j.artint.2021.103503
- Convertino et al. (2008) Gregorio Convertino, Dorrit Billman, Peter Pirolli, J. P. Massar, and Jeff Shrager. 2008. The CACHE Study: Group Effects in Computer-supported Collaborative Analysis. Computer Supported Cooperative Work (CSCW) 17, 4 (Aug. 2008), 353–393. https://doi.org/10.1007/s10606-008-9080-9
- de Fine Licht and de Fine Licht (2020) Karl de Fine Licht and Jenny de Fine Licht. 2020. Artificial Intelligence, Transparency, and Public Decision-Making: Why Explanations Are Key When Trying to Produce Perceived Legitimacy. AI Soc. 35, 4 (dec 2020), 917–926. https://doi.org/10.1007/s00146-020-00960-w
- Desiere and Struyven (2021) Sam Desiere and Ludo Struyven. 2021. Using Artificial Intelligence to classify Jobseekers: The Accuracy-Equity Trade-off. Journal of Social Policy 50, 2 (2021), 367–385. https://doi.org/10.1017/S0047279420000203
- DeVito et al. (2018) Michael A. DeVito, Jeffrey T. Hancock, Megan French, Jeremy Birnholtz, Judd Antin, Karrie Karahalios, Stephanie Tong, and Irina Shklovski. 2018. The Algorithm and the User: How Can HCI Use Lay Understandings of Algorithmic Systems?. In Extended Abstracts of the 2018 CHI Conference on Human Factors in Computing Systems (Montreal QC, Canada) (CHI EA ’18). Association for Computing Machinery, New York, NY, USA, 1–6. https://doi.org/10.1145/3170427.3186320
- Dhanorkar et al. (2021) Shipi Dhanorkar, Christine T. Wolf, Kun Qian, Anbang Xu, Lucian Popa, and Yunyao Li. 2021. Who needs to know what, when?: Broadening the Explainable AI (XAI) Design Space by Looking at Explanations Across the AI Lifecycle. In Designing Interactive Systems Conference 2021. ACM, Virtual Event USA, 1591–1602. https://doi.org/10.1145/3461778.3462131
- Ehsan et al. (2024) Upol Ehsan, Samir Passi, Q. Vera Liao, Larry Chan, I.-Hsiang Lee, Michael Muller, and Mark O. Riedl. 2024. The Who in XAI: How AI Background Shapes Perceptions of AI Explanations. https://doi.org/10.1145/3613904.3642474 arXiv:2107.13509 [cs].
- Ehsan et al. (2023a) Upol Ehsan, Koustuv Saha, Munmun De Choudhury, and Mark O. Riedl. 2023a. Charting the Sociotechnical Gap in Explainable AI: A Framework to Address the Gap in XAI. Proc. ACM Hum.-Comput. Interact. 7, CSCW1, Article 34 (apr 2023), 32 pages. https://doi.org/10.1145/3579467
- Ehsan et al. (2023b) Upol Ehsan, Philipp Wintersberger, Elizabeth A Watkins, Carina Manger, Gonzalo Ramos, Justin D. Weisz, Hal Daumé Iii, Andreas Riener, and Mark O Riedl. 2023b. Human-Centered Explainable AI (HCXAI): Coming of Age. In Extended Abstracts of the 2023 CHI Conference on Human Factors in Computing Systems. ACM, Hamburg Germany, 1–7. https://doi.org/10.1145/3544549.3573832
- European Commission (2021) European Commission. 2021. Laying Down Harmonised Rules on Artificial Intelligence. digital-strategy.ec.europa.eu/en/library/proposal-regulation-laying-down-harmonised-rules-artificial-intelligence.
- Fiesler et al. (2019) Casey Fiesler, Jed R. Brubaker, Andrea Forte, Shion Guha, Nora McDonald, and Michael Muller. 2019. Qualitative Methods for CSCW: Challenges and Opportunities. In Companion Publication of the 2019 Conference on Computer Supported Cooperative Work and Social Computing. ACM, Austin TX USA, 455–460. https://doi.org/10.1145/3311957.3359428
- Flügge (2021) Asbjørn Ammitzbøll Flügge. 2021. Perspectives from Practice: Algorithmic Decision-Making in Public Employment Services. In Companion Publication of the 2021 Conference on Computer Supported Cooperative Work and Social Computing. ACM, Virtual Event USA, 253–255. https://doi.org/10.1145/3462204.3481787
- Freiesleben and König (2023) Timo Freiesleben and Gunnar König. 2023. Dear XAI Community, We Need to Talk!. In Explainable Artificial Intelligence, Luca Longo (Ed.). Springer Nature Switzerland, Cham, 48–65. https://doi.org/10.1007/978-3-031-44064-9_3
- Grimm (2019) Stephen R. Grimm. 2019. Varieties of Understanding. In Varieties of Understanding. Oxford University Press, 1–14. https://doi.org/10.1093/oso/9780190860974.003.0001
- Guesmi et al. (2023) Mouadh Guesmi, Mohamed Amine Chatti, Shoeb Joarder, Qurat Ul Ain, Rawaa Alatrash, Clara Siepmann, and Tannaz Vahidi. 2023. Interactive Explanation with Varying Level of Details in an Explainable Scientific Literature Recommender System. International Journal of Human–Computer Interaction (Oct. 2023), 1–22. https://doi.org/10.1080/10447318.2023.2262797
- Hennink et al. (2017) Monique M. Hennink, Bonnie N. Kaiser, and Vincent C. Marconi. 2017. Code Saturation Versus Meaning Saturation: How Many Interviews Are Enough? Qualitative Health Research 27, 4 (March 2017), 591–608. https://doi.org/10.1177/1049732316665344
- Hoffman et al. (2023) Robert R. Hoffman, Shane T. Mueller, Gary Klein, Mohammadreza Jalaeian, and Connor Tate. 2023. Explainable AI: roles and stakeholders, desirements and challenges. Frontiers in Computer Science 5 (Aug. 2023), 1117848. https://doi.org/10.3389/fcomp.2023.1117848
- Janis (1972) Irving L Janis. 1972. Victims of groupthink: A psychological study of foreign-policy decisions and fiascoes. , viii, 277–viii, 277 pages.
- Johnson and Johnson (1985) David W. Johnson and Roger T. Johnson. 1985. The Internal Dynamics of Cooperative Learning Groups. Springer US, Boston, MA, 103–124. https://doi.org/10.1007/978-1-4899-3650-9_4
- Kawakami et al. (2024) Anna Kawakami, Amanda Coston, Haiyi Zhu, Hoda Heidari, and Kenneth Holstein. 2024. The Situate AI Guidebook: Co-Designing a Toolkit to Support Multi-Stakeholder, Early-stage Deliberations Around Public Sector AI Proposals. In Proceedings of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 749, 22 pages. https://doi.org/10.1145/3613904.3642849
- Keil (2019) Frank Keil. 2019. How Do Partial Understandings Work? In Varieties of Understanding. Oxford University Press, 191–208. https://doi.org/10.1093/oso/9780190860974.003.0010
- Keil (2003) Frank C. Keil. 2003. Folkscience: Coarse interpretations of a complex reality. Trends in Cognitive Sciences 7, 8 (Aug. 2003), 368–373. https://doi.org/10.1016/S1364-6613(03)00158-X
- Keil (2006) Frank C. Keil. 2006. Explanation and Understanding. Annual Review of Psychology 57, 1 (2006), 227–254. https://doi.org/10.1146/annurev.psych.57.102904.190100
- Kramer et al. (2018) Max F. Kramer, Jana Schaich Borg, Vincent Conitzer, and Walter Sinnott-Armstrong. 2018. When Do People Want AI to Make Decisions?. In Proceedings of the 2018 AAAI/ACM Conference on AI, Ethics, and Society (New Orleans, LA, USA) (AIES ’18). Association for Computing Machinery, New York, NY, USA, 204–209. https://doi.org/10.1145/3278721.3278752
- Krueger (2004) Richard A Krueger. 2004. Focus groups : a practical guide for applied research (3. ed., 6. print. ed.). Sage, Thousand Oaks, Calif. [u.a.].
- Kulesza et al. (2015) Todd Kulesza, Margaret Burnett, Weng-Keen Wong, and Simone Stumpf. 2015. Principles of Explanatory Debugging to Personalize Interactive Machine Learning. In Proceedings of the 20th International Conference on Intelligent User Interfaces. ACM, Atlanta Georgia USA, 126–137. https://doi.org/10.1145/2678025.2701399
- Kulesza et al. (2013) Todd Kulesza, Simone Stumpf, Margaret Burnett, Sherry Yang, Irwin Kwan, and Weng-Keen Wong. 2013. Too much, too little, or just right? Ways explanations impact end users’ mental models. In 2013 IEEE Symposium on Visual Languages and Human Centric Computing. IEEE, San Jose, CA, USA, 3–10. https://doi.org/10.1109/VLHCC.2013.6645235
- Langer et al. (2021) Markus Langer, Daniel Oster, Timo Speith, Holger Hermanns, Lena Kästner, Eva Schmidt, Andreas Sesing, and Kevin Baum. 2021. What do we want from Explainable Artificial Intelligence (XAI)? – A stakeholder perspective on XAI and a conceptual model guiding interdisciplinary XAI research. Artificial Intelligence 296 (July 2021), 103473. https://doi.org/10.1016/j.artint.2021.103473
- Lee et al. (2019) Min Kyung Lee, Daniel Kusbit, Anson Kahng, Ji Tae Kim, Xinran Yuan, Allissa Chan, Daniel See, Ritesh Noothigattu, Siheon Lee, Alexandros Psomas, and Ariel D. Procaccia. 2019. WeBuildAI: Participatory Framework for Algorithmic Governance. Proceedings of the ACM on Human-Computer Interaction 3, CSCW (Nov. 2019), 1–35. https://doi.org/10.1145/3359283
- Liao et al. (2020) Q. Vera Liao, Daniel Gruen, and Sarah Miller. 2020. Questioning the AI: Informing Design Practices for Explainable AI User Experiences. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–15. https://doi.org/10.1145/3313831.3376590
- Lim and Dey (2009) Brian Y. Lim and Anind K. Dey. 2009. Assessing Demand for Intelligibility in Context-Aware Applications. In Proceedings of the 11th International Conference on Ubiquitous Computing (Orlando, Florida, USA) (UbiComp ’09). Association for Computing Machinery, New York, NY, USA, 195–204. https://doi.org/10.1145/1620545.1620576
- Lima et al. (2023) Gabriel Lima, Nina Grgic-Hlaca, Jin Keun Jeong, and Meeyoung Cha. 2023. Who Should Pay When Machines Cause Harm? Laypeople’s Expectations of Legal Damages for Machine-Caused Harm. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency (Chicago, IL, USA) (FAccT ’23). Association for Computing Machinery, New York, NY, USA, 236–246. https://doi.org/10.1145/3593013.3593992
- Long et al. (2021) Duri Long, Aadarsh Padiyath, Anthony Teachey, and Brian Magerko. 2021. The Role of Collaboration, Creativity, and Embodiment in AI Learning Experiences. In Creativity and Cognition. ACM, Virtual Event Italy, 1–10. https://doi.org/10.1145/3450741.3465264
- Lopez (2019) Paola Lopez. 2019. Reinforcing Intersectional Inequality via the AMS Algorithm in Austria. In Proceedings of the 18th Annual STS Conference. Graz, 289–309. https://doi.org/10.3217/978-3-85125-668-0-16
- Lundberg and Lee (2017) Scott M. Lundberg and Su-In Lee. 2017. A Unified Approach to Interpreting Model Predictions. In Proceedings of the 31st International Conference on Neural Information Processing Systems (Long Beach, California, USA) (NIPS’17). Curran Associates Inc., Red Hook, NY, USA, 4768–4777.
- Maxwell (2010) Joseph A. Maxwell. 2010. Using Numbers in Qualitative Research. Qualitative Inquiry 16, 6 (2010), 475–482. https://doi.org/10.1177/1077800410364740
- Mercier and Sperber (2011) Hugo Mercier and Dan Sperber. 2011. Why do humans reason? Arguments for an argumentative theory. Behavioral and Brain Sciences 34, 2 (April 2011), 57–74. https://doi.org/10.1017/S0140525X10000968
- Miller (2019) Tim Miller. 2019. Explanation in artificial intelligence: Insights from the social sciences. Artificial Intelligence 267 (Feb. 2019), 1–38. https://doi.org/10.1016/j.artint.2018.07.007
- Molnar (2022) Christoph Molnar. 2022. Interpretable Machine Learning (2 ed.). https://christophm.github.io/interpretable-ml-book
- Moshman and Geil (1998) David Moshman and Molly Geil. 1998. Collaborative Reasoning: Evidence for Collective Rationality. Thinking & Reasoning 4, 3 (July 1998), 231–248. https://doi.org/10.1080/135467898394148
- Naiseh et al. (2021) Mohammad Naiseh, Dena Al-Thani, Nan Jiang, and Raian Ali. 2021. Explainable recommendation: when design meets trust calibration. World Wide Web 24, 5 (Sept. 2021), 1857–1884. https://doi.org/10.1007/s11280-021-00916-0
- Naiseh et al. (2020) Mohammad Naiseh, Nan Jiang, Jianbing Ma, and Raian Ali. 2020. Personalising Explainable Recommendations: Literature and Conceptualisation. In Trends and Innovations in Information Systems and Technologies, Álvaro Rocha, Hojjat Adeli, Luís Paulo Reis, Sandra Costanzo, Irena Orovic, and Fernando Moreira (Eds.). Vol. 1160. Springer International Publishing, Cham, 518–533. https://doi.org/10.1007/978-3-030-45691-7_49 Series Title: Advances in Intelligent Systems and Computing.
- Naiseh et al. (2024) Mohammad Naiseh, Catherine Webb, Tim Underwood, Gopal Ramchurn, Zoe Walters, Navamayooran Thavanesan, and Ganesh Vigneswaran. 2024. XAI for group-AI interaction: towards collaborative and inclusive explanation. In World conference for explainable artificial intelligence (17/07/24 - 19/07/24). https://eprints.soton.ac.uk/493227/
- Navajas et al. (2018) Joaquin Navajas, Tamara Niella, Gerry Garbulsky, Bahador Bahrami, and Mariano Sigman. 2018. Aggregated knowledge from a small number of debates outperforms the wisdom of large crowds. Nature Human Behaviour 2, 2 (Jan. 2018), 126–132. https://doi.org/10.1038/s41562-017-0273-4
- Niculae and Danescu-Niculescu-Mizil (2016) Vlad Niculae and Cristian Danescu-Niculescu-Mizil. 2016. Conversational Markers of Constructive Discussions. In Proceedings of the 2016 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies. Association for Computational Linguistics, San Diego, California, 568–578. https://doi.org/10.18653/v1/N16-1070
- Niklas et al. (2015) Jedrzej Niklas, Karolina Sztandar-Sztanderska, and Katarzyna Szymielewicz. 2015. Profiling the Unemployed in Poland: Social and Political Implications of Algorithmic Decision Making. panoptykon.org/sites/default/files/leadimage-biblioteka/panoptykon_profiling_report_final.pdf.
- Nokes-Malach et al. (2015) Timothy J. Nokes-Malach, J. Elizabeth Richey, and Soniya Gadgil. 2015. When Is It Better to Learn Together? Insights from Research on Collaborative Learning. Educational Psychology Review 27, 4 (Dec. 2015), 645–656. https://doi.org/10.1007/s10648-015-9312-8
- Patton (1990) Michael Quinn Patton. 1990. Qualitative evaluation and research methods, 2nd ed. Sage Publications, Inc, Thousand Oaks, CA, US. 532–532 pages.
- Raji et al. (2022) Inioluwa Deborah Raji, I. Elizabeth Kumar, Aaron Horowitz, and Andrew Selbst. 2022. The Fallacy of AI Functionality. In 2022 ACM Conference on Fairness, Accountability, and Transparency. ACM, Seoul Republic of Korea, 959–972. https://doi.org/10.1145/3531146.3533158
- Ribeiro et al. (2016) Marco Tulio Ribeiro, Sameer Singh, and Carlos Guestrin. 2016. ”Why Should I Trust You?”: Explaining the Predictions of Any Classifier. In Proceedings of the 22nd ACM SIGKDD International Conference on Knowledge Discovery and Data Mining (San Francisco, California, USA) (KDD ’16). Association for Computing Machinery, New York, NY, USA, 1135–1144. https://doi.org/10.1145/2939672.2939778
- Rozenblit and Keil (2002) Leonid Rozenblit and Frank Keil. 2002. The misunderstood limits of folk science: An illusion of explanatory depth. Cognitive Science 26, 5 (2002), 521–562. https://doi.org/10.1207/s15516709cog2605_1
- Sato et al. (2019) Brian K. Sato, Cynthia F. C. Hill, and Stanley M. Lo. 2019. Testing the test: Are exams measuring understanding? Biochemistry and Molecular Biology Education 47, 3 (May 2019), 296–302. https://doi.org/10.1002/bmb.21231
- Schmude et al. (2023) Timothée Schmude, Laura Koesten, Torsten Möller, and Sebastian Tschiatschek. 2023. On the Impact of Explanations on Understanding of Algorithmic Decision-Making. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency (Chicago, IL, USA) (FAccT ’23). Association for Computing Machinery, New York, NY, USA, 959–970. https://doi.org/10.1145/3593013.3594054
- Schmude et al. (2024) Timothée Schmude, Laura Koesten, Torsten Möller, and Sebastian Tschiatschek. 2024. Information That Matters: Exploring Information Needs of People Affected by Algorithmic Decisions. arXiv:2401.13324 [cs.HC]
- Scott et al. (2022) Kristen M. Scott, Sonja Mei Wang, Milagros Miceli, Pieter Delobelle, Karolina Sztandar-Sztanderska, and Bettina Berendt. 2022. Algorithmic Tools in Public Employment Services: Towards a Jobseeker-Centric Perspective. In Proceedings of the 2022 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’22). ACM, Seoul Republic of Korea, 2138–2148. https://doi.org/10.1145/3531146.3534631
- Shen et al. (2020) Hong Shen, Haojian Jin, Angel Alexander Cabrera, Adam Perer, Haiyi Zhu, and Jason I. Hong. 2020. Designing Alternative Representations of Confusion Matrices to Support Non-Expert Public Understanding of Algorithm Performance. Proceedings of the ACM on Human-Computer Interaction 4, CSCW2 (Oct. 2020), 1–22. https://doi.org/10.1145/3415224
- Shin (2023) Don Donghee Shin. 2023. Algorithms, humans, and interactions: How do algorithms interact with people? designing meaningful AI experiences (first edition ed.). Routledge, Boca Raton, FL. https://doi.org/10.1201/b23083
- Shneiderman (2022) Ben Shneiderman. 2022. Human-centered AI. Oxford University Press, Oxford.
- Shulner-Tal et al. (2022) Avital Shulner-Tal, Tsvi Kuflik, and Doron Kliger. 2022. Enhancing Fairness Perception – Towards Human-Centred AI and Personalized Explanations Understanding the Factors Influencing Laypeople’s Fairness Perceptions of Algorithmic Decisions. International Journal of Human–Computer Interaction (July 2022), 1–28. https://doi.org/10.1080/10447318.2022.2095705
- Smith et al. (2009) M. K. Smith, W. B. Wood, W. K. Adams, C. Wieman, J. K. Knight, N. Guild, and T. T. Su. 2009. Why Peer Discussion Improves Student Performance on In-Class Concept Questions. Science 323, 5910 (Jan. 2009), 122–124. https://doi.org/10.1126/science.1165919
- Speith et al. (2024) Timo Speith, Barnaby Crook, Sara Mann, Astrid Schomäcker, and Markus Langer. 2024. Conceptualizing understanding in explainable artificial intelligence (XAI): an abilities-based approach. Ethics and Information Technology 26, 2 (June 2024), 40. https://doi.org/10.1007/s10676-024-09769-3
- Sutcliffe (2005) Alistair Sutcliffe. 2005. Applying small group theory to analysis and design of CSCW systems. In Proceedings of the 2005 workshop on Human and social factors of software engineering - HSSE ’05. ACM Press, St. Louis, Missouri, 1–6. https://doi.org/10.1145/1083106.1083119
- Szymanski et al. (2021) Maxwell Szymanski, Martijn Millecamp, and Katrien Verbert. 2021. Visual, textual or hybrid: the effect of user expertise on different explanations. In 26th International Conference on Intelligent User Interfaces. ACM, College Station TX USA, 109–119. https://doi.org/10.1145/3397481.3450662
- Thiebes et al. (2021) Scott Thiebes, Sebastian Lins, and Ali Sunyaev. 2021. Trustworthy Artificial Intelligence. Electronic Markets 31, 2 (June 2021), 447–464. https://doi.org/10.1007/s12525-020-00441-4
- Wang and Yin (2021) Xinru Wang and Ming Yin. 2021. Are Explanations Helpful? A Comparative Study of the Effects of Explanations in AI-Assisted Decision-Making. In 26th International Conference on Intelligent User Interfaces (IUI ’21). Association for Computing Machinery, New York, NY, USA, 318–328. https://doi.org/10.1145/3397481.3450650
- Wason (1968) P. C. Wason. 1968. Reasoning about a Rule. Quarterly Journal of Experimental Psychology 20, 3 (1968), 273–281. https://doi.org/10.1080/14640746808400161
- Weiss (1995) Robert Stuart Weiss. 1995. Learning from strangers: The art and method of qualitative interview studies (1. free press paperback ed.). Free Press, New York, NY.
- Weitz et al. (2024) Katharina Weitz, Ruben Schlagowski, Elisabeth André, Maris Männiste, and Ceenu George. 2024. Explaining It Your Way - Findings from a Co-Creative Design Workshop on Designing XAI Applications with AI End-Users from the Public Sector. In Proceedings of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 745, 14 pages. https://doi.org/10.1145/3613904.3642563
- Wenzelburger et al. (2022) Georg Wenzelburger, Pascal D. König, Julia Felfeli, and Anja Achtziger. 2022. Algorithms in the public sector. Why context matters. Public Administration (Dec. 2022), padm.12901. https://doi.org/10.1111/padm.12901
- Wieringa (2023) Maranke Wieringa. 2023. “Hey SyRI, tell me about algorithmic accountability”: Lessons from a landmark case. Data & Policy 5 (2023), e2. https://doi.org/10.1017/dap.2022.39
- Wiggins and McTighe (2005) Grant P. Wiggins and Jay McTighe. 2005. Understanding by design (expanded 2nd ed.). Association for Supervision and Curriculum Development, Alexandria, VA.
- Xu (2019) Wei Xu. 2019. Toward human-centered AI: A perspective from human-computer interaction. Interactions 26, 4 (jun 2019), 42–46. https://doi.org/10.1145/3328485
- Yurrita et al. (2022) Mireia Yurrita, Dave Murray-Rust, Agathe Balayn, and Alessandro Bozzon. 2022. Towards a multi-stakeholder value-based assessment framework for algorithmic systems. In Proceedings of the 2022 ACM Conference on Fairness, Accountability, and Transparency (Seoul, Republic of Korea) (FAccT ’22). Association for Computing Machinery, New York, NY, USA, 535–563. https://doi.org/10.1145/3531146.3533118
- Zagzebski (2019) Linda Zagzebski. 2019. Toward a Theory of Understanding. In Varieties of Understanding. Oxford University Press, 123–136. https://doi.org/10.1093/oso/9780190860974.003.0007
- Zhang et al. (2023) Angie Zhang, Olympia Walker, Kaci Nguyen, Jiajun Dai, Anqing Chen, and Min Kyung Lee. 2023. Deliberating with AI: Improving Decision-Making for the Future through Participatory AI Design and Stakeholder Deliberation. Proceedings of the ACM on Human-Computer Interaction 7, CSCW1 (April 2023), 1–32. https://doi.org/10.1145/3579601
- Züger and Asghari (2023) Theresa Züger and Hadi Asghari. 2023. AI for the public. How public interest theory shifts the discourse on AI. AI & SOCIETY 38, 2 (April 2023), 815–828. https://doi.org/10.1007/s00146-022-01480-5
Appendix
The supplementary material provides additional information on the study and use case. In Section A, we provide the features used by the employment prediction algorithm, the mock newspaper article introducing the use case to participants, and the study task description. Section B lists the self-report and interview questions that were asked during the first and second self-reports.
Appendix A Additional information on the employment prediction algorithm
In this section, we give additional information on the features used by the employment prediction algorithm, the mock newspaper article introducing the system to participants, and the study task participants solved.999More detailed information on the deployment context of the system has been omitted to adhere to the anonymization policy but will be re-inserted for the final version.
A.1. Features used by the employment prediction algorithm
| Variable | Nominal values |
| Gender | Male/Female |
| Age group | 0–29/30–49/50+ |
| Citizenship | [Deployment country]/EU/Non-EU |
| Highest level of education | Grade school/apprenticeship, vocational school/high- or secondary school, university |
| Health impairment | Yes/No |
| Obligations of care (only women) | Yes/No |
| Occupational group | Production sector/service sector |
| Regional labor market | Five categories for employment prospects in assigned job center |
| Prior occupational career | [Variables as described below] |
| Days of gainful employment within 4 years | 75%/75% |
| Cases within four 1-year intervals | 0 cases/1 case/min. 1 case in 2 intervals/min. 1 case in 3 or 4 intervals |
| Cases with a duration longer than 180 days | 0 cases/min. 1 case |
| Measures claimed | 0/min. 1 supportive/min. 1 educational/min. 1 subsidized employment |
A.2. Mock newspaper article
A.3. Task description
Appendix B Self-reports and interview guide
B.1. Self-reports
Participants gave self-reports twice in the study, before and after the explanation phase (described in Section 3). In the following, we list each self-report question and the available answers.
Understanding I + II
“I think that I understand the system…”
(1 = very little; 2 = little; 3 = neither/nor; 4 = well; 5 = very well) -
Individual decision I + II
“In your opinion, should the system be introduced?” (Yes / No) -
Decision confidence I + II
“In making this decision, I am…”
(1 = very uncertain; 2 = uncertain; 3 = neither/nor; 4 = certain; 5 = very certain) -
Explanation helpfulness
”Which explanations did you find most helpful for your understanding?”
(choose any from: data, system details, usage, context) -
Explanation influence on decision
”Which explanations were most influential to your decision?”
(choose any from: data, system details, usage, context) -
Contributing your voice (focus groups only)
“I was able to contribute my voice in the group discussion…”
(1 = very little; 2 = little; 3 = neither/nor; 4 = well; 5 = very well) -
Influence of discussion (focus groups only)
“The group discussion influenced my decision…”
(1 = very little; 2 = little; 3 = neither/nor; 4 = strongly; 5 = very strongly)
B.2. Interview guide
During the second self-report of participants, the investigator asked interview questions about participants’ interaction with the explanations, their understanding processes, the most relevant information, and any additional situational questions. In the following, we list the questions composing the interview guide. The questions about inclusion and voice in the group were omitted in single interviews.
Understanding II
How did the explanations help you to understand the system?
What did you find difficult to understand?
And how did the collaboration help you?
Was something missing? An explanation or a question? -
Individual decision II
How do you feel about this decision?
Decision confidence II
How have the explanations and the collaboration influenced your decision confidence?
Explanation helpfulness
Which of the explanations made you realize: Ah, I’ve understood something, that’s good to know. And why? What effect did that have?
How did you communicate this to the group?
Explanation influence on discussion
Which explanation made you think: Oh, that’s important. It changes how I think about it. And why?
Contributing your voice
How did you feel about the discussion process? Was everyone able to say everything?
Influence of discussion
How do you feel about the decision the group made?

### v2

@@
-Deliberative XAI: How Explanations Impact Understanding and Decision-Making of AI Novices in Collective and Individual Settings
+Better Together? The Role of Explanations in Supporting Novices in Individual and Collective Deliberations about AI
 Abstract.
-XAI research often focuses on settings where people learn about and assess algorithmic systems individually.
-However, as more public AI systems are deployed, it becomes essential for XAI to facilitate collective understanding and deliberation.
-We conducted a task-based interview study involving 8 focus groups and 12 individual interviews to explore how explanations can support AI novices in understanding and forming opinions about AI systems.
-Participants received a collection of explanations organized into four information categories to solve tasks and decide about a system’s deployment.
-These explanations improved or calibrated participants’ self-reported understanding and decision confidence and facilitated group discussions.
-Participants valued both technical and contextual information and the self-directed and modular explanation structure.
-Our contributions include an explanation approach that facilitates both individual and collaborative interaction and explanation design recommendations, including active and controllable exploration, different levels of information detail and breadth, and adaptations to the needs of decision subjects.
+Deploying AI systems in public institutions can have far-reaching consequences for many people, making it a matter of public interest.
+Providing opportunities for stakeholders to come together, understand these systems, and debate their merits and harms is thus essential.
+Explainable AI often focuses on individuals, but deliberation benefits from group settings, which are underexplored.
+To address this gap, we present findings from an interview study with 8 focus groups and 12 individuals.
+Our findings provide insight into how explanations support AI novices in deliberating alone and in groups.
+Participants used modular explanations with four information categories to solve tasks and decide about an AI system’s deployment.
+We found that the explanations supported groups in creating shared understanding and in finding arguments for and against the system’s deployment.
+In comparison, individual participants engaged with explanations in more depth and performed better in the study tasks, but missed an exchange with others.
+Based on our findings, we provide suggestions on how explanations should be designed to work in group settings and describe their potential use in real-world contexts.
+With this, our contributions inform XAI research that aims to enable AI novices to understand and deliberate AI systems in the public sector.
 1. Introduction
-A growing number of algorithmic decision-making systems (ADM systems) are deployed and used in the public sector to decide about critical issues, such as employment, migration, and criminal justice (Scott et al., 2022; Bansak et al., 2018; Züger and Asghari, 2023; Chouldechova, 2017).
-These systems can result in discrimination and damages for decision subjects (Brown et al., 2019; Raji et al., 2022), especially when decision-making is opaque and uncontestable (Ananny and Crawford, 2018; Alfrink et al., 2023; de Fine Licht and de Fine Licht, 2020).
-The field of explainable AI (XAI) aims to reduce these risks by developing methods to make ADM systems more understandable and facilitate their assessment and control (Langer et al., 2021).
-But while much of XAI research is focused on individuals, research has shown that learning about and understanding complex topics benefits from collaboration in group settings (Smith et al., 2009; Moshman and Geil, 1998; Nokes-Malach et al., 2015; Navajas et al., 2018).
-Work on public AI further argues that multi-stakeholder deliberation is an essential step in ensuring public participation in deployment decisions of these systems (Kawakami et al., 2024; Naiseh et al., 2024; Brown et al., 2019), but how explanations can fulfill these demands is an open question in XAI.
-Crucially, many people affected by decisions of ADM systems have no technical background in developing or analyzing these systems.
-That is, a majority of decision subjects, domain experts, and other stakeholders of the public are lay people (Lima et al., 2023; DeVito et al., 2018; Shulner-Tal et al., 2022) or AI novices111We use the term in reference to ”visualization novices” (Burns et al., 2023) and synonymously with ”lay people” (Szymanski et al., 2021; Shen et al., 2020; DeVito et al., 2018)..
-Explanations for AI novices naturally need to fulfill different requirements than for AI practitioners, as they have to be adapted to their background (Ehsan et al., 2024), interests (Langer et al., 2021), and prior knowledge (Szymanski et al., 2021; Cheng et al., 2019; Schmude et al., 2024), which is why personalization (Shulner-Tal et al., 2022; Conati et al., 2021; Naiseh et al., 2020) and interactivity (Cheng et al., 2019; Bertrand et al., 2023; Guesmi et al., 2023) are central concepts in explanation design.
-Further, while explanation formats (e.g., interactive, visual, textual) are known to impact AI novices’ understanding of algorithmic systems (Cheng et al., 2019; Bove et al., 2022; Szymanski et al., 2021), the precise effect is inconsistent between studies (Bove et al., 2022) due to contextual factors such as participants’ perceptions of the use case domain.
-Previous qualitative research identified the information needs of AI novices (Schmude et al., 2024), emphasized user-preference for social and conversational formats (Bertrand et al., 2023), and employed participatory methods to design explanations in collaboration with end-users (Weitz et al., 2024; Lee et al., 2019).
-This work connects to these insights and draws from disciplines such as CSCW (Convertino et al., 2008; Fiesler et al., 2019), cognitive science (Keil, 2006, 2003), and educational psychology (Smith et al., 2009; Nokes-Malach et al., 2015) to examine how explanations can support understanding and decision-making of AI novices in individual and collaborative settings (Naiseh et al., 2024; Weitz et al., 2024).
-Developing this design approach to improve understanding and facilitate deliberation meets multiple challenges.
-Group composition and dynamics place special demands on explanation design (Naiseh et al., 2021), as explanations must cover a diverse set of information needs (Schmude et al., 2024; Bove et al., 2024) and format preferences (Bertrand et al., 2023).
-In addition, they must support collaborative interaction (Long et al., 2021), such as sharing and simultaneous processing of information.
-That is, explanations need to provide comprehensive information but remain clear and navigable, and should foster group interaction but still be suitable for individual settings.
-We address these challenges by proposing a modular explanation design that spans four information categories from which users can select information according to their needs.
-Additionally, when validating explanation approaches qualitatively, it is crucial to contact the relevant stakeholder groups, which can be especially challenging with decision subjects in minority groups or difficult situations, who arguably have the highest risk of incurring negative consequences from ADM.
-To address this, we conducted two focus groups with decision subjects to include their perspectives and voices in developing explanations for AI systems in the public sector (Brown et al., 2019).
-This paper explores how explanations can be used in individual and collaborative settings to support AI novices’ learning, interaction, and decision-making regarding the deployment of ADM systems.
-We present the findings of a task-based interview study with 43 AI novices involving 8 focus groups and 12 single interviews.
-We propose an explanation design that addresses AI novices’ information needs and applies current best practices in XAI (Byrne, 2023).
-The design consists of a collection of 36 explanations in the form of question-answer pairs organized into the four categories data, system details, usage, and context, which are further subdivided into subtopics and levels of detail (Figure 1).
-Participants used these explanations to solve the study tasks and decide about deploying an ADM system (Figure 3).
-The study’s use case is an algorithm for the prediction of job-seekers’ employability, connecting to previous work on ADM systems in employment (Scott et al., 2022; Niklas et al., 2015; Lopez, 2019).
-We analyze participants’ self-reported understanding, decision confidence, and perceptions of key information, and conduct a thematic analysis of group interactions with explanations.
-We discuss explanation design suggestions and the benefits and challenges of using explanations in collaborative settings.
-In the analysis of our findings, we put particular focus on the following research questions:
+A growing number of AI systems111We use the term ‘AI system’ to describe algorithmic systems with machine learning components.
+The nomenclature follows research on explainable AI (Langer et al., 2021) and research on AI in the context of society (Züger and Asghari, 2023; Collins et al., 2024) and regulation (Panigutti et al., 2023).
+are deployed in the public sector to decide about critical issues, such as employment, migration, and criminal justice (Scott et al., 2022; Bansak et al., 2018; Züger and Asghari, 2023; Chouldechova, 2017).
+These systems can have consequences for all stakeholders but tend to have the largest impact on their decision subjects (people the system decides over), such as discrimination or misclassification (Brown et al., 2019; Raji et al., 2022).
+These harms intensify when decision-making is opaque and uncontestable (Ananny and Crawford, 2018; Alfrink et al., 2023; de Fine Licht and de Fine Licht, 2020).
+For these reasons, public AI systems should be considered as ‘matters of public interest’ (Züger and Asghari, 2023), meaning that they need to be explainable, justifiable, and open for public deliberation (Kawakami et al., 2024; Naiseh et al., 2024; Brown et al., 2019).
+Explanations can make AI systems more understandable and easier to assess and control (Langer et al., 2021).
+Much of explainable AI (XAI) research is focused on individuals, but research has shown that group settings can facilitate the understanding of complex topics (Moshman and Geil, 1998; Nokes-Malach et al., 2015; Navajas et al., 2018).
+Further, group settings encourage the exchange of views and arguments (Stromer-Galley, 2007; Smith et al., 2009), which are vital when engaging in deliberation (collectively finding a solution to a problem (Habermas, 1991)).
+XAI has not explored in detail how explanations can be combined with group settings to leverage these benefits.
+In this paper, we aim to address this gap.
+As arguably many people affected by decisions of AI systems have no technical background in developing or analyzing these systems, this work focuses on ‘lay people’ (Lima et al., 2023; DeVito et al., 2018; Shulner-Tal et al., 2022) or ‘AI novices’ (Mohseni et al., 2020).
+Explanations for AI novices naturally have different requirements than explanations for AI practitioners, as they have different expertise (Ehsan et al., 2024), interests (Langer et al., 2021), and prior knowledge (Szymanski et al., 2021; Cheng et al., 2019; Schmude et al., 2024).
+Explanation formats (e.g., visual, textual, dialogue) are known to impact AI novices’ understanding (Cheng et al., 2019; Bove et al., 2022; Szymanski et al., 2021) but show inconsistent effects (Bove et al., 2022) due to contextual factors such as participants’ perceptions of the use case domain.
+A possible solution is the adaptation of explanations with personalization (Shulner-Tal et al., 2022; Conati et al., 2021; Naiseh et al., 2020) and interactivity (Cheng et al., 2019; Bertrand et al., 2023; Guesmi et al., 2023).
+To this end, previous research has analyzed AI novices’ information needs (Schmude et al., 2024) and developed explanations in collaboration with end-users (Weitz et al., 2024; Lee et al., 2019b).
+However, designs that can adapt to AI novices and that support their deliberation, especially in groups, are still rare (Chiang et al., 2024; Naiseh et al., 2024).
+But these explanations are essential to provide opportunities where AI novices can learn about and discuss AI systems and to realize the principles of human-centered AI (engaging stakeholders and empowering people (Shneiderman, 2022)).
+Designing explanations that support understanding and deliberation for AI novices in both group and individual settings meets multiple challenges.
+Group composition and dynamics place special demands on explanation design (Naiseh et al., 2021), as explanations need to cater to a diverse set of information and format preferences (Schmude et al., 2024; Bove et al., 2024; Bertrand et al., 2023).
+They must further support a joint understanding process and collaborative interactions (Long et al., 2021), such as sharing and combining, all while providing comprehensive information and remaining clear and navigable.
+We address these challenges by proposing a modular explanation design that spans four information categories (data, system details, usage, and context) from which users can select.
+Another challenge consists in validating explanation approaches qualitatively with the relevant stakeholder groups.
+Specifically, XAI research does not always include people from marginalized population groups, who are most likely to be affected negatively as decision subjects (Brown et al., 2019).
+To address this, we conducted two focus groups with decision subjects to include their perspectives and voices on AI systems in the public sector.
+To examine the role of explanations in supporting AI novices’ understanding and deliberation we present the findings of a task-based interview study with 43 participants, involving 8 focus groups and 12 single interviews.
+For this study, we used an explanation design comprising 36 single explanations in question-answer pairs.
+These explanations are organized into the four categories data, system details, usage, and context and further assigned to subtopics and levels of detail (Figure 1).
+Participants used these explanations to solve the study tasks and decide about deploying a public AI system (Figure 3).
+We used an employment scoring algorithm that connects to previous work on AI systems in employment (Scott et al., 2022; Niklas et al., 2015; Lopez, 2019).
+Our analysis examines participants’ self-reported understanding, decision confidence, and perceptions of key information.
+We further conducted a thematic analysis of how participants interacted with explanations in both settings.
+The following research questions guide the analysis:
 [RQ1]
-Explanations: How does a question-driven, modular explanation design support AI novices’ understanding in individual and group settings?
+Explanations: How does a question-driven, modular explanation design support AI novices’ understanding in groups and individual settings?
 [RQ2]
-Deliberation: How do the explanations support group interaction with information and deliberation processes?
-The contributions of our work include i) the validation of an explanation design building on current XAI principles, including question-driven and modular design, different levels of completeness and soundness, and suitability for both individual and collective settings; ii) an analysis of the kind and amount of information that participants requested most often and perceived as most important; iii) an analysis of how our explanations support group collaboration in solving tasks and group deliberation in forming opinions about ADM systems; and iv) recommendations for the design and use of these explanations and implications for the incorporation of group settings in XAI.
+Deliberation: How do AI novices use explanations to form opinions and make decisions about AI systems?
+Our contributions include i) an explanation design that builds on a question-driven and modular design to accommodate different levels of completeness and soundness and that is suitable for both individual and group settings; ii) an in-depth description of how explanations support participants’ understanding and deliberation processes that identifies salient differences between the two settings; iii) an analysis of which type of explanations participants requested most often and perceived as most important; and iv) recommendations regarding the design and use of explanations in group settings.
+We envision that this work can provide valuable starting points for future XAI research that aims to connect explanations to deliberation on public AI systems.
 2. Background and Related Work
-In this section, we briefly describe how our work connects to human-centered AI, research on explanation design for AI novices, and literature on understanding, interaction, and decision-making in groups.
+This section describes how our work is embedded in human-centered explainable AI and outlines the main challenges and approaches to designing explanations for AI novices.
+It further introduces the two main lenses of analysis to answer our research questions: understanding and deliberation.
 2.1. Human-centered explainable AI
-Explainability is often described as a cornerstone of responsible AI systems222While not every ADM system includes AI, we use both terms as we focus on decision-making systems with machine learning components.
-(Thiebes et al., 2021), as explanations can enable people to understand (Langer et al., 2021) and contest AI decisions (Alfrink et al., 2023).
-Similar concepts are embedded in the notion of human-centered AI (Capel and Brereton, 2023), which proposes to 1) build systems based on user observation and stakeholder engagement and 2) empower rather than replace people by making these systems controllable and autonomy-preserving (Shneiderman, 2022; Shin, 2023; Xu, 2019).
-These measures become especially important in high-risk settings (European Commission, 2021), such as employment (Scott et al., 2022; Flügge, 2021), immigration (Bansak et al., 2018), or criminal justice (Chouldechova, 2017), where erroneous and non-transparent algorithmic decisions can cause severe harm to those affected (Raji et al., 2022).
-To better understand how explanations can contribute to these principles, human-centered explainable AI places emphasis on the sociotechnical dimensions of explaining AI systems (Ehsan et al., 2023b) in that it assumes that transparency alone is not enough to make AI systems explainable (Ananny and Crawford, 2018), but that explanations need to consider the system’s social context (Wenzelburger et al., 2022), its lifecycle (Dhanorkar et al., 2021) and its different stakeholder groups (Ehsan et al., 2023b).
-In the context of this work, human-centered explainability is realized by testing and validating a design approach that has been developed based on qualitative exploration of AI novices’ information needs (Schmude et al., 2024).
-Further, our work aims to examine suitable formats for explanations that can accompany the deployment of AI systems in public institutions (de Fine Licht and de Fine Licht, 2020) and inform approaches of practicing responsible AI in high-risk application settings (Ehsan et al., 2023b).
-2.2. Explanation design for AI novices
-Explanations for non-technical audiences have different requirements than those for AI practitioners.
-Explanation methods like LIME (Ribeiro et al., 2016), SHAP (Lundberg and Lee, 2017), and surrogate models (Molnar, 2022) are better suited to experts due to their reliance on previous knowledge and technical literacy, while effective methods for AI novices are still being developed.
-As transparency does not equal understanding (Ananny and Crawford, 2018), simply making all information about a system available is no valid explanation strategy.
-Instead, explanations must be selected and designed to be suitable to the recipient’s stakeholder role (Langer et al., 2021), prior knowledge (Schmude et al., 2024), beliefs (Miller, 2019), and explanatory stance (Byrne, 2023; Keil, 2006).
-Various studies examined how such explanations for AI novices might be realized: Cheng et al.
-(2019) showed that ”white-box” explanations (transparent models) can improve ”objective”333We use the term in reference to Cheng et al.
+Explainability is often described as a cornerstone of responsible AI systems (Thiebes et al., 2021), as explanations can enable stakeholders such as users and decision subjects to understand (Langer et al., 2021) and contest AI decisions (Alfrink et al., 2023).
+A similar focus is set by the domain of human-centered AI (Capel and Brereton, 2023), which proposes to build AI systems that 1) are based on user-experience design and stakeholder engagement, and 2) empower rather than replace people by being controllable and autonomy-preserving (Shneiderman, 2022; Shin, 2023; Xu, 2019).
+These principles become especially important in high-risk settings (European Commission, 2024), such as employment (Scott et al., 2022; Flügge, 2021), immigration (Bansak et al., 2018), or criminal justice (Chouldechova, 2017), where erroneous or non-transparent algorithmic decisions can cause severe harm to those affected (Raji et al., 2022).
+In response to these risks, the domain of human-centered explainable AI (HCXAI) examines how explanations can contribute to “equitable and ethical Human-AI interaction” (Ehsan et al., 2023b).
+It assumes that transparency alone is not enough to make AI systems explainable (Ananny and Crawford, 2018), but that explanations need to consider the system’s social context (Wenzelburger et al., 2022), its lifecycle (Dhanorkar et al., 2021) and its different stakeholder groups (Ehsan et al., 2023b).
+In the context of this work, human-centered explainability is realized by testing and validating a design approach intended to support AI novices in understanding AI systems and deciding about their deployment in public institutions (Züger and Asghari, 2023).
+2.2. Designing explanations for AI novices
+The majority of people who interact and are involved with AI systems in public institutions are lay people or AI novices, here defined as “users who [might] use AI products in daily life but have no (or very little) expertise on machine learning systems” (Mohseni et al., 2020).
+Established explanation methods, like LIME (Ribeiro et al., 2016), SHAP (Lundberg and Lee, 2017), and surrogate models (Molnar, 2022) are tailored to experts and require technical knowledge; hence, they do not address the needs of AI novices.
+To better cover these needs and match information to them, it is necessary to understand how non-experts conceive of AI systems.
+Previous HCI research has analyzed lay understandings to explore user perception and understanding of several algorithmic systems (DeVito et al., 2018; Eslami et al., 2016).
+Similarly, XAI research has begun to explore the information needs of AI novices to design suitable explanations for a broader audience (Schmude et al., 2024).
+However, few explanation designs have been proposed that truly assume the perspectives of AI novices (Szymanski et al., 2021; Cheng et al., 2019).
+In the following, we summarize current approaches with respect to AI novices’ information needs and current practices of explanation design.
+Regarding information needs, previous qualitative research outlined that AI novices value information about the context and intention of a system’s deployment (Schmude et al., 2024; Kawakami et al., 2024) as well as about the responsible institution (Brown et al., 2019).
+In contrast to traditional XAI approaches, which focus on descriptive information about the system’s workings and outputs, explanations for AI novices thus also require normative information, such as justifications (Biran and McKeown, 2017) for design choices.
+Regarding information coverage, previous work posits that transparency does not equal understanding (Ananny and Crawford, 2018) and that simply making all information about a system available is no valid explanation strategy.
+Empirical evaluations of this claim showed that “white-box” explanations (transparent models) can improve “objective”222We use the term in reference to Cheng et al.
 (2019) and Bove et al.
 (2022), it means to describe factual or testable understanding.
-understanding but may overwhelm users, reducing perceived understanding.
-However, Bove et al.
-(2022) used similar explanations and found that they had the opposite effect on participant understanding, which they attribute to a difference in the studies’ use case domains (student admission vs.
+understanding but may overwhelm non-expert users and reduce perceived understanding (Cheng et al., 2019).
+However, later work (Bove et al., 2024) used similar explanations and found that they had the opposite effect on understanding, attributed to a difference in the studies’ use case domains (student admission vs.
 finance).
-Szymanski et al.
+This indicates that the amount of information should be adaptable to the given context.
+Regarding explanation format, Szymanski et al.
 (2021) examined how expert and lay users rated explanations of different formats and found that while lay users favored visual explanations, they performed better with textual ones.
-Other studies underscore this variability, indicating that comprehension varies with personal and demographic factors and domain knowledge (Wang and Yin, 2021; Shulner-Tal et al., 2022; Ehsan et al., 2024) and that personalized explanations could be effective in addressing this (Conati et al., 2021; Shulner-Tal et al., 2022).
-Further aspects to be considered include the explanation’s purpose (Freiesleben and König, 2023) and the recipient’s familiarity with AI (Kramer et al., 2018).
-Regarding the information needs, previous qualitative research outlined that AI novices value information about the context and intention of a system’s deployment (Schmude et al., 2024; Kawakami et al., 2024) as well about the responsible institution (Brown et al., 2019).
-Notably, Lee et al.
-(2019) and Weitz et al.
-(2024) conducted participatory workshops to design explanations with end-users in the public sector, finding that participants valued the collaborative settings and their improved understanding of algorithmic systems.
-Drawing from these approaches, this work focuses on the design and qualitative validation of explanations such that they address the information needs of AI novices (Schmude et al., 2024) and support their understanding (Langer et al., 2021) and decision-making (Bertrand et al., 2023) in individual and collective settings.
-Our explanation design draws from various approaches in previous work.
-The information base used for the explanations was compiled from different sources documenting the development and planned deployment of employment prediction algorithms (Scott et al., 2022; Allhutter et al., 2020), producing an extensive collection of ”scavenged” (Wieringa, 2023) material.
-To structure this collection, we drew from Lim and Dey (2009)’s work on intelligibility types, Liao et al.
-(2020)’s concept of question-driven explanation design, and Schmude et al.
-(2024)’s separation of information categories into data, system details, usage, and context.
-We further apply Kulesza et al.
-(2013)’s principles on explanation soundness (fidelity, complexity) and completeness (coverage, density) and draw from works that apply these principles in explanation design (Chatti et al., 2022; Guesmi et al., 2023; Kulesza et al., 2015), by introducing a structure of sub-topics and a hierarchy of explanation levels.
-This combination of question-driven explanations, different levels of information detail and breadth, and user-controlled selection of information aims to introduce modularity and interactivity (Cheng et al., 2019; Schmude et al., 2023; Guesmi et al., 2023) while allowing for adjusting of explanations to users’ needs (Conati et al., 2021; Shulner-Tal et al., 2022).
-2.3. Understanding and decision-making in collaborative settings
-Improving understanding is a central goal of explanations (Langer et al., 2021) and enables people to pursue their goals, such as assessing or contesting a system’s decisions (Hoffman et al., 2023).
+Other studies confirm this discrepancy and posit that comprehension varies with demographic factors and domain knowledge (Wang and Yin, 2021; Shulner-Tal et al., 2022; Ehsan et al., 2024).
+These issues are assumed to be addressed with personalization of explanations (Conati et al., 2021; Shulner-Tal et al., 2022), meaning that they are selected and designed according to the user’s stakeholder role (Langer et al., 2021), prior knowledge (Schmude et al., 2024), beliefs (Miller, 2019), and explanatory stance (Byrne, 2023; Keil, 2006).
+Further aspects to be considered include the explanation’s purpose (Freiesleben and König, 2023) and the user’s familiarity with AI (Kramer et al., 2018).
+These approaches guided the conceptual development of the explanation design presented in this study.
+We compiled information from different sources documenting employment prediction algorithms (Scott et al., 2022; Allhutter et al., 2020), producing an extensive collection of “scavenged” (Wieringa, 2023) material.
+To structure this collection, we drew from work on intelligibility types (Lim and Dey, 2009), question-driven explanation design (Liao et al., 2020), and the separation of information categories into data, system details, usage, and context (Schmude et al., 2024).
+We further applied the principles of explanation soundness (fidelity, complexity) and completeness (coverage, density) (Chatti et al., 2022; Guesmi et al., 2023; Kulesza et al., 2015) by introducing a structure of sub-topics and a hierarchy of explanation levels.
+This combination of question-driven explanations, levels of detail, and user-controlled selection of information aims to support modularity and interactivity (Cheng et al., 2019; Schmude et al., 2023; Guesmi et al., 2023) as well as the adjustment of explanations to users’ needs (Conati et al., 2021; Shulner-Tal et al., 2022).
+Section 3.2 describes how these principles were realized in the explanation design.
+2.3. Analytical lenses: Understanding and deliberation
+In the following, we introduce understanding and deliberation to serve as the main analytical lenses for this paper.
+Section 3.5 then operationalizes them for the evaluation of the explanation design and settings.
+2.3.1. Individual and collaborative understanding of AI systems
+Improving understanding of an AI system is the primary purpose of explanations, as understanding is thought to enable assessment (e.g., of a system’s fairness) (Langer et al., 2021) and action (e.g., contestation) (Henin and Le Métayer, 2022) for the system’s stakeholders.
 However, understanding can be defined in numerous ways (Grimm, 2019; Baumberger et al., 2017; Zagzebski, 2019; Keil, 2006).
-This work draws from literature in the learning sciences, cognitive sciences, and XAI to define understanding as i) connecting and applying information, whereas ”knowledge” only stores information (Grimm, 2019; Baumberger et al., 2017), ii) being the attempt to grasp the underlying structure of a phenomenon by way of simplification (Zagzebski, 2019), iii) consisting of several ”facets” that include both the analytical and the emotional connection to information (Wiggins and McTighe, 2005), and iv) rarely being exhaustive, but rather a ”working” mental model that is attained by recognizing and filling gaps until the learner deems it sufficient (Keil, 2006).
-Due to the challenge of defining and measuring understanding (Sato et al., 2019), recent research has proposed an ”abilities-based” approach (Speith et al., 2024) which defines understanding through the skills the learner possesses, connecting to comparable operationalizations by the learning sciences (Wiggins and McTighe, 2005).
-While individual understanding has been the subject of many studies in XAI (Chen et al., 2023; Cheng et al., 2019; Wang and Yin, 2021; Schmude et al., 2023), understanding in group settings has been less explored.
-However, different disciplines have investigated collaborative understanding in detail: In the cognitive sciences, distributed cognition (Keil, 2003) and outsourced understanding (Keil, 2006) are components for the analysis of explanatory understanding, whereas educational psychology produced insight on the effects of group tutoring (Nokes-Malach et al., 2015; Smith et al., 2009; Moshman and Geil, 1998; Navajas et al., 2018) and comparisons to one-on-one tutoring (Bloom, 1984).
-Further, research in the field of computer-supported collaborative work provides approaches for the analysis of collaborative settings (Fiesler et al., 2019), group composition (Convertino et al., 2008), and group interactions (Sutcliffe, 2005).
-Recent research on public AI (Züger and Asghari, 2023) has emphasized the importance of including stakeholders by introducing deliberation early in the design process (Kawakami et al., 2024; Yurrita et al., 2022; Zhang et al., 2023).
-Further, recent work in XAI has begun to consider how explanations for group interactions could be approached, describing that ”many-to-one” interactions (multiple people interacting with an explanation) will likely differ enormously from ”one-to-one” interactions due to ”complexities in group dynamics, cognitive bias amplification, trust issues within the group, and group-centric evaluation” (Naiseh et al., 2024).
-Drawing from these conceptualizations, this work aims to explore empirically how explanations can support group deliberation by improving understanding and informing decision-making processes.
+This work draws from research in learning sciences, cognitive sciences, and explainable AI to define understanding as i) connecting and applying information (Grimm, 2019; Baumberger et al., 2017), ii) being the attempt to grasp the underlying structure of a phenomenon by way of simplification (Zagzebski, 2019), iii) consisting of several “facets” that include both the analytical and the emotional connection to information (explain, interpret, apply, take perspective, empathize, self-reflect) (Wiggins and McTighe, 2005), and iv) being a “working” mental model that is attained by recognizing and filling gaps until the learner deems it sufficient (Keil, 2006).
+Due to the challenge of defining and measuring understanding (Sato et al., 2019), recent research has proposed an “abilities-based” approach (Speith et al., 2024), connecting to comparable operationalizations by the learning sciences (Wiggins and McTighe, 2005).
+We examine understanding by analyzing which facets of understanding participants use to answer the study tasks and make a confident deployment decision (Section 3.4).
+While individual understanding has been the subject of many studies in XAI (Chen et al., 2023; Cheng et al., 2019; Wang and Yin, 2021; Schmude et al., 2023), understanding in group settings has been less explored (Chiang et al., 2023).
+We thus draw from disciplines that have investigated collaborative understanding: The cognitive sciences have examined distributed cognition (sharing cognitive load) and outsourcing (Keil, 2006) (delegating understanding) as fruitful mechanisms for collaborative settings, such as the navigation of a ship (Keil, 2003).
+The prerequisite is that groups achieve “cognitive symbioses with mutually supporting roles” (Keil, 2006), i.e., a constructive working dynamic.
+Similarly, educational psychology has found that peer discussion (Smith et al., 2009), collaborative reasoning (Moshman and Geil, 1998), and aggregated knowledge (Navajas et al., 2018) leads groups to perform better than individuals on the same tasks.
+However, whether groups perform well depends on their interactions, which can be described with cognitive and social mechanisms of collaborative success and failure (Nokes-Malach et al., 2015).
+When groups perform worse than individuals, the associated mechanisms include increased memory load and retrieval disruption (losing train of thought).
+In contrast, members tend to have established common ground and shared task-related information when they perform better.
+Thus, while it is not clear from the outset if groups are better for learning than one-on-one settings (Bloom, 1984), their advantages, such as sharing of cognitive load and exchange of views, likely support finding solutions to complex problems and present a valuable testing ground to deliberate deployment of AI systems.
+The field of computer-supported cooperative work (CSCW) has long examined collaborative settings about group composition and interactions (Fiesler et al., 2019; Convertino et al., 2008; Sutcliffe, 2005).
+However, work in XAI has only begun to consider how explanations for group interactions could be approached, describing that “many-to-one” interactions (multiple people interacting with an explanation) will likely differ from “one-to-one” interactions due to “complexities in group dynamics, cognitive bias amplification, trust issues within the group, and group-centric evaluation” (Naiseh et al., 2024).
+Lastly, previous work in XAI has examined individual versus group understanding in AI-assisted decision-making but surprisingly found little effect on understanding (Chiang et al., 2023).
+Following up on these findings, we use a ‘triangulation of methods’ (Carter et al., 2014), as described in Section 3.5), to empirically explore and compare the effects of explanations on the understanding processes of AI novices in groups and individual settings.
+2.3.2. Deliberating on AI systems
+Deliberation, in the sense of informed reasoning and decision-making, is based on understanding (de Fine Licht and de Fine Licht, 2020) and is key in enabling citizens to debate public sector AI proposals and their potential consequences (Kawakami et al., 2024; Züger and Asghari, 2023).
+Habermas (1991) describes deliberation as the exchange of rational-critical arguments on a problem to the end of finding a solution.
+These rational-critical arguments are grounded in truth or a shared understanding of reality, are open for judgment, and can be defended.
+This connection between shared understanding and deliberation is central to our examination of explanations’ effects.
+Deliberation takes place in many areas that shape politics and life in society (Lupia, 2023).
+Examples include public referendums that let inhabitants vote on jurisdictional changes (such as Swiss federal and state laws (Swiss Confederation, 2025)), citizen forums addressing matters of public importance (such as water supply in California (Innes and Booher, 2003)), and community-based grassroots formats where citizens support each other (such as the right to repair movement (Collins et al., 2024)).
+These settings have in common that they involve “social entities made up of people who are in one way or another engaged with their environment” and who use deliberation and productive conflict to negotiate and change policy issues (Hajer and Wagenaar, 2003).
+While these participation formats are not perfect and potentially incur cognitive biases such as groupthink333Prioritizing group harmony over real argumentation and discussion.
+(Naiseh et al., 2024; Janis, 1971; Baron, 2005), they create spaces where the general public can gather, discuss, form opinions, and decide on public interests.
+We argue that AI systems in public institutions constitute such public interests, encapsulated in the term public AI.
+Züger and Asghari (2023) employ the term to make explicit that AI systems in public institutions must fulfill obligations to prove their benefit.
+These obligations include being justifiable, equal, open to validation, technically secure, and the result of a deliberation or co-design process.
+Identifying formats supporting this deliberation on public AI systems is an open research challenge.
+Prior work has investigated how ’mini-publics’ (Fung, 2003) can be used to support the co-design of algorithmic policy (Lee et al., 2019b) and procedural justice in algorithmic resource allocation (Lee et al., 2019a).
+HCI research has further shown that participatory formats can connect communities and institutions in public service transformation (Crivellaro et al., 2019).
+And in XAI, studies showed that group discussions can facilitate the critical analysis of an AI system’s recommendations and that supplying information on both pros and cons of an AI’s recommendation lead to more frequent and more productive group deliberation (Chiang et al., 2023).
+However, settings that allow participants to deliberate in person on the deployment of high-stakes public AI are underexplored.
+We aim to address this gap by implementing mini-publics as focus groups with three different compositions (domain experts, decision subjects, and members of the general public444Participants who were neither directly affected as job-seekers nor were potential users of the system.), thus including stakeholders of different backgrounds and degrees of involvement.
+We further compare group deliberation processes with those in single interviews, which can be described as “internal deliberation” (Mercier and Landemore, 2012).
+On this basis, we aim to provide insight into suitable explanation designs and social formats to support deliberation on public AI systems.
+Section 3.5 describes the concrete analysis approach to this end.
 3. Methods
 In this section, we describe our methods and study procedure.
-We conducted a task-based semi-structured interview study with 43 participants recruited from civil society organizations, a job agency, and the author’s extended network (Section 3.5), structured into 8 focus groups (3–5 participants) and 12 single interviews.
-Participants were presented with the study’s employment prediction use case (Section 3.1) and received a collection of explanations about this system (Section 3.2) before solving four tasks and making decisions about the system’s deployment (Section 3.3).
+We conducted a task-based semi-structured interview study with 43 participants (Section 3.6), structured into 8 focus groups with 3–5 participants each and 12 single interviews (Section 3.3).
+Participants were presented with the study’s employment prediction use case (Section 3.1) and a collection of explanations about this system (Section 3.2) before solving four tasks and deciding about the system’s deployment (Section 3.4).
 The study closed with an interview, lasting 90–120 minutes for focus groups and 60 minutes for single interviews.
-We analyzed the explanations’ effect, participants’ prioritization of information, and the individual and collective interaction with the explanations (Section 3.4).
+We analyzed individual and collective interactions with the explanations, self-reports, and deliberation processes (Section 3.5).
 The university’s research ethics committee approved this study.
-3.1. Use case: The employment prediction algorithm
-The employment prediction algorithm444More detailed information on the use case has been omitted to adhere to the anonymization policy but will be re-inserted for the final version.
-is a system meant to support counselors of a public employment agency in assessing job-seekers’ employability.
-It uses a logistic regression model trained on historical data to assess how demographic features (e.g., age, education, nationality) influence employment chances, producing short-term and long-term employment scores as recommendations.
-Counselors use these recommendations to decide about adequate support measures for the job-seeker (courses, stabilizing measures, etc.) but can override the system’s recommendations with justification.
-The system is comparable to algorithmic profiling tools that are used in various countries to assist in job-seeker assessment and resource allocation (Scott et al., 2022; Lopez, 2019; Niklas et al., 2015) and which repeatedly highlight the limits of automation in interpersonal interaction (Bundesagentur für Arbeit, 2021; Niklas et al., 2015; Desiere and Struyven, 2021; Scott et al., 2022).
-Due to its documentation and exemplification of typical ADM systems’ challenges in public employment, the employment prediction algorithm was thus chosen as the use case for this study.
+3.1. Use case: The AMS employment prediction algorithm
+3.1.1. Description
+The AMS algorithm 555AMS stands for the Public Employment Agency (Arbeitsmarktservice).
+is a system developed to calculate the employability of job-seekers in Austria.
+It was created by a private company for the Austrian Public Employment Agency between 2015 and 2021 but was never used as a live system and put on hold in 2021 due to privacy objections (Allhutter et al., 2020).
+The system uses a logistic regression model trained on historical data to predict job-seekers’ employment chances based on demographic features (such as age, education, nationality, etc.) and work history.
+The outputs are a short-term and long-term employment score for each job-seeker (Gamper et al., 2020).
+These scores would serve as recommendations for the job-seekers’ counselors at the employment agency to assist in deciding about suitable support measures.
+Counselors could overwrite the system’s predicted scores of job-seekers but would need to give a reason for doing so (Allhutter et al., 2020; Holl et al., 2018).
+More information is provided in the appendix.
+3.1.2. Choice of use case
+Algorithmic tools that assist in assessing job-seekers and resource allocation have been deployed in various countries, including Germany (Bundesagentur für Arbeit, 2021), Austria (Allhutter et al., 2020), Poland (Niklas et al., 2015), and the Netherlands (Desiere and Struyven, 2021).
+However, the introductions of these applications also repeatedly led to sociotechnical conflicts (Scott et al., 2022).
+The deployment of the AMS algorithm was motivated by three overarching goals: a) increasing consultation efficiency, b) increasing support measure effectiveness, and c) reducing arbitrariness (Gamper et al., 2020).
+Detailed reports warned that counselors might over-rely on the algorithm or hesitate to overrule its suggestions (Allhutter et al., 2020).
+Further, the algorithm’s model and underlying data structure were predicted to discriminate against marginalized groups, who would lack the option to contest the system itself (Lopez, 2019).
+Transparency and ongoing scrutiny of the algorithm were listed as necessary measures to prevent these risks (Allhutter et al., 2020).
+As the AMS algorithm represents a larger class of algorithmic decision-making systems that spark public debate around their deployment in public institutions (Raji et al., 2022), it exemplifies how AI systems become matters of public interest and presents a suitable use case for our study.
 3.2. Explanation design
-Our explanation design was composed of 36 question-answer pairs about the employment prediction algorithm.
-Each question belonged to one of four categories, data (format, content, limitations), system details (features, model, examples), usage (operation by and interaction with users), and context (intention of deployment, target group, responsible actors), which were informed by work about the information needs of AI novices (Schmude et al., 2024).
-Each category was further divided into subtopics with three levels of increasing detail (base level, level 2, level 3).
-Every explanation contained a question (e.g., ”Who operates the system?”) answered with a brief text or image printed on an A5 paper sheet (Figure 2).
+3.2.1. Description
+The explanation design comprised 36 question-answer pairs about the AMS algorithm.
+Each question belonged to one of four categories, data (format, content, limitations), system details (features, model, examples), usage (operation by and interaction with users), and context (intention of deployment, target group, responsible actors).
+Each category was further divided into topics with three levels of increasing detail (base level, level 2, level 3).
+Every explanation was printed on an A5 paper sheet and contained a question (e.g., ”Who operates the system?”) answered with a brief text or image (cf.
+Figure 2).
 Participants first received an explanation overview (Figure 1) and the four base explanations and could request levels 2 and 3 at any time during the explanation phase (as depicted in Figure 3).
-The structure of the explanations was intended to i) facilitate uptake by providing information in a modular format and ii) provide different levels of both soundness (information levels) and completeness (subtopics) of information (Kulesza et al., 2013, 2015; Chatti et al., 2022).
-The selection of subtopics and information levels was guided by work on ADM systems in public employment (Scott et al., 2022; Niklas et al., 2015; Allhutter et al., 2020) and research on the prioritization of information by AI novices (Schmude et al., 2024; Weitz et al., 2024).
-The self-directed provision of information allowed participants to select explanations according to their interests (Conati et al., 2021) and information needs.
-The question-answer style was motivated by work on question-driven explanations (Liao et al., 2020) and was intended to encourage engagement.
-The explanations made use of different techniques that include feature importance, local and global explanations, examples, counterfactuals, and argumentative approaches (Figure 2).
-Lastly, the explanations were presented in an analog paper format to encourage interactions such as handing around paper sheets, pointing, and reading aloud.
-3.3. Study setup and procedure
-The study procedure consisted of several steps that are depicted in Figure 3.
-This section describes the introduction of the use case, the explanation phase comprising exploration, tasks, and the group decision, and the mechanisms for participants’ self-reports on understanding, confidence, and other aspects.
-Use case introduction.
-Participants received initial information about the employment prediction algorithm in the form of a mock newspaper article.
-The article provided key information and featured the perspectives of employers and employee associations.
-Exploration phase.
-After receiving the explanation overview and base level explanations of each category, participants had 15 minutes (focus groups) or 20 minutes (interviews) to become familiar with the structure and mark relevant explanations.
-Participants could freely decide which explanations to read and request and were encouraged to discuss them.
-Tasks.
-Participants then received a fictional job-seeker case example describing Mr.
-Harald G.555Details on the inspiration for this case example where omitted to adhere to anonymization policy, but will be re-inserted for the final version.
-, a fictional job-seeker with a brief backstory and a list of features that would be used to calculate his employment chances.
-Participants solved four tasks formulated as questions covering i) the job-seeker’s ability to change information stored about him (such as education, physical impairment, etc.), ii) his employability scoring by the employment prediction algorithm, iii) the support measures he would receive, and iv) his ability to contest the decision.
-All questions had multiple answers with one correct choice and could be answered with information present in the explanations in different categories and detail levels.
-Whereas tasks 1, 3, and 4 required mostly information retrieval, task 2 could be solved in two ways: by either giving an estimate based on the rough weightings in the system details base explanations or by calculating the precise employment score.
-Participants had 15 (focus groups) or 20 (single interviews) minutes to solve the tasks.
-During that time, they could access and request all explanations and discuss possible solutions.
-Decision task.
-Lastly, participants in focus groups had 10 minutes to discuss the system’s deployment and make a collective yes / no decision.
-If no consensual decision was reached in time, participants either decided themselves how to resolve the situation (e.g., majority vote) or received the option of stating conditions for the system’s deployment.
-Self-reports.
-Participants reported their understanding, decision confidence, and inclusion and voice in the group with 5-point Likert scales, their decision on the system’s deployment with yes / no voting ballots, and their choice of the explanation categories most ”helpful for their understanding” and ”influential to their decisions” by selecting the categories that applied.
-Self-reports were conducted using a color-coded system to ensure anonymity during focus groups.
-Each participant was assigned a color and given matching reporting material.
-During the second self-report, the investigator also asked interview questions about participants’ interaction with the explanations, their understanding processes, the most relevant information, and any additional situational questions.
-In single interviews, the collective decision and questions about inclusion and voice in the group were omitted.
+The explanations were presented in an analog paper format that allowed participants to interact with them physically and that facilitated social interactions, such as sorting, exchanging, pointing, and reading to each other.
+The full collection of explanations is depicted in the appendix.
+3.2.2. Design foundations
+The explanation design was intended to allow the users to explore information in a flexible and self-directed manner.
+To this end, the design used a modular structure, meaning that the explanations were divided into four information categories, which were again subdivided into topics and three levels of detail (base level, level 1, level 2).
+The four information categories, data, system details, usage, and context were based on research on AI novices’ information needs and covered both technical and sociotechnical system aspects (Schmude et al., 2024).
+The subdivision of explanation categories into topics and levels of detail organized this broad supply of information while accommodating different needs of information completeness and soundness (Kulesza et al., 2013) and introducing a degree of personalization (Chatti et al., 2022).
+The goal of the modular design was thus to create explanations that offered information on every aspect of the AI system, from which participants could select the most relevant according to their information needs and preferences.
+It further aimed to avoid limitations of “groupware” systems (Mandviwalla and Olfman, 1994) by supporting both individual and collaborative interaction, providing multiple user perspectives (e.g., user, decision subject), and synchronizing interaction with the material.
+The question-answer style was motivated by explanation design research (Liao et al., 2020, 2021) and was intended to improve user engagement and understanding by matching their thought processes.
+For example, when users learn that the system uses features to calculate job-seekers’ employment scores, they might ask what the exact weights of these features are and how they are calculated.
+This corresponds to the three levels of detail in topic A of system details (Figure 1).
+The explanations further used different explanation methods (Speith, 2022) such as feature importance, local and global explanations, examples, counterfactuals, and argumentative approaches (Figure 2).
+Information was presented in different formats but mostly relied on textual information and used highlighting, colors, and illustrations to emphasize key points.
+3.3. Study procedure
+We describe the procedure in the focus group and single interview setting (depicted in Figure 3).
+Like previous work in XAI (Chiang et al., 2023), we conducted individual and group settings to compare how the social setting would affect participants’ understanding and deliberation processes.
+3.3.1. Focus group procedure.
+Throughout the study, participants sat together with the investigator and could freely interact with each other.
+They first completed consent forms and questionnaires about demographics and knowledge about employment (domain knowledge) and AI systems (technical literacy).
+A round of introductions followed, where each group member stated their name and last interaction with AI to break the ice.
+The investigator then explained the study procedure and distributed a mock newspaper article introducing the AI use case (included in the appendix).
+Participants indicated their understanding, deployment decision, and decision confidence for the first time (3.4).
+The study’s explanation phase followed (including orientation, task, and decision phase), throughout which participants received and kept access to all explanations.
+In the beginning, the group received an overview of the explanations and all base level explanations, all other explanations could be requested at any time.
+After 15 minutes, participants received task sheets and had another 15 minutes to complete them, deciding independently whether they wanted to work together or individually.
+Finally, the group had 10 minutes to make a joint deployment decision (yes/no, with conditions allowed).
+A second round of individual reports followed, described in the next section, and then the investigator concluded the study.
+Focus group studies took around 90 minutes.
+3.3.2. Single interview procedure.
+The majority of the study procedure remained the same in single interviews.
+However, the explanation phase did not include a group decision phase; instead, the orientation and task phases were prolonged to 20 minutes each to provide individuals with the same total time as focus groups to interact with the explanations and individual reports about the group setting were omitted.
+The individual study setting was included to analyze how understanding processes changed depending on whether participants worked in a group or alone and whether the explanation design would support both settings.
+Single interviews took around 60 minutes.
+3.4. Study elements
+This section provides descriptions and motivation for the study elements: the introduction of the use case, the explanation phase comprising orientation, tasks, and the group decision, and participants’ individual reports.
+3.4.1. Use case introduction.
+Participants received initial information about the AMS algorithm in the form of a mock newspaper article inspired by an Austrian newspaper publication from 2019 (Szigetvari, 2018).
+The article provided key information about the system’s basic workings, goals, and deployment context and featured the opinions of employers and employee associations about its merits and risks.
+The presentation format was chosen to provide an introductory summary of the AI system using a familiar layout and non-technical language while highlighting both the pros and cons of the system’s deployment.
+Thus, The article served as a basic introduction to the use case, which aimed to approximate the amount of information participants might receive through the media.
+This way, participants received a baseline of information with which to assess their initial understanding, deployment decision, and decision confidence.
+Further, this introduction served to outline relevant aspects that could be explored using the explanations.
+3.4.2. Explanation phase (orientation, tasks, and group decision).
+For orientation, participants received the explanation overview (Figure 1) and base-level explanations (the first row of explanations in each category) and had 15 minutes (focus groups) or 20 minutes (single interviews) to become familiar with the structure and explore topics of interest.
+Explanations were provided as single A5 sheets to promote the physical sharing and exchanging of explanations.
+Participants could freely decide which explanations to request and read, and whether to share and discuss information with others.
+For the tasks, participants received the case of Mr.
+Harald G.666The case example was inspired by Allhutter et al.
+(2020) and adapted to this study, as depicted in the appendix., a fictional job-seeker with a brief backstory and a list of features.
+Participants had 15 minutes (focus groups) or 20 minutes (single interviews) to solve four tasks pertaining to this case.
+All questions could be answered with information from explanations in different categories and levels of detail.
+Whereas tasks 1, 3, and 4 required locating information, task 2 could be solved in two ways (aside from guessing): by either giving an estimate based on the rough weightings in the system details base explanations or calculating the precise employment score.
+Participants could access and request all explanations and discuss possible solutions.
+These were the four tasks (correct answers underlined):
+Task 1: Can Harald change the data stored about him (e.g., to correct it)? (yes / no)
+Task 2: In which group of employment chance does the system categorize Harald?
+(high (¿66%) / medium (¡66% & ¿25%) / low (¡25%))
+Task 3: Which support measures will Harald receive?
+(qualifying measures, such as courses and training / stabilization and increased supervision / none)
+Task 4: Can Harald appeal against this decision? (yes / no)
+For the group decision, focus group participants had 10 minutes to discuss the system’s deployment and were asked to collectively decide whether to accept or reject it.
+This was meant to simulate a small referendum in which each participant’s voice counted for the final outcome.
+If no consensual decision was reached in time, participants were asked how the situation should be resolved (e.g., majority vote).
+They were further made aware of the option to state conditions for the system’s deployment.
+3.4.3. Individual reports I and II
+Participants were asked for individual reports before and after the explanation phase.
+At both points, participants reported understanding (5-point scale), deployment decision (yes / no), and decision confidence (5-point scale) to examine the effect of the explanation phase.
+In report II, participants also reported the explanation categories that were most helpful for their understanding (multiple choice) and most influential to their decisions (multiple choice); focus group participants additionally reported perceived inclusion in the group and the discussions’ influence on their decision (5-point scales).
+During report II, the investigator asked participants interview questions about their interaction with the explanations, understanding processes, prioritized information, and situational aspects.
+The list of interview questions is included in the appendix.
+To prevent influence between participants’ reports, individual reports in focus groups were conducted anonymously and re-assigned by the study examiner using a color-coded reporting system (Figure 4).
+Each participant was assigned a color and received the material for all ten individual reports.
+Participants took the corresponding paper slip for each report, wrote their answer, and threw it in a gathering container that hid it from view.
+The gathered reports were then collected and recorded by the investigator at the end of the study.
 3.5. Analysis
 All focus groups and interviews were audio-recorded and transcribed.
-These transcripts, participants’ self-reports during the studies, and the investigators’ field notes, provided the data basis for analysis.
-For RQ1-Explanations, we focus on how the explanations helped and impacted participants’ understanding of the ADM system and examine the differences between individual and group settings.
-For this, we compare participants’ self-reported understanding before and after the explanation phases, their performance in the study tasks, and which and how many explanations were requested.
-Further, we conduct an inductive thematic analysis (Braun and Clarke, 2006) of participants’ responses to the investigator’s interview questions, capturing articulations about the strengths and weaknesses of the explanations’ design, the impact of the time limit, understanding barriers, and information perceived as most relevant.
-While the quantitative items thus serve to characterize the diversity of participants’ perceptions (Maxwell, 2010) and facilitate qualitative exploration (Weiss, 1995), they are not intended to invoke ”inference […] of greater generality” (Maxwell, 2010) nor impose a mental model based on variance theory (Patton, 1990).
-For RQ2-Deliberation, we first focus on how participants shared and combined information to solve the study tasks and, second, on how explanations enabled group deliberation.
-For the first part, we conduct an inductive and deductive thematic analysis of the groups’ interaction processes with explanations, connecting our descriptions of these interaction processes to mechanisms for collaborative success and failures as described by Nokes-Malach et al.
-(2015).
-For the second part, we compare self-reported decision confidence before and after participants received and discussed the explanations, summarize their feedback about the group setting, and conduct an inductive thematic analysis of focus group conversations in which explanations drove group deliberation and convergence, facilitating the opinion-forming of participants.
+These transcripts provided the data basis for the thematic analysis.
+Participants’ individual reports, task solutions, decisions, and the investigators’ field notes provided further data for within- and between-subject comparisons of understanding and decision-making.
+3.5.1. Thematic analysis.
+For both research questions, we conducted thematic analysis (Braun and Clarke, 2006) of participants’ articulations to develop a qualitative account of their understanding and decision-making processes.
+To this end, we developed inductive codes in the first pass and refined them in the second pass on the transcriptions.
+The resulting inductive code base was structured along the overarching categories of understanding, deliberation (decision-making processes, arguments), opinions (e.g., about AI and policy choices), and experiences (e.g., anecdotes and lived situations).
+The full code-book is provided in the appendix.
+We further highlight that while the quantitative items in participants’ self-reports serve to characterize the diversity of participants’ perceptions and facilitate qualitative exploration (Weiss, 1995), they are not intended to invoke “inference […] of greater generality” (Maxwell, 2010) nor impose a mental model based on variance theory (Patton, 1990).
+3.5.2. RQ1-Explanations.
+RQ1 focuses on the explanations’ impact on participants’ understanding of the study’s use case and the differences between individual and group settings.
+We employ a triangulation approach (Carter et al., 2014) by using three ways of analyzing the explanations’ effect on participant understanding.
+Firstly, comparing participants’ individual reports before and after the explanation phase was a subjective indicator of changes in their understanding and decision confidence.
+Secondly, participants’ answers to the four study tasks indicated their factual or testable understanding.
+Lastly, participants’ verbal reports during and after the explanation phase were used to analyze their understanding processes and barriers thematically.
+In a deductive analysis, we further compared their interactions to mechanisms of “collaborative success and failure” (Nokes-Malach et al., 2015) and the “six facets of understanding” (Wiggins and McTighe, 2005).
+With this three-part combination, we examined participants’ subjective understanding, their information gain, and the cognitive processes of their understanding.
+This choice was motivated by educational psychology research outlining that understanding cannot only be elicited through test questionnaires (Sato et al., 2019) but involves emotional (Wiggins and McTighe, 2005) and meta-cognitive processes (Veenman et al., 2006) that are equally important.
+Our focus on understanding is motivated by previous XAI research, which highlights the importance of understanding in decision-making processes (Langer et al., 2021; Shin, 2023; Hoffman et al., 2023).
+3.5.3. RQ2-Deliberation.
+RQ2 focuses on how participants formed opinions about the AMS algorithm, weighed the pros and cons of its deployment, and settled on a deployment decision.
+To this end, we compare participants’ decision confidence before and after the explanation phase.
+We further conduct an inductive and deductive thematic analysis of participants’ interactions in both settings to connect them to the “elements of deliberation” (Stromer-Galley, 2007) – a set of characteristics that outline deliberation processes.
+Based on this, we analyze when participants used arguments (grounded, defensible positions), opinions (personal judgments on things, values, states), and personal experiences (Stromer-Galley, 2007; Mercier and Landemore, 2012) to consider the system’s deployment.
+For single interviews, we examine participants’ responses to interview questions during the study to examine their reasoning process and “internal deliberation” (Mercier and Landemore, 2012).
+Lastly, to account for one of the most prevalent cognitive biases in group settings, we examine focus groups for occurrences of groupthink (Janis, 1971; Baron, 2005) – an effect that sets in when concurrence-seeking in groups overrides realistic argumentation and discussion.
 3.6. Participants
-Recruitment.
-Table 1+2 provides an overview of the study participants.
-Participants were recruited through cooperation with civil society organizations, an employment agency, and the authors’ extended network.
-Recruitment criteria included full legal age and no technical knowledge about AI systems, i.e., AI novices.
+3.6.1. Recruitment.
+Table 1 and 2 provide an overview of the study participants.
+Participants were recruited through cooperation with civil society organizations, an employment agency, public calls for participation, and the authors’ extended network.
+For focus groups, the authors contacted staff from these organizations known from previous studies or used channels of general inquiry to describe the study and invite participation.
+Interested organizations all offered to support the recruitment process by coordinating with the authors on selecting and inviting potential participants and finding a place and time to conduct the studies.
+Groups A, B, C, E, F, G, and H were organized this way.
+Group D was recruited through the authors’ network and was equally composed of people who had previously been job-seeking.
+For individual studies, participants were recruited using the same channels, and calls for participation were further advertised on screens and information boards throughout different city districts.
 All studies were conducted in person in office or public spaces.
@@
 Concerning the participant sample size, we are guided by research on qualitative methods, which suggests that the number of participants should be determined by code and meaning saturation (Hennink et al., 2017).
-Choice of participants.
-Participants were selected to be representatives of one of three roles: domain experts, decision subjects, or members of the general public.
-We define domain experts as people who are competent in the field that the ADM system is used in, such as job counselors or advisors (groups 2, 3, 7, and 8).
-We define decision subjects as people who would potentially be impacted by the system’s decision, such as job-seekers and people who had previously been job-seeking (groups 5 and 6).
-All remaining participants are considered as members of the general public and were included to test changes in explanation effects and participants’ perceptions (groups 1 and 4).
+3.6.2. Recruitment criteria.
+Participants were required to be of full legal age and AI novices, i.e., to have no technical knowledge or expertise about machine learning systems as described in Section 2.2.
+These criteria were screened in a pre-questionnaire before invitation to the study using two questions: “How would you rate your knowledge of algorithms?” and “How would you rate your knowledge of artificial intelligence (AI)?”.
+Each question could be answered on a scale corresponding to “no knowledge at all” to “professional and detailed knowledge”.
+Here, the first question elicited technical knowledge, as participants familiar with AI tools might have rated their AI expertise as high but were unlikely to know about algorithms without a strong technical interest or background.
+3.6.3. Group composition.
+Participants were further selected to be representatives of one of three roles: domain experts, decision subjects, or members of the general public.
+We define domain experts as people who are competent in the field that the AI system is used in, such as job counselors or advisors (groups B, C, G, and H).
+We define decision subjects as people who would potentially be impacted by the system’s decision, such as job-seekers and people who had previously been job-seeking (groups E and F).
+All remaining participants are considered members of the general public and were included to test changes in explanation effects and participants’ perceptions (groups A and D).
 The study was conducted with separate participants in three pilot groups to test and refine the explanation design and study procedure.
@@
 B2 | 76 | A-levels | Retired | Group G | G1 | 37 | University | Executive staff | |
-| B3 | 46 | University | Social counselor | G2 | 49 | University | DSGVO officer | ||
+B3 | 46 | University | Social counselor | G2 | 49 | University | GDPR officer | ||
 B4 | 70 | A-levels | Retired | G3 | 44 | Secondary school | Training counselor | ||
@@
 4. Results
-In this section, we present our results, structured according to our research questions: how question-driven, modular explanations support understanding in individual and group settings (RQ1, Section 4.1) and how groups used explanations666We describe every question and answer pair as one ”explanation” and their entirety as a ”collection of explanations” (Section 3).
-to solve the study tasks and form opinions about the system’s deployment (RQ2, Section 4.2).
-Participant labels denote the study setting (focus group: A–H / single interviews: S) and the individual participant, as listed in Tables 1+2.
-4.1. RQ1-Explanations: How does a question-driven, modular explanation design support AI novices’ understanding in individual and group settings?
-Participants received a collection of 36 explanations to solve the study tasks and decide about the system’s deployment individually and (in focus groups) collectively (Section 3).
-The following subsections describe the explanations’ effect on participants’ understanding in both settings (4.1.1), design aspects that impeded understanding (4.1.2), the time limit’s effect on prioritization of information (4.1.3), the interaction differences between single and group settings (4.1.4), and participants’ ratings of explanations’ helpfulness and influence (4.1.5).
-4.1.1. Explanations had mixed effects on participants’ understanding.
-| Focus gropus | ||||||||||||||
-| Increased understanding | F2 | C3 | F1 | B4 | A2 | H1 | H2 | E1 | H4 | |||||
-| Self-reported understanding | +++ | ++ | ++ | + | + | + | + | + | + | |||||
-| Decision confidence | - | ++ | = | +++ | + | + | + | = | = | |||||
-| Unchanged understanding | H3 | B3 | A3 | G1 | H5 | A1 | A4 | C2 | D3 | E2 | C1 | E3 | F4 | F5 |
-| Self-reported understanding | = | = | = | = | = | = | = | = | = | = | = | = | = | = |
-| Decision confidence | +++ | ++ | + | + | + | = | = | = | = | = | - | - | - | - |
-| Decreased understanding | D1 | G4 | D2 | G2 | B1 | B2 | F3 | G3 | ||||||
-| Self-reported understanding | - | - | - | - | - - | - - | - - - | - - - | ||||||
-| Decision confidence | + | + | = | = | +++ | = | + | = | ||||||
-| Single studies | S6 | S12 | S8 | S7 | S4 | S9 | S1 | S3 | S11 | S5 | S10 | S2 | ||
-| Self-reported understanding | +++ | + | + | = | = | = | = | = | = | - | - | - - | ||
-| Decision confidence | ++ | + | = | +++ | + | + | = | - - | - - | ++ | = | - |
-| Focus groups | Single interviews | |||||||||
-| ID | 1st decision | Collective decision | 2nd decision | ID | 1st decision | Collective decision | 2nd decision | ID | 1st decision | 2nd decision |
-| A1 | No | Yes | F1 | Yes | Yes | S1 | Yes | Yes | ||
-| A2 | No | Yes | F2 | Yes | Yes | S2 | No | No | ||
-| A3 | Yes | Yes | F3 | Yes | No | S3 | No | No | ||
-| A4 | No | Yes | Yes | F4 | No | No | S4 | No | No | |
-| B1 | No | No | F5 | Yes | No | Yes | S5 | Yes | No | |
-| B2 | Yes | No | G1 | Yes | Yes | S6 | No | No | ||
-| B3 | Yes | No | G2 | No | No | S7 | No | No | ||
-| B4 | Yes | No | No | G3 | Yes | No | S8 | No | No | |
-| C1 | No | No | G4 | No | No | No | S9 | No | No | |
-| C2 | No | No | H1 | No | No | S10 | Yes | Yes | ||
-| C3 | No | No | Yes | H2 | No | No | S11 | Yes | No | |
-| D1 | No | No | H3 | No | No | S12 | No | No | ||
-| D2 | No | No | H4 | No | No | |||||
-| D3 | No | No | Yes | H5 | No | No | No | |||
-| E1 | No | No | ||||||||
-| E2 | No | No | ||||||||
-| E3 | No | No | Yes | |||||||
-Participants reported their understanding before and after receiving and discussing the explanations.
-Table 3 shows that self-reported understanding decreased and increased for comparable numbers of participants, while also often remaining unchanged.
-Participants gave divergent feedback on how they perceived the explanations, describing their structure and design as ”nicely presented” (A2, C2), ”good to get an overview” (C3, H4), ”well-selected” regarding the questions (G2), ”active and controllable” (S8), and as a good primer (F4).
-Critical comments described them as ”too much” (D1, S4), ”at first overwhelming” (H5), ”confusing” (B1, D1), ”theoretical” (D2), and ”quite complex” (C2).
-However, this was not always seen as an issue: ”I believe that it is complex.
-But I also believe that it needs to be.” (D2).
-Participants gave nuanced reasons why their perceived understanding decreased.
-Several focus group participants stated they did not fully understand the technical explanations, including numbers and a formula (B3, C3, G4).
-In contrast, participants in single interviews with comparable backgrounds met fewer challenges with technical explanations, often stating that they appreciated applying this knowledge when solving the tasks (S2, S8, S11).
-In these cases, the setting likely mediated the ease of interaction, as participants in single interviews could concentrate without distraction (further described in Section 4.1.4+ 5.2).
-However, irrespective of the setting, confrontation with mathematical details seemed to re-adjust participants’ benchmarks for ”high understanding” and calibrate their perceptions.
-Likewise, several participants stated that the different subtopics and levels of detail in the explanations helped them identify the information they were still missing to judge the whole system (B1, F3, G3).
-Both observations indicate that the explanations revealed participants’ blind spots and helped them to overcome the ”illusion of explanatory depth” (Rozenblit and Keil, 2002) – an overestimation of one’s understanding due to limited knowledge and misled intuition – which is exemplified by S2’s comment: ”The more I look at the explanations now, the more I realize that I simply assumed a lot of things but didn’t actually know how it worked.” In short, participants gauged their understanding relative to the amount of information available.
-That is, since participants only had time to read through parts of all available explanations and were aware that they had not seen all of them, they reported a decrease in perceived understanding when it, in fact, improved.
-The implications for measuring explanations’ effect on understanding are discussed in Section 5.
-4.1.2. The explanation design was not equally suited to every participant group.
-In contrast to participants whose perceived understanding was improved and calibrated by the explanations, we observed that the explanations did not fully support understanding or interaction for some groups.
-In these cases, the explanations’ language and textual format were listed as impeding factors.
-The language style was described discrepantly as ”well understandable” (P3, E3, H1) but also ”very difficult” (E2, F2) and ”too technical” (B3, B4) depending on participants’ background.
-Crucially, in Groups E and F (decision subjects), participants stated that the language posed challenges.
-E2 compared the explanations to official letters they receive: ”I get letters that I don’t understand from the court.
-I understand every single word, but I don’t understand the context.” In line with this, job counselors in Groups C, G, and H stated that explanations in simple language would be required if they were supposed to be useful for their job-seeking clients.
-Similar comments were made about the explanations’ heavy reliance on textual information, described as demanding and requiring plenty of concentration.
-While previous work found that textual explanations can be effective despite users’ dispreference (Szymanski et al., 2021) or are even preferred (Weitz et al., 2024), a more interactive and visual design would likely facilitate interaction: ”If there’s a program where, when you click on it, it opens as a window and you can read about it, that would be great, and you can deal with it at your own pace.” (H5) Both, simpler language options and an interactive version, are thus useful suggestions for improving the explanation design to ensure that explanations are helpful for a broader range of stakeholder groups.
-Further, Section 4.2.1 describes how these understanding challenges impacted group interaction.
-4.1.3. Time limits led to forced prioritization of information, with different effects on participants.
-All participants had 40 minutes to become familiar with the explanation structure, request new explanations, solve the given tasks, and, in focus groups, make a collective decision (Table 4).
-The time limit was meant to incentivize participants to prioritize information and, where possible, work together to process larger amounts.
-This succeeded for most focus group’s participants, who stated that ”when you have little time, the teamwork is extremely good” (B2) and that ”even a brief exchange is enough, otherwise it becomes too much” (C3).
-However, the time limit also led to induced stress, especially in single interviews (S2, S6, S7, S8, S9, S11).
-S9 commented that this reminded them of their work as a social center manager: ”Basically, the time to read about something new does not exist.” (S9).
-This points to the need for brief and compact synopses complementing the detailed explanations.
-Regarding prioritization, many participants liked choosing explanations according to their own interest, which gave the feeling of being in control (S8), provided a direction for exploration (H1), and justified skipping explanation categories in favor of others (S6, S9).
-Allowing for the self-directed selection of relevant information by users could thus be beneficial to encourage interaction and engagement.
-4.1.4. Some single participants worked through surprising amounts of explanations.
-Compared to focus groups, participants in single interviews often requested and engaged with equal or even higher amounts of explanations (as depicted in Figure 4).
-Often, participants in single interviews requested ”the whole context” (S3) or ”all of level 3” (S4) of one category, appearing to deal with the amount of information easily.
-A likely reason for this is that single interviews supported more focused and in-depth interaction with explanations compared to group settings.
-Participants in single interviews also performed better in the study tasks overall and, as mentioned above, often calculated the job-seeker’s precise employment chance in task 2, which required following a multi-step process.
-This is noteworthy, as none of the groups, aside from the pilot groups, attempted to calculate the precise chance.
-This cannot be explained by a difference in educational or professional background, as even focus groups with university graduates in domain expert professions did not perform these calculations, again indicating that the reason might be the difference in supported interaction.
-For example, the parallel processing of information might have impeded understanding of the system’s entirety in groups, as described by G3: ”You only pick out the most important things.
-And then you haven’t gathered information about the thing as a whole.” This highlights the benefits and trade-offs of both settings, suggesting that individual engagement with explanations might be helpful before discussing them in a group setting.
-The implications of this are discussed further in Section 5
-4.1.5. All explanation categories were deemed helpful and influential, but system details and context were most often mentioned.
-In the second self-reports, participants were asked which explanation category was most helpful for their understanding, most influential for their decisions, and whether specific explanations stood out.
-Regarding understanding, participants in focus groups responded with a rather even distribution over all categories, as depicted in Figure 5, and frequently mentioned that ”all of them [are relevant]… I don’t think you can leave anything out, really” (D3).
-Both participants in focus groups and single interviews further stated that system details were crucial, as they were perceived as ”tangible” (S6), ”concrete” (S8), and to be ”the devil in the detail” (B4).
-Specifically, participants found explanations about the features and weighting most important (D1, G1, H3, S5), which G4 commented with: ”That is the central point, the basis of the whole system.” Even though context was less often reported as an influential explanation category, participants requested these explanations most often (Figure 6) and frequently mentioned affected persons’ inability to contest decisions (C1, D1, E1) and the system’s political background (A1, S3) as relevant – both belonging to context.
-Notably, participants in single interviews found data much less helpful and less influential, stating, e.g., that GDPR would likely cover the important aspects of it (S8) or that they prioritized another category in the time available: ”I didn’t deal with it in such detail […].
-You read it [the base level] and already feel informed.
-As is often the case with data protection, ’yeah, close enough’.” (S6).
-Other participants stated that the technical details were not too important for them but that they instead focused on how the system was used practically and which values it represented (S3, S10), corresponding to the categories usage and context.
-4.2. RQ2-Deliberation: How do the explanations support group interaction with information and deliberation processes?
-In this section, we describe how our explanations supported group collaboration in exchanging information (Section 4.2.1) and in discussing and forming opinions about the system’s deployment (Section 4.2.2).
-4.2.1. Explanations support collaborative success if group dynamic is constructive.
-We identify several processes that groups used to interact with information collectively.
-Our descriptions of these interaction processes include splitting (by category, order, priority), sharing (swapping, reading out loud, proposing solutions), analyzing (discussing, explaining), and combining (merging, locating, correcting).
-The degree to which groups used these processes appeared to depend on the familiarity between group members, their domain expertise, and their understanding.
-In the following, we analyze three exemplary conversation excerpts that showcase different forms of interaction in focus groups while also demonstrating the diversity in collective interaction.
-For each excerpt, we describe the interaction processes that emerged and then connect them to mechanisms that drive collaborative failure or success (marked by quotation marks) as outlined by Nokes-Malach et al.
-(2015).
-Case 1: Information sharing leads to increased engagement and construction of common ground.
-Group G was composed of participants who hold executive positions in a civil society organization and are responsible for the management and training of employees.
-Participants in this group showed high engagement with information, frequently sharing explanations and exchanging their perspectives by tapping on their professional experience.
-The excerpt begins several minutes after participants received the explanations just as G4 examines an explanation comparing the algorithmic score for two job-seekers:
-G4: That’s bad, the two of them. Look, I’ve got this [system details-C-base], ”What chances would the system calculate for Martin and Schifteh?” (sharing: reading out loud)
-G2: Schifteh is probably worse off, isn’t she? (analyzing: discussing)
-G4: Schifteh has 30% chance of employment and Martin 52%, even though Schifteh has a degree and would be working in the IT sector.
+In this section, we present our results as answers to our research questions: How question-driven, modular explanations777We note again that with “explanation” we mean a question and answer pair and with “explanations” we mean the collection of all 36 explanations (Section 3).
+support understanding in individual and group settings (RQ1, Section 4.1) and how AI novices used explanations to form opinions and decide about the system’s deployment (RQ2, Section 4.2).
+Participant labels denote the study setting (focus group: A–H / single interviews: S) and the participant ID, as listed in Tables 1+2.
+To distinguish themes in the analysis, inductive themes are italicized, while ‘deductive themes’ are put in quotes.
+4.1. RQ1-Explanations: How does a question-driven, modular explanation design support AI novices’ understanding in groups and individual settings?
+To examine how AI novices used the explanations to understand the study’s use case, we analyzed their self-reports, articulations, and interactions in both settings.
+We found that each setting supported different aspects of understanding, suggesting a trade-off.
+We first describe how the explanations contributed to shared understanding and ‘collaborative success’ in groups (4.1.1) and continue with the explanations’ role in instances of ‘collaborative failure’ (Nokes-Malach et al., 2015) (4.1.2), summarized in Figure 5.
+We then describe individuals’ interactions with the explanations (4.1.3), participants’ feedback on the explanation design (4.1.4), and summarize the benefits and drawbacks of both settings for XAI (4.1.5).
+4.1.1. Groups’ benefits: Shared understanding and increased engagement.
+In the best cases, groups leveraged the modular explanation structure to use distributed cognition (Keil, 2003), meaning that participants processed information in parallel and then combined it.
+We use the term shared understanding to capture interactions that realized distributed cognition.
+Examples of such interactions included locating information together, sharing information with others, discussing interpretations, debating task solutions, and querying and explaining (a question by a group participant invites other participants to contribute).
+The explanations only afforded this set of interactions to groups, as they required social interaction with other participants.
+For example, in Group C, participant C1 read the first study task aloud and asked for input (querying), after which the group discussed solutions (explaining):
+C1 Can Harald change the data stored about him? Yes, he can certainly change it, can’t he? […]
+C3 Which stored data, the one down there? [points at Harald’s demographic features]
+C1 Yes, just that.
+C3 49 – no, male – no.
+The apprenticeship – no, Austria – he can still change that.
+Duty of care – he could get married or have children.
+He could change his service sector.
+He could change his career.
+Impairment…
+C1 Well, what is meant by ‘change’? When he enters the data, he can change the data. He doesn’t have to specify the knee problem. […]
+C3 So he can change it.
+C2 Yes.
+Interactions such as querying and explaining and discussing interpretations rely on collaboration between participants to ‘share working memory resources’, ‘complement others’ knowledge’, ‘re-expose information’, and ‘correct errors’.
+These aspects of collaboration are described as cognitive mechanisms of ‘collaborative success’ (Nokes-Malach et al., 2015) and provide groups with multiple ways to tackle explanations.
+For example, participants tended to work through information about usage and context alone or in pairs but raised explanations with the group when they were difficult or piqued their interest.
+This was often the case with explanations about system details, which included the most numerical information but also were an important key to understanding the system’s calculations.
+We describe the process of using other’s understanding to close gaps in one’s own as outsourcing (Keil, 2006).
+As understanding AI systems involves interacting with a variety of different information categories (e.g., technical, political, social), outsourcing provides a way to hand information to the team member most competent in this category.
+For example, in Group B, B2 expressed their appreciation for B3’s help in solving study task 2: “It was a math problem.
+You [B3] filtered it out well.
+It was very analytical.
+With your help, we were able to recognize these weak points.” In contrast to querying and explaining, which participants used to invite input or spark conversation, outsourcing was thus used for the active delegation of an impeded understanding process.
+Whether groups used this collaboration depended on participants’ relationships and the group’s social dynamics.
+The explanations also served to support these ‘social mechanisms’ of collaboration, by drawing the group’s attention to certain aspects of the AI system and encouraging them to share their experiences and opinions.
+We present an excerpt from Group G as an illustration.
+Here, participant G4 shared an explanation that documented the algorithm’s impact on two job-seekers (‘joint management of attention’), which prompted G2 and G3 to discuss interpretations (‘increased engagement’).
+This interaction established ‘common ground’ that the group later used for deliberation:
+G4 That’s bad, the two of them. Look, “What chances would the system calculate for Martin and Schifteh?”
+G2 Schifteh is probably worse off, isn’t she?
+G4 Schifteh has a 30% chance of employment and Martin 52%, even though Schifteh has a degree and would be working in the IT sector.
 And Martin has compulsory schooling and works in the cleaning sector.
 Martin’s chances of employment are almost twice as high as Schifteh’s.
-(sharing: reading out loud; analyzing: explaining) […]
-G3: I think that’s a bit weird. […] Because if she can speak English very well and has the specialist knowledge that our IT sector needs… (analyzing: discussing)
-G4: She gets two minuses for living in [a specific city district777Information was omitted to adhere to anonymization policy and will be re-inserted for the final version.].
-People who live in [this district], or in a district that is not well regarded, have totally negative… (analyzing: explaining)
-G1: Yes, and here you have it in writing, I’ll have to look at that too. (sharing: swapping)
-Framed in the context of mechanisms that drive collaborative success, G4 first draws the group’s attention to an explanation they deem relevant, describing the algorithm’s treatment of two different cases of job-seekers.
-G4 reads the information to the group, which prompts G2 and G3 to contribute their own perspectives, ”increasing engagement” while also ”cueing their prior knowledge”.
-As a result, G1 is also engaged by the information and articulates interest in looking at it.
-In combination, the interaction has established ”common ground” by sharing information the group can use for further collaboration.
-Interactions like these were common in Groups G and H (job counselors and trainers), indicating that the explanations’ format and content suited these compositions.
-Case 2: Discussion of solutions leads to correction.
-Group B consisted of both staff members and volunteers of a civil society organization.
-The group interaction was mostly driven by one participant who requested and shared most of the information and thus acted as a guide and informer, potentially driven by her role as deputy manager.
-While this led to a more one-sided dynamic, the other participants later appreciated outsourcing part of the understanding process to B3.
-In the excerpt below, B3 proposes a solution for task 2 (chance of employment that the system assigns Harald G.; group low, medium, or high) that is then discussed:
-B3: I chose “group low” for task 2. (sharing: proposing solutions)
-B4: I put him [Harald G.] in ”group medium”. (sharing: proposing solutions)
-B1: Why not medium, B3? (analyzing: discussing) […]
-B3: With 49 years he has one minus. And the knee problems could be a “health impairment”. (analyzing: explaining)
-B2: But if he has a job where he sits, that doesn’t matter. (analyzing: discussing)
-B3: But I don’t see that this is considered anywhere.
-And I think the fact that he hasn’t been employed much in the last 4 years has a very negative effect.
-What information do you have that speaks against that?
-(combining: correcting; analyzing: discussing)
-B1: Because with 50, he would still be in ”group medium”.
-[…] Then he’s male and from [this country888Information was omitted to adhere to anonymization policy and will be re-inserted for the final version.].
-(analyzing: explaining)
-B3: Male is only 0, the citizenship is also only 0. That doesn’t make up for anything. (combining: correcting)
-[B1 agrees] […]
-B4: Then I’ll correct it to low again. (combining: merging)
-The group members in this situation ”negotiate multiple perspectives” on the correct solution, notice discrepancies in their ”common ground,” and ”re-expose” themselves to information to ”correct errors” in their knowledge.
-B3 led the interaction and corrected other participants repeatedly, which could have caused adverse reactions from them (e.g., due to being ”evaluated”).
-However, B2 and B4 in fact later articulated appreciation for this interaction, stating that ”B3 filtered that out well” (B2) and that ”it was helpful that you [B3] always read it out and corrected it straight away” (B4).
-Correction was therefore seen as constructive to the discussion, possibly because of B3’s leading role in the organization and because participants already knew each other.
-Case 3: Attempted error-correction fails to create common ground.
-Group E was composed of people who were job-seeking at the time of the interview or in the past.
-Members of this group were not familiar with each other.
-Interaction was mostly fragmented and participants rarely shared or discussed information.
-In the excerpt below, participants were asked about their responses to task 4 (whether Harald G.
-is able to contest the system’s decision):
-E2: I say you can always object. Even if you’re locked up, you can go to court somewhere. (sharing: proposing solutions)
-E3: Yes, I think so too. (sharing: proposing solutions)
-E1: Well, I’ve already read that [the explanation], and it says that you can’t appeal against it.
-But it says that you could if a legal basis were first created, which doesn’t exist at the moment.
-(combining: correcting; analyzing: explaining)
-Investigator: Okay. Is there anything you [E2, E3] would like to change now, based on this information?
-[E2 & E3 indicate that they do not]
-Investigator: You would say that he can contest the decision?
-[E2 & E3 agree]
-Despite E1’s reference to an explanation and their attempt to ”correct errors” in what E2 and E3 believe to be the solution, the interaction fails to establish ”common ground” or ”negotiate multiple perspectives”, as the different viewpoints are neither elaborated nor discussed collectively.
-Possibly, E2 and E3 did not believe that E1 or the explanation were correct, or they felt embarrassed for being corrected.
-Further, E1 had no particular motivation to persuade the other participants of the correct solution, as E1 later self-described as ”introverted” and stated that they usually preferred individual over group settings.
-Comparable mechanisms unfolded only in Group F, where participants tended to read explanations themselves or disengage rather than interact with other participants.
-This might have had multiple reasons: First, both Groups E and F indicated a low interest in the topic of labor; second, participants in both groups had trouble understanding the explanations due to language barriers and the heavy reliance on text; and third, participants in both groups were unknown to each other.
-In combination, this might have created a group dynamic that was not constructive, i.e., neither argumentative (Mercier and Sperber, 2011), solution-oriented (Niculae and Danescu-Niculescu-Mizil, 2016), or familiar (Johnson and Johnson, 1985), which disrupted collaborative interaction with the explanations as well as discussion processes.
-Summary.
-The conversation excerpts in cases 1 and 2 demonstrate that the explanation design supported various collaborative interaction processes.
-These processes led participants to create common ground and find a task solution through error correction, which are markers of ”collaborative success” (Nokes-Malach et al., 2015).
-In contrast, case 3 demonstrates that the explanations did not always achieve their purpose of facilitating interaction, highlighting that the interplay of explanation format and group dynamics must be considered in the design of explanations in group settings.
-4.2.2. Explanations in group settings allow for informed discussion of pivotal topics and opinion formation about system deployment.
-For most participants, deciding about the employment prediction algorithm’s deployment was a clear choice: 7 out of 8 groups voted ”No”, and a majority of participants also voted ”No” in their individual decisions (Table 4).
-Confidence in these decisions increased strongly after the explanation phase (Table 3), with most participants indicating that they now felt better informed.
-Comparing single interviews with focus groups, two effects were notable: First, a third of focus group participants changed their vote in the second self-report, and second, participants in single interviews stated repeatedly that they would have found a discussion helpful.
-This is further underscored by participants’ feedback on the focus groups: Most felt included in the discussions (26 of 31) and were influenced in their decisions by them (18 of 31).
-Specifically, participants appreciated the group size, which allowed for balanced contributions and productive discussion (C1, D1, E2), and valued the exchange of perspectives: ”Everyone understands a little differently and sees a different focus, I think it’s really important to talk about that.” (H5).
-We see this as evidence that the explanations were suitable for the collaborative setting and supported group deliberation.
-This section thus aims to describe the synergy effects that emerge when combining explanations with group discussion and how this affected participants’ perceptions.
-In the following, we describe three conversation excerpts in which explanations lead groups to discuss their core concerns and beliefs.
-Case 1: Group discusses changes of perspectives on the system’s deployment.
+[…]
+G3 I think that’s a bit weird. […] Because if she can speak English very well and has the specialist knowledge that our IT sector needs…
+G4 She even gets two minuses for living in Favoriten [a city district]. […]
+G1 Yes, and here you have it in writing, I’ll have to look at that too.
+A later excerpt in Section 4.2.3 further shows that explanations also led participants to ‘negotiate multiple perspectives’.
+Addressing both sides of collaboration, cognitive and social, here is an important goal when supporting group understanding.
+It can be assumed that when explanations succeed in doing both, they can provide participants with more comprehensive and more complex information than individual settings.
+In line with this, previous work in HCI has found that group interaction boosts task performance compared to individual settings (Karadzhov et al., 2023).
+However, in our study, individuals surprisingly performed better in the study tasks than groups, as discussed in Section 4.1.3.
+Even so, we argue that the set of interactions (shared understanding) enabled through the combination of our explanation design and the group setting presents important pathways to help AI novices understand algorithmic systems.
+These interactions can be especially useful when group members have different domain expertise and information needs, as they can use complementary knowledge, memory, and perspectives to make sense of information.
+At the same time, as noted earlier, these interactions are partly dependent on the group dynamic.
+Positive interactions like those described were especially frequent in Groups G and H (job counselors and trainers), where participants knew and trusted each other.
+In contrast, the next subsection describes instances where groups encountered challenges in understanding, illustrating the importance of social mechanisms.
+4.1.2. Groups’ drawbacks: Process loss and susceptibility to social dynamics
+In some of the focus groups, participants lost track of information, forgot their train of thought, or abandoned understanding altogether.
+We summarize these effects under the term impeded understanding and its final result as abandoned understanding.
+We found that impeded understanding occurred due to explanation design flaws and co-occurred with adverse social dynamics, resulting in ‘process loss’ (groups falling short of their potential performance (Kerr and Tindale, 2004)).
+For some participants, the benefits of the explanations’ modular structure turned into disadvantages when it hampered them in navigating and retrieving information.
+Such impeded interactions included cumbersome information uptake, being overwhelmed by information, and relying on intuition over information.
+Further, for some participants, the group setting contributed to these impediments.
+For example, participant B3 stated that “For me, it doesn’t make sense to […] split up [the explanations], and everyone reads a part, that’s actually not enough.” As before, these impediments can be connected to cognitive mechanisms of ‘collaborative failure’ (Nokes-Malach et al., 2015).
+When groups had difficulties in interacting with the explanations, they also incurred ‘memory coordination cost’ (increased cognitive load) and ‘retrieval strategy disruption’ (losing train of thought).
+We illustrate these mechanisms with an exchange from Group G.
+Although this group was composed of participants with university education, it did not succeed in calculating the employment chance for study task 2, in contrast to single interview participants with the same education.
+G4 We can go through the features briefly. Where is the piece of paper with this terrible matrix? […]
+G2 I still don’t understand which value to put. To calculate it, I need an exact value for the weighting.
+G1 You can calculate it with this. The apprenticeship has 52%. I believe that he [Harald] has over 25%.
+G3 Yes, definitely, I mean, roughly speaking…
+G1 She [Shifteh] has over 30%. And she also has 2 minuses […] and a plus.
+G3 That’s also how I estimated it. […]
+G2 But how do you calculate it? […] And why are there differences between the general weighting and the exact calculation? That doesn’t click for me right now.
+Here, the levels of detail in the explanations acted against participants’ understanding by obscuring the actual feature weights, which were only accessible in level 2 of system details.
+While improved navigation might solve this issue, it also shows the difficulty of simplifying information about AI systems without omitting key aspects.
+In the intention to provide an easier reading of the feature weights, which was an advantage in other cases, the explanations’ clarity was reduced, and information was obscured.
+Addressing all information needs of AI novices (Schmude et al., 2024) thus leads to problems with information overload, as observed in previous work on ‘white-box’ explanations (Cheng et al., 2019).
+However, impeded understanding alone did not mean that collaboration failed; rather, it depended on how groups dealt with these issues.
+Here, the key aspects were group cohesion (Kerr and Tindale, 2004) and constructiveness (Niculae and Danescu-Niculescu-Mizil, 2016).
+When the group dynamic was unfamiliar, it gave room to negative social mechanisms, such as ‘social loafing’ (group loses motivation) and ‘fear of evaluation’ (being criticized by others), and participants began to abandon understanding.
+This suggests that when interaction between participants stopped, interaction with the explanations stopped as well.
+Limiting these negative dynamics and promoting positive ones must be a goal of both explanation design and setting.
+Notably, these adverse social dynamics occurred most often in Groups E and F, which were composed of job-seekers.
+Participants had trouble engaging with the explanations and abandoned interactions and understanding by saying: “Probably [you can solve it] with that, but I don’t know, I’m too stupid for that.” (E3) or “I don’t know what I should say.
+Everything has already been said.” (F2).
+Here, two things failed: The explanations failed to make crucial information accessible, and the group setting failed to uplift members who were discouraged.
+Interactions that offset this discouragement, such as locating information together and outsourcing, were not realized in Groups E and F.
+We thus propose to use co-design approaches to make explanations viable for decision subjects, as has been done with public servants (Weitz et al., 2024).
+Further, XAI should employ methods that create a productive social dynamic, which we identify as the second key aspect to support ‘collaborative success’ and shared understanding (Figure 5).
+4.1.3. Why individuals performed better in the study tasks but still felt the absence of collaboration.
+Despite the different advantages and disadvantages that group and individual settings offer in learning contexts (Section 2.3.1), these differences are rarely examined empirically in XAI888In one study that made this comparison, participants met in a chatroom and decided in collaboration with the AI and not about it (Chiang et al., 2023).
+We address this gap by comparing in-person focus groups and single interviews to examine whether the social setting impacted participants’ understanding.
+We use a triangulation999Using multiple methods of data collection (here: self-reports, task performance, articulations/actions) to explore a phenomenon (Carter et al., 2014).
+approach by investigating participants’ understanding with respect to three aspects: interactions with the explanations, task performance, and self-reported understanding.
+Overall, the explanation afforded its core functionality to both individuals and groups.
+Interestingly, participants in single interviews tended to request the same number or more explanations than focus groups (Figure 6), despite having seemingly less working memory.
+S5 described: “Well, maybe [it was overwhelming] at the very beginning […] But I then realized that I could get through it to some extent.” Single interview participants also regularly consumed explanations in bulk, i.e., read through the whole of an explanation category rapidly.
+This interaction was nearly nonexistent in focus groups.
+However, even though participants in single interviews performed better in the study tasks, they often stated that they missed the “exchange with people, with other perspectives” (S3).
+S8 explained that this exchange would allow for a different form of understanding:
+I think that, on your own, you can think about it very intensely and […] make up your own mind.
+But that’s also the disadvantage, making up your own mind.
+Others may have completely different thoughts and a different professional background.
+And that would probably have been an exciting exchange.
+(S8)
+Participants in individual settings further performed better in the study tasks than groups (Table 3).
+A possible explanation is that participants in single interviews engaged differently with the study tasks, as they often calculated the exact employment chance in study task 2.
+None of the focus groups completed this step, but rather made educated guesses.
+This might be explained by the degree of focused attention the settings afforded participants.
+As single interviews incurred no distractions, participants could immerse themselves in the explanations.
+Notably, this difference is not represented in self-reported understanding.
+Most participants in both settings reported unchanged understanding after the explanation phase (Table 3).
+Paradoxically, they verbally stated that it improved.
+E3 commented: “I don’t think I understood it the way you can understand it yet, but it’s definitely better than before.” And S3 explained that: “I would still say my understanding is ‘good’, but this ‘good understanding’ is much more informed now than the first superficial one.” This indicates that participants tended to judge their understanding relative to the information available, not necessarily in relation to their previous report.
+We describe this process as calibrating understanding.
+Previous research in cognitive science has documented similar effects (Keil, 2006), which were also reproduced in an XAI study on white-box explanations (Cheng et al., 2019).
+However, participants’ verbal reports, their feedback on the explanations (Section 4.1.4), and the calibration process itself indicate that the explanations improved understanding.
+Including additional measures, such as information gain, could further capture the calibration process, which is discussed in Section 5.2.
+To compare individual and group interactions with the explanations, we lastly draw from the ‘six facets of understanding’ (Wiggins and McTighe, 2005).
+The framework describes that understanding is represented by the ability to ‘explain, interpret, apply, take perspective, empathize, and self-reflect’ with respect to a topic.
+The more facets are covered, the better the understanding.
+Seeing that individuals had a clear advantage in solving the study tasks suggests that the individual setting supported the ‘apply’ facet.
+In contrast, the group settings often led participants to ‘explain’ information to others and ‘interpret’ it (Section 4.1.1), and to ‘take perspective’ and ‘empathize’ through the exchange of views and experiences (Section 4.2).
+As explanations aim to improve understanding of a given AI system, combining both settings to cover more facets of understanding could thus be a fruitful approach.
+Further, explanations for individuals in particular can benefit from information covering facets usually dependent on social interaction.
+Our design aimed to implement this through explanations such as What chance would the system calculate for me?
+(interpret – making it personally relevant), and How could the system make the work of counselors easier or more difficult?
+(take perspective: provide multiple angles and arguments).
+Based on these findings, we argue that both group and individual settings can contribute to participant understanding and should ideally be combined.
+In particular, focused attention can facilitate the application of information, while shared understanding and the exchange of opinions and arguments (Section 4.2) aid encouragement, reflection, and collective action.
+Considering this trade-off between settings can inform how explanations can be combined with social settings to cover as many understanding facets as possible.
+Focus groups | |||||||||
+ID | Und. I | Und. II | Change | Task Performance | ID | Und. I | Und. II | Change | Task Performance |
+A1 | 2 | 2 | 0 | F1 | 2 | 4 | +2 | ||
+A2 | 1 | 2 | +1 | F2 | 1 | 4 | +3 | ||
+A3 | 4 | 4 | 0 | F3 | 5 | 2 | -3 | ||
+A4 | 4 | 4 | 0 | F4 | 3 | 3 | 0 | ||
+B1 | 4 | 2 | -2 | F5 | 4 | 4 | 0 | ||
+B2 | 4 | 2 | -2 | G1 | 5 | 5 | 0 | ||
+B3 | 5 | 5 | 0 | G2 | 3 | 2 | -1 | ||
+B4 | 2 | 3 | +1 | G3 | 5 | 2 | -3 | ||
+C1 | 4 | 4 | 0 | G4 | 5 | 4 | -1 | ||
+C2 | 2 | 2 | 0 | H1 | 4 | 5 | +1 | ||
+C3 | 1 | 3 | +2 | H2 | 4 | 5 | +1 | ||
+D1 | 4 | 3 | -1 | H3 | 4 | 4 | 0 | ||
+D2 | 4 | 3 | -1 | H4 | 4 | 5 | +1 | ||
+D3 | 4 | 4 | 0 | H5 | 4 | 4 | 0 | ||
+E1 | 3 | 4 | +1 | ||||||
+E2 | 3 | 3 | 0 | ||||||
+E3 | 4 | 4 | 0 | ||||||
+Single interviews | |||||||||
+ID | Und. I | Und. II | Change | Task Performance | ID | Und. I | Und. II | Change | Task Performance |
+S1 | 5 | 5 | 0 | S7 | 4 | 4 | 0 | ||
+S2 | 4 | 2 | -2 | S8 | 3 | 4 | +1 | ||
+S3 | 4 | 4 | 0 | S9 | 5 | 5 | 0 | ||
+S4 | 4 | 4 | 0 | S10 | 4 | 3 | -1 | ||
+S5 | 5 | 4 | -1 | S11 | 4 | 4 | 0 | ||
+S6 | 1 | 4 | +3 | S12 | 2 | 4 | +2 |
+4.1.4. Reflections on the explanation design: Modularity, levels of detail, and most important information.
+To examine how the explanation design was received, we asked participants for feedback on the explanations’ structure, content, style of expression, and information coverage.
+We report and summarize the participants’ criticisms as a basis to formulate design improvement suggestions in Section 5.
+Strengths and weaknesses of the design.
+Positive comments described the explanations’ structure as “nicely presented” (A2, C2) and “good to get an overview” (C3, H4) while being “active and controllable” (S8).
+Critical comments described the information coverage as “too much” (D1, S4), and the structure as “confusing” (B1, D1) and “demanding” (D2).
+Participants saw the design’s strengths in its four-category structure, question-driven presentation, active selection, and information scope.
+However, the scope and depth of information also led to information overload and loss of overview.
+Further, the explanations’ many and complicated texts were described as “very difficult” (E2, F2).
+E2 compared the language to “letters […] from the court.
+I understand every single word, but I don’t understand the context.” Previous work has found that textual explanations can effectively convey information but tend to raise aversion with users (Szymanski et al., 2021; Schmude et al., 2023).
+However, Weitz et al.
+(2024) paradoxically found that users preferred textual explanations.
+This points to the need for further research on textual formats in XAI, like the automated adaptation of text to different difficulties.
+Most helpful and influential information.
+Participants in focus groups stated that all explanation categories helped their understanding and influenced their decision evenly ( Figure 7), often mentioning that “all of them [are relevant]… I don’t think you can leave anything out, really” (D3).
+In contrast, participants in single interviews found data much less helpful and less influential, stating, e.g., that they prioritized another category in the time available.
+Notably, participants emphasized that two categories were central: system details and context.
+System details were perceived as “tangible” (S6) and “concrete” (S8), and explanations about the features and weighting were perceived as especially important: “That is the central point, the basis of the whole system.” (G4) In turn, explanations from the category context were requested the most (Figure 6).
+Here, participants appreciated explanations that described decision subjects’ inability to contest decisions and the system’s political background.
+Drawing from the concept of ‘intelligibility types’ (Lim and Dey, 2009), we argue that system details provided descriptive information to the question “What did the system do?”, while context provided normative information to the question “Why did the system do [this]?”.
+Future research should investigate how both information types can be integrated into explanations for AI novices.
+4.1.5. Summary RQ1: Group and individual settings address different understanding facets in explainable AI
+In RQ1-Explanations, we asked how a question-driven, modular explanation design supports AI novices’ understanding in groups and individual settings.
+We found that the explanations support both settings but differ in how understanding develops.
+In groups, we found that explanations facilitated interactions that produced shared understanding and involved cognitive and social mechanisms of ‘collaborative success’ (Nokes-Malach et al., 2015).
+In groups with a trusting social dynamic, participants tackled explanations together, which acted against discouragement.
+When groups had a negative social dynamic, the explanation design could become overwhelming and understanding issues were left unchecked, leading to ‘process loss’ (Kerr and Tindale, 2004) and abandoned understanding.
+Participants in single interviews interacted with the explanations in a more focused and self-directed manner.
+This had advantages for task performance and explanation engagement, which aligns with research on tutoring methods (Bloom, 1984).
+However, the positive effects of aggregated knowledge (Navajas et al., 2018) and peer discussion (Smith et al., 2009) in group settings should not be disregarded.
+Our findings showed that group settings can bridge understanding issues by boosting morale and letting participants share knowledge, interpretations, and experiences.
+We thus argue that individual and group settings support different understanding facets (Wiggins and McTighe, 2005), meaning that they provide different grounds for understanding AI systems.
+While individual settings can make it easier to understand technical and numerical details that require much attention (‘apply’), group settings can support understanding of the deployment context and consequences through the exchange of expertise and lived experiences (‘interpret’, ‘take perspective’, ‘empathize’).
+Consequently, individual and group settings should be combined to leverage their different modes of interaction and understanding facets when explaining AI systems.
+In cases where both social settings cannot be provided, explanations of AI systems should aim to reinforce facets that are not covered in the corresponding setting.
+4.2. RQ2-Deliberation: How do AI novices use explanations to form opinions and make decisions about AI systems in groups and individual settings?
+Before and after the explanation phase, participants decided if the study’s case should be deployed, and groups additionally made a collective decision (Table 4).
+We compared participants’ decisions and decision confidence and how they deliberated deployment to examine the impact of explanations and the social setting.
+We first describe participants’ confidence and decision changes in single interviews (4.2.1) and then present three cases of group discussion illustrating ‘elements of deliberation’ (Stromer-Galley, 2007), including reasoned arguments (4.2.2), disagreement (4.2.3), and groupthink (4.2.4).
+Section 4.2.5 summarizes the findings.
+4.2.1. Explanation phase led to increased decision confidence and decision swings.
+For most participants, deciding about the AMS algorithm’s deployment was a clear choice: 7 out of 8 groups and the majority of single participants voted “No” (Table 4).
+Many participants reported increased decision confidence after the explanation phase and stated that they felt better informed due to the explanations and, where applicable, the group discussion.
+Reasons for these increases included a better understanding of the system’s “fundamental idea” (S4) and the “exchange of different opinions and things that catch your eye” (G2).
+Participant B3 emphasized that the explanations, although they only contained factual information, provided a stark contrast to public narratives:
+Well, I changed my mind – you think you understand something when you see it in the media.
+You have a political opinion about it.
+But you don’t know the background information.
+And when you get to the background information, you can have a completely different opinion.
+(B3)
+This contrast highlights how explanations of an AI system can impact decision-making by correcting lay understandings (DeVito et al., 2018) and is in line with previous work that advocates for making algorithmic complexity visible to users to tackle lay understandings (Eslami et al., 2016).
+It further directly connects the explanations and changes in participants’ decision confidence.
+Few participants reported decreased decision confidence, and only S3 and S11 reported a stronger decrease of -2 (Table 4).
+S3 explained that the system might have benefits, but too much depended on the conditions for deployment.
+In particular, it should be deployed “responsibly, with a pilot project, in a selected group, for three months”, and not haphazardly, where “you sit around for a day or eight hours and then training is finished” (S3).
+S11, who together with S3 requested the most explanations out of all participants, paradoxically stated that their confidence decreased because “I’m still missing so much information.
+Especially […] how tedious it is for the counselors if they have to disagree with the system.” In consequence, S11 changed their decision from ‘Yes’ to ‘No’.
+Notably, this fear of an algorithmic imprint was a prevalent theme throughout all studies and was often connected to past experiences with digitization projects, and the corresponding institutional deficiencies.
+S5, who also changed from ‘Yes’ to ‘No’, similarly stated that the explanations helped them to scrutinize the system: “You don’t have to introduce anything that’s extra bad”.
+While the explanations thus made the decision more uncertain for some, they undoubtedly encouraged critical reflection about the use case and triggered decision changes.
+Despite only having “their own mind” (S8), single participants could make use of the explanations to weigh pros and cons and adjust their mental model.
+This form of “internal deliberation” is supported by exposure to different views, as provided through the argumentative explanations in usage and context.
+This suggests that explanations can substitute at least small parts of public deliberation, which is thought to be the more salient driver for “reasoning towards good outcomes” (Mercier and Landemore, 2012).
+To contrast these findings with focus groups, we illustrate this public deliberation with three conversation excerpts that showcase elements of deliberation in focus groups.
+Focus groups | |||||||||
+ID | Decision I | Group decision | Decision II | Decision Conf. | ID | Decision I | Group decision | Decision II | Decision Conf. |
+A1 | No | Yes | 0 | F1 | Yes | Yes | 0 | ||
+A2 | No | Yes | +1 | F2 | Yes | Yes | -1 | ||
+A3 | Yes | Yes | +1 | F3 | Yes | No | +1 | ||
+A4 | No | Yes | Yes | 0 | F4 | No | No | -1 | |
+B1 | No | No | +3 | F5 | Yes | No | Yes | -1 | |
+B2 | Yes | No | 0 | G1 | Yes | Yes | +1 | ||
+B3 | Yes | No | +2 | G2 | No | No | 0 | ||
+B4 | Yes | No | No | +3 | G3 | Yes | No | 0 | |
+C1 | No | No | -1 | G4 | No | No | No | +1 | |
+C2 | No | No | 0 | H1 | No | No | +1 | ||
+C3 | No | No | Yes | +2 | H2 | No | No | +1 | |
+D1 | No | No | +1 | H3 | No | No | +3 | ||
+D2 | No | No | 0 | H4 | No | No | 0 | ||
+D3 | No | No | Yes | 0 | H5 | No | No | No | +1 |
+E1 | No | No | 0 | ||||||
+E2 | No | No | 0 | ||||||
+E3 | No | No | Yes | -1 | |||||
+Single interviews | |||||||||
+ID | Decision I | – | Decision II | Decision Conf. | ID | Decision I | – | Decision II | Decision Conf. |
+S1 | Yes | – | Yes | 0 | S7 | No | – | No | +3 |
+S2 | No | – | No | -1 | S8 | No | – | No | 0 |
+S3 | No | – | No | -2 | S9 | No | – | No | +1 |
+S4 | No | – | No | +1 | S10 | Yes | – | Yes | 0 |
+S5 | Yes | – | No | +2 | S11 | Yes | – | No | -2 |
+S6 | No | – | No | +2 | S12 | No | – | No | +1 |
+4.2.2. Case 1 - reasoned arguments: Group B discusses whether to deploy the system.
 Group B was composed of staff members and volunteers of a civil society organization.
-Three participants in this group changed their votes from ”Yes” in the first report to ”No” in the second report.
-In the excerpt below, B2 gives their perspective on the system’s deployment, and B3 responds with theirs, referring to explanations about the system’s features and weightings (System details, Base, B + Level 2, C), which fuel the discussion that follows:
-B2: I’m skeptical, but I’m still in favor of introducing it. Because it could be an aid and a relief for the staff working there.
-B3: I was originally in favor for these reasons, but since I’ve seen these parameters, I would be very much against it.
+Three participants in this group changed their votes from ‘Yes’ in the first report to ‘No’ in the second report.
+We found this change to be driven by three main deliberation elements (Stromer-Galley, 2007): ‘sourcing’ information, ‘reasoned arguments’ (opinion claims grounded in the information), and ‘engagement’ with the topic and between participants.
+In the excerpt, B2 and B3 weigh pros and cons of deployment.
+B3 grounds their arguments in explanations about the system’s features and weightings (system details), changing the discussion’s course:
+B2 I’m skeptical, but I’m still in favor of introducing it. Because it could be an aid and a relief for the staff working there.
+B3 I was originally in favor for these reasons, but since I’ve seen these parameters, I would be very much against it.
 Because I think there’s a lot of ideology in it.
-I think it’s no longer acceptable that men are favored over women and that duty of care only applies to women – this comes from a time that should be long gone.
-B2: Those are strong arguments.
-B3: The things that come out are so absurd as well.
-For example, the apprenticeship of our Harald [the example job-seeker] was rated positively.
-He can’t use the apprenticeship at all for retraining.
-These things are given far too little consideration.
+I think it’s no longer acceptable that men are favored over women and that duty of care only applies to women.
+This comes from a time that should be long gone.
+B2 Those are strong arguments.
+B3 The things that come out are so absurd as well.
+For example, Harald’s apprenticeship was rated positively, but he can’t even use the apprenticeship for retraining.
 […] As much as I like the idea, I don’t like the parameters.
-B1: Did you vote yes first?
-B3: I ticked yes at first, but I was really shocked at what was in there. […]
-B1: What I’m wondering is, what would be the real benefit of introducing the system?
-What could be found out with much less effort?
-After all, employees should be able to see which degree you have quite quickly.
-B4: It’s a grid, a structure for the people who work at the agency so that they can quickly find a box.
-The excerpt highlights how the provided information changed B3’s attitude toward the system and how this change, in turn, affected the group discussion.
-The combination of explanations and group discussion here had a major effect on participants’ perceptions, as three participants in this focus group changed their vote from ”yes” in the first self-report to ”no” in the second.
-Interactions like B2’s appreciation of B3’s arguments and B1’s questions about the system’s benefits depict processes of constructive group deliberation directly supported by the explanations.
-Case 2: Group debates the meaning of ”unalterable” facts.
-Group D was composed of participants who were neither job-seekers nor domain experts and thus would be affected only indirectly by the system’s deployment.
-Even so, when D1 referred to an explanation describing how two different job-seekers would be judged by the system (System details, Base, C), it led to a discussion about the objectivity and adequacy of using data to represent a person:
-D3: I believe that the system can form the initial basis, based on the unalterable facts, which are of course weighted, but then it has to be enriched by a human being. […]
-D1: But I don’t believe that there are unalterable facts – well, not in this area. It’s all a question of representation and the lens through which you see the world.
-D3: When the job-seeker says, I only have four years of elementary school, then that’s four years of elementary school… […]
-D2: That doesn’t mean that he can’t still be a very educated person.
-D3: But that is hard to sell to an employer, right? […]
-D2: I’m skeptical about the data. You [D3] said the ”basis”, I think there are cracks in this basis. And I’m afraid of these cracks, that something will be pre-determined…
-D3: But the human decision is always subjective.
-D2: That has to be weighed up. On the one hand, you have the arbitrariness of the individual employee, yes, and on the other hand you have an incomplete picture of a person.
-D1: Or a false image.
-D2: An incomplete one, I would say.
-In this excerpt, the discussion surfaces discrepancies in how participants perceived the factuality and value of data.
-While the discussion in this focus group did not lead to a consensual decision in the time given, we still see this exchange as an important process in the deliberation on public AI systems.
-The merit of this discussion was later acknowledged by D1, who found the explanations confusing but stated that these exchanges were the ”centerpiece” and the most intriguing part.
-This highlights the role of explanations in the surfacing and debating of divergent opinions.
-Case 3: Group discusses consequences of the system’s deployment for affected people.
-Group H was composed of staff members of a civil society organization specializing in helping young adults to take up apprenticeships.
-While all participants were clearly against the system’s deployment, the group decision task still led to a discussion about why specifically they opposed the system.
-The excerpt begins with H2’s comment about an explanation describing that job-seekers do not have a right of contestation (Context, Level 3, B):
-H2: Just the fact that you can’t contest the decision, that’s a no-go.
-H1: I think that having such systems introduced by institutions that themselves have very little idea of the real work – they may even want to help and support work systems, but that completely misses the mark because they are making decisions that are again driven by economic interests.
-H4: And the life of the people who are then affected by the system is not represented at all, the psychosocial situation of the people is not taken into account.
-H3: It is if they are depressed…
-H4: Then it’s minus, minus, exactly.
-It simply deprives people of the opportunities they deserve.
-And that’s probably also my personal experience, I’ve worked with long-term unemployed adults who also had physical or mental impairments.
-The problem is that the longer I’m out of the job market, the fewer opportunities I have and people are already in a bad enough situation.
-And then you’re totally stuck in this [algorithmic] system.
-I just think that’s unfair.
-H2: In terms of labor market policy, it may be justified, but in terms of social policy, it’s totally wrong.
-The excerpt shows that even participants with unanimous attitudes toward the system found value in discussing and sharing their reasons for disapproving its deployment.
-H4 in particular appeared to be incensed by the system’s representation of soft factors.
-Their account of working with long-term job-seekers was later appreciated by the other participants, who stated that this perspective was valuable to them.
-The explanations in this case served to facilitate group conversation about the meaning and consequences of the system’s deployment and encouraged participants to share their views, stances, and experiences.
-Summary.
-While Section 4.2.1 describes that explanations were effective in supporting group collaboration, this Section shows that their role in providing information for discussions between participants is equally important.
-Explanations helped participants in debating topics that are pivotal to the acceptance, trust, and application of ADM systems, in effect supporting their collaborative reasoning (Moshman and Geil, 1998).
-In this way, the explanations support what Kawakami et al.
-(2024) describe as ”deliberations around whether and under what conditions to move forward with developing or deploying” a public sector AI tool.
-This deliberation is important, as it allows participants to express and reflect on their opinions, feel heard, and consider other perspectives when making decisions.
+B1 Did you vote yes first?
+B3 I ticked yes at first, but I was really shocked at what was in there [in the system]. […]
+B1 What I’m wondering is, what would be the real benefit of introducing the system? […]
+B4 It’s a grid, a structure for the people who work at the agency, so that they can quickly find a box.
+The excerpt highlights how the explanations led B3 to change their deployment decision and served as discussion triggers.
+In the resulting discussion, participants state both arguments (discrimination, what’s the benefit?) and opinions (disagreement with policy choices, AI can assist in decisions).
+Note that there is a difference between arguments (expression of reasoning processes that can be defended against critique) and opinions (expression of the speaker’s belief) (Stromer-Galley, 2007; Mercier and Landemore, 2012).
+While conceiving arguments to persuade interlocutors can result in confirmation bias (interpreting evidence such that it confirms existing beliefs) (Mercier and Sperber, 2011), the fact that B3 changed their attitude, in fact, indicates that the explanations acted against this bias.
+We argue that the excerpt thus shows a positive synergy in that the explanations provided grounds for ‘arguments’, which entered the discussion and provoked ‘collective reasoning’ and three decision swings.
+However, considering the large argumentative influence of B3, it should also be considered how the discussion would unfold if B3 had advocated for deployment.
+A case with comparable dynamics is described in 4.2.4.
+4.2.3. Case 2 - disagreement: Group D debates normative positions regarding the algorithmic representation of people.
+Group D was composed of participants who had been job-seeking in the past.
+When the group discussed the AI system’s deployment, the conversation shifted to how features that represent job-seekers’ profiles should be selected and weighed.
+This produced disagreement, an “important marker for deliberation” (Stromer-Galley, 2007) that displays heterogeneity of viewpoints, acts against polarization, and involves close examination of others’ reasoning.
+In the excerpt, D3 argues for the system’s deployment, while D1 argues against, and D2 acts as a mediator:
+D3 I believe that the system can form the initial basis, based on the unalterable facts, which are of course weighted, but then it has to be enriched by a human being. […]
+D1 But I don’t believe that there are unalterable facts – well, not in this area. It’s all a question of representation and the lens through which you see the world.
+D3 When the job-seeker says, ‘I only have four years of elementary school’, then that’s four years of elementary school… […]
+D2 That doesn’t mean that he can’t still be a very educated person.
+D3 But that is hard to sell to an employer, right? […]
+D2 I’m skeptical about the data. You [D3] said it’s the ‘basis’, I think there are cracks in this basis. And I’m afraid […] that something will be pre-determined…
+D3 But the human decision is always subjective.
+D2 That has to be weighed up. On the one hand, you have the arbitrariness of the individual employee, yes, and on the other hand, you have an incomplete picture of a person.
+D1 Or a false image.
+D2 An incomplete one, I would say.
+The group here discusses diverging views and expresses opinions.
+While these opinions are meant to persuade and defend, they are grounded in lived experiences rather than in the explanations.
+The central conflict develops between D1’s belief that the system misrepresents reality and D3’s viewpoint that it can increase objectivity and assist in decisions.
+The discussion here did not lead to a consensus on the deployment decision in the given time, but resulted in a majority vote for ‘No’.
+We argue that it still illustrates an important process in the deliberation on public AI systems: Participants again ‘sourced’ information that was turned into arguments, but the debate led to a more fundamental topic that surfaced discrepancies which would impede finding a collective decision.
+The fact that participants then engaged in ‘disagreement’ is a sign of productive deliberation, as it shows that there were diverse viewpoints, that no polarization or ‘groupthink’ (Janis, 1971) occurred, and that the proposal was closely examined based on the information given (Stromer-Galley, 2007).
+In a real setting, this form of debate could serve as a fruitful basis to investigate whether the system is in the ‘public interest’ (Züger and Asghari, 2023) and to host ‘early-stage deliberations’ (Kawakami et al., 2024) on the system during development.
+The merit of this debate was further later acknowledged by D1, who found the explanations confusing but stated that these exchanges were the study’s “centerpiece” and most intriguing part.
+We argue that the interplay between explanations and group discussion here supported a (simulated) evidence-informed policy-making process (Mair et al., 2019).
+4.2.4. Case 3 - groupthink? Group A follows a minority position and votes for system deployment
+Group A was composed of volunteers from a civil society organization.
+Three participants in this group changed their decisions, shifting from ‘No’ to ‘Yes’ after the explanation phase.
+We explain these changes with three aspects: First, Group A focused on the explanation category data and did not interact much with other categories (Figure 6).
+This meant less attention was paid, for example, to the system’s feature selection and weightings that were decisive in Cases 1 and 2.
+Second, participants of Group A stated that they were not directly affected by the system, as they were retired, implying low ‘engagement’: “It doesn’t affect me anymore and I think to myself, yeah…” (A1).
+Third, participants prioritized group concurrence above a “careful, critical scrutiny” (Janis, 1971).
+The following excerpt illustrates the tipping point for the collective decision:
+A3 I still think the system is better, even if there are still mistakes in it, than sitting opposite someone [a counselor] who doesn’t like you… […]
+A2 So rather ‘no’?
+A1 Yes, as A3 says, it’s… I don’t know.
+A3 Yes and no… […]
+A2 I mean, it can’t be avoided, it will happen. I’m convinced of that, whether we like it or not, it’s done.
+A1 It won’t affect us anymore, at least not in the employment office. […] I agree with the majority.
+A2 But that’s difficult now.
+A3 I’ll stick with ‘yes’. My daughter would say I shouldn’t think so negatively, especially with AI. […]
+A2 I say ‘yes’ too. […] You, A1 and A4, can tip the scales.
+A1 I say ‘yes’ now too, but not because I’ve changed my mind, but because I want an overall solution.
+A4 I say ‘yes’ but I’m leaning towards ‘no’.
+Despite articulated reservations, all participants ultimately decided to vote for deployment.
+We compared this excerpt with characteristics of ‘groupthink’, a “mode of thinking” in which people value concurrence higher than consideration of alternative courses of action (Janis, 1971).
+This mode produces defective decision-making processes due to three key aspects: strong social identification with the group, salient norms, and a perceived low self-efficacy to make the decision (Baron, 2005).
+The excerpt clearly demonstrates two of these aspects: A1 changes their decision due to a desire for group harmony, and A4 follows suit (group identification).
+Further, both A1 and A3 express their uncertainty and sway between options (low self-efficacy).
+While A2’s statement that AI is inevitable is an opinion rather than an argument (neither ‘sourced’ nor the product of evident reasoning), it triggers the group to make a quick decision that disregards any remaining ‘disagreement’.
+Although the process can be connected to aspects of ‘groupthink’ (Janis, 1971), such as rationalizations of flawed logic and self-censorship, these aspects are not nearly as pronounced as in the literature (Janis, 1971, 1972; Baron, 2005).
+For example, the group did not share an illusion of unanimity, and the uncertainty among participants suggests no guiding salient norms.
+Still, as participants avoided ‘disagreement’ and instead followed decisions of other, the excerpt presents a suboptimal deliberation process (Baron, 2005).
+In part, this can be attributed to the explanation’s failure to make all fundamental information easily available and to not encourage analytical thinking over intuitive, heuristical thinking (Buçinca et al., 2021).
+In addition, the group might have missed a role that explicitly takes the opposing viewpoint to fuel discussion, which was identified to benefit deliberation in previous XAI research (Chiang et al., 2024).
+The implications of these findings are discussed in Section 5.
+4.2.5. Summary RQ2
+The findings in this Section demonstrate how explanations supported deliberation in focus groups and single interviews.
+Many participants reported improved decision confidence and changed their deployment decisions based on the explanations, often due to a disillusionment regarding the AMS algorithm’s assumed merits.
+These changes occurred in both settings, suggesting that the explanations supported public and internal deliberation (Mercier and Landemore, 2012).
+In group settings, participants used explanations when discussing deployment, as illustrated in Case 1.
+Case 2 further highlights that explanations surfaced discrepancies in personal beliefs and produced productive conflict.
+In contrast, Case 3 shows a deployment decision based more on concurrence-seeking than on ‘collaborative reasoning’ (Moshman and Geil, 1998).
+However, we hesitate to label the exchange as ‘groupthink’, as it does not align with all factors that characterize the phenomenon (Baron, 2005).
+Based on these findings, we argue that explanations can support people in considering if AI systems are in the public interest and to discuss “whether and under what conditions to move forward with developing or deploying” them (Züger and Asghari, 2023).
+To achieve this, both the explanations and the group setting need to i) be designed so that they allow for the easy sourcing of information for arguments, ii) make all relevant information available as soon as possible, and iii) include mechanisms that encourage participants to examine both the proposal and their positions closely.
+Matching explanations and social setting to support ‘elements of deliberation’ (Stromer-Galley, 2007) thus presents promising starting points for future research on how explainable AI can promote public deliberation on AI.
 5. Discussion
-In this section, we discuss the implications of our findings for the design of explanations for AI novices (Section 5.1) as well as suggestions for their applications in collaborative settings (Section 5.2), which are summarized in Figure 7.
-5.1. Reflections on the explanation design
-Explanations improved or calibrated participants’ understanding, but measurements should be reconsidered.
-In Section 4.1.1, we described that while participants gave discrepant responses in their understanding self-reports (Table 3), they rarely felt that their understanding truly decreased or remained the same.
-Instead, we argue that the explanations changed their frame of reference for ”understanding” by revealing blind spots and making participants realize that their initial understanding was worse than assumed.
-Drawing from definitions in the cognitive sciences, the explanations thus supported participants in developing understanding by i) making participants realize that their initial understanding was incomplete (Rozenblit and Keil, 2002) and ii) supporting them in locating and closing relevant understanding gaps (Keil, 2006).
-The effect of participants reporting decreased perceived understanding following the explanations can be attributed to the fact that these explanations led them to attain a ”partial understanding” (Keil, 2019), but also showed them the difference to a complete, comprehensive understanding.
-The difficulty of self-assessing one’s understanding is a known issue (Rozenblit and Keil, 2002) and previous studies in XAI have shown similar effects caused by white-box explanations (Cheng et al., 2019), which, due to their high information density, led to increases in objective understanding but decreases in self-perceived understanding.
-This is not necessarily an issue regarding participants’ understanding, and together with the substantial increases in decision confidence, it can be seen as evidence that our explanations effectively improved and calibrated participants’ understanding.
-However, it also demonstrates that the relativity of internal cognitive states should be accounted for when measuring understanding.
-For this reason, recent work in XAI has proposed measurement based on abilities (Speith et al., 2024).
-This approach appears promising, considering that our single interview participants often found applying their knowledge in the study tasks helpful.
-Still, it likely requires modification when used in collective settings, as, for example, participants in focus groups never calculated the precise employment chance in task 2.
-The operationalization of the abilities-based approach for both individual and group settings is thus a direction for future research.
-Participants considered all four explanation categories relevant and appreciated technical and contextual information.
-Our explanation design covered the categories data, system details, usage, and context.
-In the second self-reports, participants indicated which of these categories were ”helpful for their understanding” and ”influential to their decisions” (Section 3.3).
-In Section 4.1.5, we describe that focus groups reported that all categories helped their understanding but that system details more often than other categories influenced their decisions.
-In contrast, participants in single interviews reported all categories equally often as helpful and influential except data.
-Possible explanations for this effect include the forced prioritization due to the time limit, which led participants in single interviews to skip information about data as it was believed to be less relevant.
-In comparison, some focus groups proceeded chronologically through the explanations and thus spent more time on data as the first category.
-Further, the findings highlight that participants welcomed non-technical information about the system’s intention, ethical values, and practical usage (Figure 6) but also assigned high importance to the model’s exact feature weights.
-We see this as confirmation that participants’ information needs cover the entirety of an ADM system and that explanations for AI novices should aim to provide both technical (data, system details) and contextual (usage, context) information.
-Explanations for decision subjects require a co-design approach.
-In Section 4.1.2, we describe that our explanations were most effective for participants who did not face challenges with either the language, large amounts of text, or the explanations’ structure.
-While these attributes are not entirely congruent with participants’ education levels, we observe that university-educated participants faced comparatively little difficulty interacting with the explanations, as the conversation excerpts in Sections 4.2.1+4.2.2 show.
-The differences were most apparent in Groups E and F, where participants stated that the explanations did not help them fully grasp the system as they required a high language and reading proficiency.
-We see this as evidence that a co-design approach similar to Weitz et al.
-(2024)’s work is required to build explanations for groups composed like Groups E and F, as it might be difficult to identify which content and format of explanations are required by decision subjects.
-On the other hand, the findings demonstrate the explanation design’s suitability for groups of domain experts, who also stated that this format of conveying information would be useful as a training measure.
-Suggestions for improving the interaction design and combining individual and collective settings.
-In Sections 4.1.1 and 4.1.2, we described that according to participants’ feedback, the explanation design offers comprehensive and flexible information selection and encourages self-directed and active exploration, but that it also has a high access threshold and requires adjustment to the subdivided structure.
-For focus group participants, this design resulted in participants splitting and swapping explanations, sharing relevant parts, and discussing surprising information, which overall resembled an associative approach that supported group interaction but also led to a more fragmented picture of the system.
-In single interviews, participants could select and interact with information at their own pace and in a focused manner but, on the other hand, missed exchange and external perspectives.
-These effects can be described in the context of collaborative success and failure (Nokes-Malach et al., 2015), which states that groups require extra cognitive effort to coordinate and manage interactions, which can lead to memory limitations, information retrieval disruption, and blocking of thought processes if these coordination costs are too high.
-In contrast, individuals do not incur these cognitive costs, but of course can also not benefit from collaborative success mechanisms such as increased working memory or complementary knowledge.
-A future iteration of the explanation design could thus follow an individual with a collective interaction phase to combine the advantages and offset limitations.
-Further, a digital version would facilitate overview (summaries) and navigation (menus) while allowing for simple language options and cross-references.
-Lastly, the essential information of each explanation category, which in our design was contained in the base levels (Section 3.2), could use a balance of textual and visual design to ensure that essential information does not rely only on textual understanding.
-On a side note, we emphasize that identifying explanation subtopics and splitting them up into levels of detail is challenging.
+In this section, we discuss how our findings answer our two main research questions: Whether a question-driven, modular explanation design supports AI novices’ understanding in groups and individual settings (RQ1) and how AI novices used these explanations to deliberate about AI systems (RQ2).
+We describe the advantages of both social settings for explainable AI, outline which real-world use cases would benefit from our explanation design, discuss whether the explanations improved understanding, and provide suggestions for their design improvements.
+We summarize the implications of our findings in Figure 8.
+5.1. Do AI novices learn and deliberate about AI better together or individually?
+In Section 4.1, we described that explanations produced shared understanding in groups, involving both cognitive and social mechanisms of “collaborative success” (Nokes-Malach et al., 2015).
+Section 4.2 further showed that explanations improved participants’ decision confidence and provided grounds for different elements of deliberation (Stromer-Galley, 2007), such as reasoned arguments and disagreement.
+In the best cases, focus groups in our study had a familiar (Johnson and Johnson, 1985) and solution-oriented (Niculae and Danescu-Niculescu-Mizil, 2016) atmosphere that facilitated sharing and discussing information.
+In these settings, the modular explanation structure showed its strengths by allowing for the distribution of tasks among group members, providing high levels of detail and breadth if needed, and offering different viewpoints that could be used as argumentative and conversational starting points.
+In this sense, the explanations fulfilled their aim of supporting learning and deliberation about a public AI system (Kawakami et al., 2024).
+The interaction between group members is the differentiating factor compared to “one-to-one” (Naiseh et al., 2021) explanation settings.
+In our study, single interviews allowed for more focused engagement with explanations and a form of “internal deliberation” (Mercier and Landemore, 2012) but lacked the exchange of knowledge and perspectives with others that is deemed central for deliberation about public AI (Züger and Asghari, 2023).
+Regarding learning and deliberation, XAI would thus benefit from researching how group settings can be used to leverage collective reasoning (Moshman and Geil, 1998), wisdom of the crowds (Navajas et al., 2018), and performance increases through peer discussion (Smith et al., 2009).
+However, the benefits of group settings have several preconditions, such as the containment of cognitive biases (groupthink (Janis, 1971), equality bias (Naiseh et al., 2024)) and, crucially, a trusting social dynamic (Chiang et al., 2023).
+The importance of the social dynamic became evident in groups where members were not familiar and had trouble engaging with the explanations.
+In groups G and H, for example, the social dynamic bridged understanding issues of individual participants and acted against discouragement.
+In groups E and F, in contrast, these understanding issues eventually led participants to abandon understanding, as the social atmosphere did not support them in overcoming them.
+Here, a lack of trust or simply unfamiliarity between participants likely amplified effects such as social loafing and the fear of being evaluated (Nokes-Malach et al., 2015).
+This underscores the importance of creating trust between group members in collaborative XAI settings (Johnson and Johnson, 1985).
+Intuitive measures could be the introduction of a simple task that the group solves collaboratively before engaging with explanations, such as the Wason card selection task (Wason, 1968).
+Another measure could be the introduction of roles (e.g., proponents and opposition), as has been done with the “devil’s advocate” in previous work (Chiang et al., 2024), to facilitate discussion and close examination of the proposal.
+Future work should examine how such measures can be incorporated into explanation design to support interaction in groups of comparable compositions.
+Lastly, regarding cognitive biases, we observed an effect resembling some aspects of “groupthink” (Janis, 1971) when participants in Group A changed their vote to “Yes” to reach a group decision.
+We argue that this effect originated in the lack of detailed interaction with explanations and, possibly, a perceived low degree of personal affection by the system’s deployment.
+However, this is contrasted by participants in Group D, who debated at length about the system’s deployment without reaching a consensus, despite not being directly affected.
+Potential measures to avoid groupthink in discussion could thus be to encourage debate, which again could be the introduction of roles to improve the “dialectic argumentation” (Mercier and Sperber, 2011), and to explain the system in a way that makes it more personally relevant to participants (Wiggins and McTighe, 2005), e.g., by emphasizing connections to their own experiences.
+5.2. Did the explanations improve participants’ understanding?
+In Section 4.1.5, we described that the explanations helped participants develop different ‘facets of understanding’ (Wiggins and McTighe, 2005).
+In groups, participants were encouraged to ‘explain’ information to each other and ‘empathize’ with others’ experiences, while individuals could better ‘apply’ information in the study tasks.
+We further described that groups’ interactions with explanations realized mechanisms of “collaborative success” (Nokes-Malach et al., 2015).
+We thus conclude that the explanations had a positive effect on understanding.
+However, a more complete answer requires that we consider the difference between measurement methods and true cognitive states.
+In Section 4.1, we described that a majority of participants reported unchanged understanding after the explanation phase (Table 3) but, paradoxically, described verbally that their understanding improved; two seemingly incongruent pieces of evidence.
+We explain these contradictory findings with a process we call calibrating understanding.
+The term describes that participants tend to report understanding not in absolute terms, or even in relation to past understanding, but in relation to the currently available information.
+Participants explicitly stated that they calibrated their interpretation of ’good understanding’ according to their knowledge of the information basis, which differed before and after the explanation phase.
+The calibration process can be traced by using concepts from the cognitive sciences: Participants i) reported their initial understanding after reading the use case introduction, they then ii) saw the explanations and realized that they had understanding gaps (Rozenblit and Keil, 2002), which they iii) proceeded to locate and close (Keil, 2006), however, they iv) also realized that they could not look at every available explanation and would develop at most a “partial understanding” (Keil, 2019), which they v) rated accordingly in the second self-reports.
+Previous studies in XAI documented similar effects caused by white-box explanations (Cheng et al., 2019), which, due to their high information density, led to increased ‘objective’ understanding but decreased self-reported understanding.
+Importantly, (Papenmeier et al., 2022) found similar discrepancies when participants who received no explanation gave higher understanding scores than participants who received faithful explanations; a discrepancy our findings might explain.
+In the same study, similar discrepancies also occurred between trust self-reports and observations of behavior (Papenmeier et al., 2022).
+In line with these findings, we argue that the calibration effect should be accounted for when measuring understanding, for example, by eliciting an additional metric that captures the perceived scope of available information.
+A potential reporting question could be “How much information do you feel you currently have about the presented AI system?”, combined with a 5-point scale ranging from “very little” to “very much”.
+Self-reported understanding could then be compared with self-reported information scope and verbal responses to acquire a more complete picture.
+Recent work in XAI has further proposed understanding measurement based on participants’ abilities (Speith et al., 2024).
+This approach appears promising, as the ability to calculate study task 2 was a relevant metric in our study.
+We thus see eliciting understanding via multiple measures and exploring how these measures can be combined in individual and group settings as a direction for future research.
+5.3. Which real-world settings would benefit from explainable AI in groups?
+In Section 2.3.2, we described several settings where citizens gather to discuss and form opinions on matters of public interest.
+These included referendums, forums, and community-based spaces.
+This paper investigates settings suitable for deliberating the deployment of public AI systems, an issue that we frame as a matter of public interest due to the scope and severity of its potential consequences.
+Having established that using explanations in group settings benefits participants’ understanding, decision confidence, and decision-making processes, it is worthwhile to consider how this setting could be employed in real-world contexts.
+One answer can be given based on participants’ feedback, who stated that training in their job agency should employ a similar format to educate about AI.
+Notably, this feedback was given by domain experts (Groups C, G) and decision subjects (Group E), suggesting that the setting would suit both stakeholder groups for an educational intervention.
+Similarly, P3 explained that the explanation approach could be helpful if a similar system were used in their care facility by embedding it in the team’s regular meetings, in which difficult cases are discussed and joint decisions are made.
+These insights are in line with previous work on XAI in public institutions.
+Notably, Lee et al.
+(2019b) and Weitz et al.
+(2024) conducted participatory workshops to design explanations with end-users in the public sector, finding that co-designing explainable AI helps in considering the needs of both clients and end-users.
+We envision that collaborative settings and ‘mini-publics’ (Fung, 2003) could be useful in many contexts that aim to strengthen participatory democracy with respect to AI.
+Potential areas of application could be professional consultation workshops for citizens affected by algorithmic decisions, comparable to legal clinics (Ontario, 2025), community-based education and training interventions, such as “contestation cafés” (Collins et al., 2024), or union forums that inform and organize employees’ voices about the use of AI in their institution (Kaur et al., 2022).
+On a different note, Crivellaro et al.
+(2019) found that participatory formats that aim to connect communities to public institutions can suffer from a lack of crucial information (e.g., budgets), which could be alleviated by an information structure such as the presented explanation design.
+In short, we propose that explainable AI in collective settings could be a valuable engagement format for contexts in which public AI could impact people’s lives.
+Future work could explore how collective XAI settings could be implemented in these contexts as part of responsible AI initiatives and in connection to both institutionalized (Costanza-Chock et al., 2022) and user-based (Shen et al., 2021) auditing practices.
+5.4. What’s missing from the explanation design and how could it be improved?
+In Sections 4.1.4, we described that participants appreciated the explanations’ comprehensive and flexible information selection and self-directed and active exploration.
+However, they also noted that the explanations have a high access threshold and require adjustment to the modular structure, making oversight difficult.
+A digital version of the explanation design could improve the overview through summaries and navigation while allowing for simple language options and cross-references.
+As the simple awareness of the scope of information also seemed to overwhelm participants, approaches to condense the scope and selection would be beneficial.
+An example could be a recommendation system that would suggest to participants explanations from different categories and levels of detail based on their stated interests and technical knowledge.
+However, we also emphasize that identifying explanation subtopics and splitting them up into levels of detail is challenging.
 Our explanations’ structure provides different levels of soundness and completeness (Kulesza et al., 2013), but deciding how the available information is allocated into this structure requires a subjective choice.
-In our study, detail levels 2 and 3 were supposed to convey more background rather than essential information gradually.
+In our study, detail levels 2 and 3 were supposed to convey more detailed but also more difficult explanations than the base levels.
 Still, some participants stated that the most critical information for them resided in level 3.
-We see avenues for future work in selecting and hierarchically structuring information to be included in explanations for ADM systems, such that this structure allows for exploration while not obscuring essential or complex information.
-5.2. Benefits and challenges of using explanations in collaborative settings
-In Sections 4.2.1+4.2.2, we described that focus groups’ interactions with information led to ”collaborative success” (Nokes-Malach et al., 2015) when the group dynamic was constructive, i.e., argumentative (Mercier and Sperber, 2011), familiar (Johnson and Johnson, 1985), and solution-oriented (Niculae and Danescu-Niculescu-Mizil, 2016), and that explanations can inform conversations about central topics in public AI deployment (Kawakami et al., 2024).
-In the following, we discuss these findings by outlining the resulting benefits and challenges for collaborative XAI.
-Benefits: Explanations support collaborative reasoning, problem-solving, and deliberation.
-In the best cases, focus groups in our study had an open, communicative, and investigative atmosphere that facilitated the up-taking, sharing, and discussion of information.
-In these settings, the explanations showed their strengths by allowing for the distribution of reading tasks among group members, providing high levels of detail and breadth if needed, and offering different viewpoints that could be used as argumentative and conversational starting points.
-In this sense, the explanations fulfilled their aim of supporting collaboration and informed decision-making around deploying a public AI system (Kawakami et al., 2024) by creating common ground that participants used as a basis for deliberation.
-This deliberation process is the differentiating factor compared to ”one-to-one” (Naiseh et al., 2021) explanation settings, which might allow for more concentrated learning (Bloom, 1984), but miss the exchange of knowledge and perspectives that could become central for societal discourse about public AI (Züger and Asghari, 2023).
-While this collaborative atmosphere has several preconditions, such as containing cognitive biases (groupthink (Janis, 1972), equality bias (Naiseh et al., 2024)) and requiring trust between group members (Johnson and Johnson, 1985), our findings suggest that explanations can be an effective enabler in supporting these crucial group interactions.
-Regarding learning and understanding, XAI could further benefit from connecting to research showing that small groups can exhibit collective rationality (Moshman and Geil, 1998), outperform the wisdom of the crowds (Navajas et al., 2018), and solve tasks correctly even if no group member knows the solution beforehand (Smith et al., 2009).
-Challenges: Explanations must adapt to different groups and require constructive dynamics.
-While our explanations were effective in supporting interaction and opinion formation in most focus groups, we also observed several challenges resulting from group compositions and trust between participants.
-While the explanations typically worked well in groups of domain experts, especially G (executive staff) and H (job trainers and counselors), special interaction situations occurred in Groups B and H due to two participants’ leading roles in the organizations.
-This could have resulted in disrupted interaction due to fear of being evaluated by someone perceived as superordinate, but, in fact, had no negative impact on collaboration.
-In Groups E and F (decision subjects), however, the unfamiliarity of participants and the explanation’s high access threshold led to a lack of interactions and discussion, indicating that the explanation did not fully support these compositions.
-In these groups, a lack of trust between participants likely disrupted interactions when participants were unknown to each other, amplifying effects such as the fear of being evaluated (Nokes-Malach et al., 2015).
-This underscores the importance of creating trust between group members in collaborative XAI settings (Johnson and Johnson, 1985).
-Intuitive measures could be the introduction of a simple task that the group solves collaboratively before engaging with explanations, such as the Wason card selection task (Wason, 1968), or creating an atmosphere that allows participants to interact in roles (e.g., proponents and opposition) to facilitate discussion without revealing personal opinions.
-Future work should examine how such measures can be incorporated into explanation design to support interaction in groups of comparable compositions.
-Lastly, and surprisingly, we did not observe many cognitive biases such as ”groupthink” (Janis, 1972) (aiming for consensus at the cost of negotiation and alternatives) or the ”equality bias” (Naiseh et al., 2024) (downplaying one’s expertise to weigh everyone’s opinion) in our study.
-In one example of groupthink, participants in Group A changed their vote to ”Yes” to reach a group decision, perhaps driven by their (stated) lack of being personally affected by the system’s deployment.
-However, this is contrasted by participants in Group D, who debated at length about the system’s deployment without reaching a consensus, despite also not being affected personally.
-Potential measures to avoid groupthink in discussion could thus be to encourage debate, which again could be the introduction of roles to improve the ”dialectic argumentation” (Mercier and Sperber, 2011), and to explain the system in a way that makes it more personally relevant to participants (Wiggins and McTighe, 2005), e.g., by emphasizing connections to their own experiences.
-5.3. Explanation design suggestions
-We summarize the implications of our findings in the form of suggestions for the design of explanations that are suited to AI novices in individual and group settings, as depicted in Figure 7.
+We see avenues for future work in selecting and hierarchically structuring information to be included in explanations for AI systems, such that they allow for exploration while not obscuring essential information.
+This essential information should balance textual and visual design to ensure its uptake does not rely only on textual understanding.
+Due to the modular structure and the complexity of some of the information, several participants used intuition in answering the study tasks instead of truly searching for the explanations for solutions.
+Framing this in the dual-process theory of cognition, we observed that participants in these moments used System 1 (intuitive heuristics) rather than System 2 thinking (analytical reasoning) (Buçinca et al., 2021; Kahneman, 2011).
+Lambe et al.
+(2016) listed strategies to counteract this tendency and encourage analytical thinking in the medical domain, including checklists, cognitive forcing mechanisms (consideration of alternative diagnoses, reconsideration of diagnoses), guided reflection, and use of particular reasoning approaches.
+Buçinca et al.
+(2021) further tested cognitive forcing mechanisms in an AI-assisted decision-making scenario, in which they used three interventions: showing participants an AI decision only on demand, showing the AI decision only after the participants made their own predictions, and letting them wait before showing them the AI decision.
+These mechanisms improved the performance of the human-AI teams but led users to dislike the interface’s usability, presenting a trade-off.
+Nonetheless, we concur that these strategies should be considered in future explanation design to “ensure that people will exert effort to attend to those explanations” (Buçinca et al., 2021).
+5.5. Summary of explanation design suggestions
+We summarize the implications of our findings in the form of suggestions for the design of explanations suited to AI novices in individual and group settings in Figure 8.
 6. Limitations
-In this section, we briefly describe limitations of our study that the reader should consider.
+Like any research, this study had limitations.
 Due to the limited sample size, we did not analyze the impact of sex and/or gender on our results, limiting the results’ generalizability regarding these aspects but not their overall validity.
 Further, our participants were recruited from organizations and networks in the same geographical region, perhaps resulting in regional or cultural biases.
-The presented use case is embedded in a specific sociotechnical context (Ehsan et al., 2023a) that might affect participants’ understanding and perceptions (e.g., perceptions might differ between employability prediction and credit approval), and thus a change in the domain might also change the explanations’ effect.
+The presented use case is further embedded in a specific sociotechnical context (Ehsan et al., 2023a) that might affect participants’ understanding and perceptions (e.g., perceptions might differ between employability prediction and credit approval), and thus, a change in the domain might also change the explanations’ effect.
 However, this does not limit the transferability of the explanation design, which can be seen as a template that can be adjusted to other use cases.
-We further note that our participant sample has a bias toward university education in the single interviews, which we addressed by comparing these participants mostly with university-educated participants in the focus groups.
+We further note that our participant sample is biased toward university education in the single interviews, which we addressed by comparing these participants mostly with university-educated participants in the focus groups.
+We are further aware that the cooperation with civil society organizations in the recruitment of participants could have led to selection biases, especially in the form of convenience sampling (over-representation of readily available participants), self-selection bias (over-representation of strongly motivated participants), and interviewer bias (over-representation of agreeable or compatible participants) (Collier, 1995).
+We aimed to counteract these biases by defining research goals and methods clearly before recruiting participants, by using multiple recruitment sources and methods, ensuring that group composition was diversified, and by reflecting on possible sampling influences in the analysis of results.
+Due to the qualitative approach, our recruitment strategy further did not aim for statistical generalizability but instead intended to cover a variety of “theoretically relevant cases” (Collier, 1995) and “careful contextualization” (Collier and Mahoney, 1996) to examine our research questions.
 7. Conclusion
-In this paper, we proposed and empirically validated a question-driven, modular explanation design for individual and collaborative settings.
-We conducted an interview study with AI novices in 8 focus groups and 12 single interviews, analyzing the explanations’ effect on understanding and decision confidence, participants’ perceptions of key information, and the interaction processes in both settings.
-We found that explanations were effective at supporting and calibrating participants’ understanding and at informing individual and collective decision-making.
-Our findings further showed that our explanations support focused and detail-oriented interaction in individual settings while also being suited to be shared and discussed in group settings.
-In groups with constructive dynamics, interactions with the explanations led to ”collaborative success” (Nokes-Malach et al., 2015) and informed group discussions about pivotal topics regarding the system’s deployment.
-For groups in which the explanations were less effective, we suggest adjusting the group dynamic, modifying the explanation’s format, and using co-design approaches to better adapt explanations to various needs.
-Based on these findings, we present suggestions for the design of explanations for AI novices and outline how their application can support collaborative reasoning and discussion.
-With this work, we aim to contribute an explanation design approach to the field of collaborative XAI and present implications for how explanations can support stakeholder deliberation about public AI systems.
+This paper tested a question-driven, modular explanation design with AI novices in groups and individual settings.
+We conducted an interview study involving 8 focus groups and 12 single interviews.
+We analyzed them to examine the effect of explanations on understanding, decisions, and decision confidence, participants’ perceptions of key information, and the interaction processes in both settings.
+We found that explanations supported participants’ understanding and decision-making differently, encouraging focused interaction in individual settings and shared understanding in group settings.
+Even though individuals could not exchange with others, the explanations still led to increased decision confidence and changes by supporting internal deliberation.
+In groups, the explanation design afforded a set of interactions that allowed participants to support each other’s understanding, and further provided grounds for exchanging arguments about key aspects regarding the system’s deployment.
+For groups that experienced collaborative failure, we suggest the modification of the explanation’s design to highlight essential information and measures to create a more productive social dynamic.
+With this work, we aim to showcase the potential of combining explanations with group settings to enable AI novices to understand and deliberate about public AI systems.
 Acknowledgements.
@@
 https://doi.org/10.1126/science.aao4408
+- Baron (2005) Robert S.
+Baron.
+2005.
+So Right It’s Wrong: Groupthink and the Ubiquitous Nature of Polarized Group Decision Making.
+In Advances in Experimental Social Psychology.
+Vol.
+37.
+Elsevier, 219–253.
+https://doi.org/10.1016/S0065-2601(05)37004-3
 - Baumberger et al.
@@
 https://doi.org/10.1145/3593013.3594053
+- Biran and McKeown (2017) Or Biran and Kathleen McKeown.
+2017.
+Human-Centric Justification of Machine Learning Predictions.
+In Proceedings of the Twenty-Sixth International Joint Conference on Artificial Intelligence, IJCAI-17.
+1461–1467.
+https://doi.org/10.24963/ijcai.2017/202
 - Bloom (1984) Benjamin Bloom. 1984. The 2 Sigma Problem: The Search for Methods of Group Instruction as Effective as One-to-One Tutoring. (1984).
@@
 - Bundesagentur für Arbeit (2021) Bundesagentur für Arbeit. 2021. Bearbeiten von Bewerberdaten durch Träger. arbeitsagentur.de/datei/dok_ba013193.pdf.
-- Burns et al.
-(2023) Alyxander Burns, Christiana Lee, Ria Chawla, Evan Peck, and Narges Mahyar.
-2023.
-Who Do We Mean When We Talk About Visualization Novices?.
-In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23).
-Association for Computing Machinery, New York, NY, USA, Article 819, 16 pages.
-https://doi.org/10.1145/3544548.3581524
+- Buçinca et al.
+(2021) Zana Buçinca, Maja Barbara Malaya, and Krzysztof Z.
+Gajos.
+2021.
+To Trust or to Think: Cognitive Forcing Functions Can Reduce Overreliance on AI in AI-assisted Decision-making.
+Proceedings of the ACM on Human-Computer Interaction 5, CSCW1 (April 2021), 1–21.
+https://doi.org/10.1145/3449287
 - Byrne (2023) Ruth M.J.
@@
 https://doi.org/10.1145/3544548.3580959
+- Carter et al.
+(2014) Nancy Carter, Denise Bryant-Lukosius, Alba DiCenso, Jennifer Blythe, and Alan J.
+Neville.
+2014.
+The Use of Triangulation in Qualitative Research.
+Oncology Nursing Forum 41, 5 (Sept.
+2014), 545–547.
+https://doi.org/10.1188/14.ONF.545-547
 - Chatti et al.
@@
 https://doi.org/10.1145/3290605.3300789
+- Chiang et al.
+(2023) Chun-Wei Chiang, Zhuoran Lu, Zhuoyan Li, and Ming Yin.
+2023.
+Are Two Heads Better Than One in AI-Assisted Decision Making?
+Comparing the Behavior and Performance of Groups and Individuals in Human-AI Collaborative Recidivism Risk Assessment.
+In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems.
+ACM, Hamburg Germany, 1–18.
+https://doi.org/10.1145/3544548.3581015
+- Chiang et al.
+(2024) Chun-Wei Chiang, Zhuoran Lu, Zhuoyan Li, and Ming Yin.
+2024.
+Enhancing AI-Assisted Group Decision Making through LLM-Powered Devil’s Advocate.
+In Proceedings of the 29th International Conference on Intelligent User Interfaces.
+ACM, Greenville SC USA, 103–119.
+https://doi.org/10.1145/3640543.3645199
 - Chouldechova (2017) Alexandra Chouldechova.
@@
 https://doi.org/10.1089/big.2016.0047
+- Collier (1995) David Collier. 1995. Translating quantitative methods for qualitative researchers: The case of selection bias. American Political Science Review 89, 2 (1995), 461–466.
+- Collier and Mahoney (1996) David Collier and James Mahoney.
+1996.
+Insights and Pitfalls: Selection Bias in Qualitative Research.
+World Politics 49, 1 (Oct.
+1996), 56–91.
+https://doi.org/10.1353/wp.1996.0023
+- Collins et al.
+(2024) Robert Patrick Collins, Johan Redström, and Marco Rozendaal.
+2024.
+The right to contestation: Towards repairing our interactions with algorithmic decision systems.
+(2024).
+https://doi.org/10.57698/V18I1.06 Publisher: International Journal of Design.
 - Conati et al.
@@
 https://doi.org/10.1007/s10606-008-9080-9
+- Costanza-Chock et al.
+(2022) Sasha Costanza-Chock, Inioluwa Deborah Raji, and Joy Buolamwini.
+2022.
+Who Audits the Auditors?
+Recommendations from a field scan of the algorithmic auditing ecosystem.
+In 2022 ACM Conference on Fairness, Accountability, and Transparency.
+ACM, Seoul Republic of Korea, 1571–1583.
+https://doi.org/10.1145/3531146.3533213
+- Crivellaro et al.
+(2019) Clara Crivellaro, Rob Anderson, Daniel Lambton-Howard, Tom Nappey, Patrick Olivier, Vasilis Vlachokyriakos, Alexander Wilson, and Pete Wright.
+2019.
+Infrastructuring Public Service Transformation: Creating Collaborative Spaces between Communities and Institutions through HCI Research.
+ACM Trans.
+Comput.-Hum.
+Interact.
+26, 3, Article 15 (May 2019), 29 pages.
+https://doi.org/10.1145/3310284
 - de Fine Licht and de Fine Licht (2020) Karl de Fine Licht and Jenny de Fine Licht.
@@
 https://doi.org/10.1145/3544549.3573832
-- European Commission (2021) European Commission.
-2021.
-Laying Down Harmonised Rules on Artificial Intelligence.
-digital-strategy.ec.europa.eu/en/library/proposal-regulation-laying-down-harmonised-rules-artificial-intelligence.
+- Eslami et al.
+(2016) Motahhare Eslami, Karrie Karahalios, Christian Sandvig, Kristen Vaccaro, Aimee Rickman, Kevin Hamilton, and Alex Kirlik.
+2016.
+First I ”like” It, Then I Hide It: Folk Theories of Social Feeds.
+In Proceedings of the 2016 CHI Conference on Human Factors in Computing Systems (San Jose, California, USA) (CHI ’16).
+Association for Computing Machinery, New York, NY, USA, 2371–2382.
+https://doi.org/10.1145/2858036.2858494
+- European Commission (2024) European Commission. 2024. Laying Down Harmonised Rules on Artificial Intelligence and Amending Regulations. .
 - Fiesler et al.
@@
 https://doi.org/10.1007/978-3-031-44064-9_3
+- Fung (2003) Archon Fung.
+2003.
+Survey Article: Recipes for Public Spheres: Eight Institutional Design Choices and Their Consequences.
+Journal of Political Philosophy 11, 3 (2003), 338–367.
+https://doi.org/10.1111/1467-9760.00181 arXiv:https://onlinelibrary.wiley.com/doi/pdf/10.1111/1467-9760.00181
+- Gamper et al.
+(2020) Jutta Gamper, Günter Kernbeiß, and Michael Wagner-Pinter.
+2020.
+Das Assistenzsystem AMAS: Zweck, Grundlagen, Anwendung.
+https://www.ams-forschungsnetzwerk.at/downloadpub/2020_Assistenzsystem_AMAS-dokumentation.pdf
 - Grimm (2019) Stephen R. Grimm. 2019. Varieties of Understanding. In Varieties of Understanding. Oxford University Press, 1–14. https://doi.org/10.1093/oso/9780190860974.003.0001
@@
 https://doi.org/10.1080/10447318.2023.2262797
+- Habermas (1991) Jurgen Habermas. 1991. The structural transformation of the public sphere: An inquiry into a category of bourgeois society. MIT press.
+- Hajer and Wagenaar (2003) Maarten A.
+Hajer and H.
+Wagenaar (Eds.).
+2003.
+Deliberative policy analysis: understanding governance in the network society.
+Cambridge University Press, Cambridge, UK ; New York, USA.
+- Henin and Le Métayer (2022) Clément Henin and Daniel Le Métayer.
+2022.
+Beyond explainability: justifiability and contestability of algorithmic decision systems.
+AI & SOCIETY 37, 4 (Dec.
+2022), 1397–1410.
+https://doi.org/10.1007/s00146-021-01251-8
 - Hennink et al.
@@
 https://doi.org/10.3389/fcomp.2023.1117848
+- Holl et al. (2018) Jürgen Holl, Günter Kernbeiß, and Michael Wagner-Pinter. 2018. Das AMS-Arbeitsmarktchancen-Modell.
+- Innes and Booher (2003) Judith E.
+Innes and David E.
+Booher.
+2003.
+Collaborative policymaking: governance through dialogue.
+In Deliberative Policy Analysis (1 ed.), Maarten A.
+Hajer and Hendrik Wagenaar (Eds.).
+Cambridge University Press, 33–59.
+https://doi.org/10.1017/CBO9780511490934.003
+- Janis (1971) IL Janis. 1971. Groupthink/Janis. IL//Psychology Today 5 (1971), 6.
 - Janis (1972) Irving L Janis. 1972. Victims of groupthink: A psychological study of foreign-policy decisions and fiascoes. , viii, 277–viii, 277 pages.
@@
 https://doi.org/10.1007/978-1-4899-3650-9_4
+- Kahneman (2011) Daniel Kahneman. 2011. Thinking, fast and slow. macmillan.
+- Karadzhov et al.
+(2023) Georgi Karadzhov, Tom Stafford, and Andreas Vlachos.
+2023.
+DeliData: A Dataset for Deliberation in Multi-party Problem Solving.
+Proceedings of the ACM on Human-Computer Interaction 7, CSCW2 (Sept.
+2023), 1–25.
+https://doi.org/10.1145/3610056
+- Kaur et al.
+(2022) Harmanpreet Kaur, Eytan Adar, Eric Gilbert, and Cliff Lampe.
+2022.
+Sensible AI: Re-imagining Interpretability and Explainability using Sensemaking Theory.
+In 2022 ACM Conference on Fairness, Accountability, and Transparency.
+ACM, Seoul Republic of Korea, 702–714.
+https://doi.org/10.1145/3531146.3533135
 - Kawakami et al.
@@
 - Keil (2006) Frank C. Keil. 2006. Explanation and Understanding. Annual Review of Psychology 57, 1 (2006), 227–254. https://doi.org/10.1146/annurev.psych.57.102904.190100
+- Kerr and Tindale (2004) Norbert L.
+Kerr and R.
+Scott Tindale.
+2004.
+Group Performance and Decision Making.
+Annual Review of Psychology 55, 1 (Feb.
+2004), 623–655.
+https://doi.org/10.1146/annurev.psych.55.090902.142009
 - Kramer et al.
@@
 https://doi.org/10.1109/VLHCC.2013.6645235
+- Lambe et al.
+(2016) Kathryn Ann Lambe, Gary O’Reilly, Brendan D Kelly, and Sarah Curristan.
+2016.
+Dual-process cognitive interventions to enhance diagnostic reasoning: a systematic review.
+BMJ quality & safety 25, 10 (2016), 808–820.
 - Langer et al.
@@
 - Lee et al.
-(2019) Min Kyung Lee, Daniel Kusbit, Anson Kahng, Ji Tae Kim, Xinran Yuan, Allissa Chan, Daniel See, Ritesh Noothigattu, Siheon Lee, Alexandros Psomas, and Ariel D.
+(2019a) Min Kyung Lee, Anuraag Jain, Hea Jin Cha, Shashank Ojha, and Daniel Kusbit.
+2019a.
+Procedural justice in algorithmic fairness: Leveraging transparency and outcome control for fair algorithmic mediation.
+Proceedings of the ACM on Human-Computer Interaction 3, CSCW (2019), 1–26.
+- Lee et al.
+(2019b) Min Kyung Lee, Daniel Kusbit, Anson Kahng, Ji Tae Kim, Xinran Yuan, Allissa Chan, Daniel See, Ritesh Noothigattu, Siheon Lee, Alexandros Psomas, and Ariel D.
 Procaccia.
-2019.
+2019b.
 WeBuildAI: Participatory Framework for Algorithmic Governance.
@@
 https://doi.org/10.1145/3313831.3376590
+- Liao et al.
+(2021) Q.
+Vera Liao, Milena Pribić, Jaesik Han, Sarah Miller, and Daby Sow.
+2021.
+Question-Driven Design Process for Explainable AI User Experiences.
+arXiv:https://arxiv.org/abs/2104.03483 [cs.HC]
 - Lim and Dey (2009) Brian Y.
@@
 Curran Associates Inc., Red Hook, NY, USA, 4768–4777.
+- Lupia (2023) Arthur Lupia. 2023. By Design: How People Adapt to Cognitive Limitations in Politics. Topics in Cognitive Science (Sept. 2023), tops.12690. https://doi.org/10.1111/tops.12690
+- Mair et al.
+(2019) David Mair, Laura Smillie, Giovanni La Placa, Florian Schwendinger, Milena Raykovska, Zsuzsanna Pasztor, René van Bavel, and European Commission (Eds.).
+2019.
+Understanding our political nature: how to put knowledge and reason at the heart of political decision-making.
+Publications Office, Luxembourg.
+https://doi.org/10.2760/374191
+- Mandviwalla and Olfman (1994) Munir Mandviwalla and Lorne Olfman.
+1994.
+What do groups need?
+A proposed set of generic groupware requirements.
+ACM Trans.
+Comput.-Hum.
+Interact.
+1, 3 (Sept.
+1994), 245–268.
+https://doi.org/10.1145/196699.196715
 - Maxwell (2010) Joseph A. Maxwell. 2010. Using Numbers in Qualitative Research. Qualitative Inquiry 16, 6 (2010), 475–482. https://doi.org/10.1177/1077800410364740
+- Mercier and Landemore (2012) Hugo Mercier and Hélène Landemore.
+2012.
+Reasoning Is for Arguing: Understanding the Successes and Failures of Deliberation.
+Political Psychology 33, 2 (April 2012), 243–258.
+https://doi.org/10.1111/j.1467-9221.2012.00873.x
 - Mercier and Sperber (2011) Hugo Mercier and Dan Sperber.
@@
 - Miller (2019) Tim Miller. 2019. Explanation in artificial intelligence: Insights from the social sciences. Artificial Intelligence 267 (Feb. 2019), 1–38. https://doi.org/10.1016/j.artint.2018.07.007
+- Mohseni et al.
+(2020) Sina Mohseni, Niloofar Zarei, and Eric D.
+Ragan.
+2020.
+A Multidisciplinary Survey and Framework for Design and Evaluation of Explainable AI Systems.
+arXiv:1811.11839 [cs.HC] https://arxiv.org/abs/1811.11839
 - Molnar (2022) Christoph Molnar. 2022. Interpretable Machine Learning (2 ed.). https://christophm.github.io/interpretable-ml-book
@@
 https://doi.org/10.1007/s10648-015-9312-8
+- Ontario (2025) Legal Aid Ontario. 2025. What is a legal clinic? - Legal Aid Ontario — legalaid.on.ca. https://www.legalaid.on.ca/faq/what-is-a-legal-clinic/. [Accessed 03-04-2025].
+- Panigutti et al.
+(2023) Cecilia Panigutti, Ronan Hamon, Isabelle Hupont, David Fernandez Llorca, Delia Fano Yela, Henrik Junklewitz, Salvatore Scalzo, Gabriele Mazzini, Ignacio Sanchez, Josep Soler Garrido, and Emilia Gomez.
+2023.
+The Role of Explainable AI in the Context of the AI Act.
+In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency (Chicago, IL, USA) (FAccT ’23).
+Association for Computing Machinery, New York, NY, USA, 1139–1150.
+https://doi.org/10.1145/3593013.3594069
+- Papenmeier et al.
+(2022) Andrea Papenmeier, Dagmar Kern, Gwenn Englebienne, and Christin Seifert.
+2022.
+It’s Complicated: The Relationship between User Trust, Model Accuracy and Explanations in AI.
+ACM Trans.
+Comput.-Hum.
+Interact.
+29, 4, Article 35 (March 2022), 33 pages.
+https://doi.org/10.1145/3495013
 - Patton (1990) Michael Quinn Patton. 1990. Qualitative evaluation and research methods, 2nd ed. Sage Publications, Inc, Thousand Oaks, CA, US. 532–532 pages.
@@
 - Shen et al.
-(2020) Hong Shen, Haojian Jin, Angel Alexander Cabrera, Adam Perer, Haiyi Zhu, and Jason I.
-Hong.
-2020.
-Designing Alternative Representations of Confusion Matrices to Support Non-Expert Public Understanding of Algorithm Performance.
-Proceedings of the ACM on Human-Computer Interaction 4, CSCW2 (Oct.
-2020), 1–22.
-https://doi.org/10.1145/3415224
+(2021) Hong Shen, Alicia DeVos, Motahhare Eslami, and Kenneth Holstein.
+2021.
+Everyday Algorithm Auditing: Understanding the Power of Everyday Users in Surfacing Harmful Algorithmic Behaviors.
+Proceedings of the ACM on Human-Computer Interaction 5, CSCW2 (Oct.
+2021), 1–29.
+https://doi.org/10.1145/3479577
 - Shin (2023) Don Donghee Shin.
@@
 https://doi.org/10.1126/science.1165919
+- Speith (2022) Timo Speith.
+2022.
+A Review of Taxonomies of Explainable Artificial Intelligence (XAI) Methods.
+In 2022 ACM Conference on Fairness, Accountability, and Transparency.
+ACM, Seoul Republic of Korea, 2239–2250.
+https://doi.org/10.1145/3531146.3534639
 - Speith et al.
@@
 https://doi.org/10.1007/s10676-024-09769-3
+- Stromer-Galley (2007) Jennifer Stromer-Galley. 2007. Measuring Deliberation’s Content: A Coding Scheme. Journal of Deliberative Democracy 3, 1 (July 2007). https://doi.org/10.16997/jdd.50
 - Sutcliffe (2005) Alistair Sutcliffe.
@@
 https://doi.org/10.1145/1083106.1083119
+- Swiss Confederation (2025) Swiss Confederation. 2025. The referendum. ch.ch/en/votes-and-elections/referendum.
+- Szigetvari (2018) András Szigetvari.
+2018.
+AMS bewertet Arbeitslose künftig per Algorithmus.
+Der Standard (2018).
+https://www.derstandard.at/story/2000089095393/ams-bewertet-arbeitslose-kuenftig-per-algorithmus
 - Szymanski et al.
@@
 https://doi.org/10.1007/s12525-020-00441-4
+- Veenman et al.
+(2006) Marcel V.
+J.
+Veenman, Bernadette H.
+A.
+M.
+Van Hout-Wolters, and Peter Afflerbach.
+2006.
+Metacognition and learning: conceptual and methodological considerations.
+Metacognition and Learning 1, 1 (April 2006), 3–14.
+https://doi.org/10.1007/s11409-006-6893-0
 - Wang and Yin (2021) Xinru Wang and Ming Yin.
@@
 - Xu (2019) Wei Xu. 2019. Toward human-centered AI: A perspective from human-computer interaction. Interactions 26, 4 (jun 2019), 42–46. https://doi.org/10.1145/3328485
-- Yurrita et al.
-(2022) Mireia Yurrita, Dave Murray-Rust, Agathe Balayn, and Alessandro Bozzon.
-2022.
-Towards a multi-stakeholder value-based assessment framework for algorithmic systems.
-In Proceedings of the 2022 ACM Conference on Fairness, Accountability, and Transparency (Seoul, Republic of Korea) (FAccT ’22).
-Association for Computing Machinery, New York, NY, USA, 535–563.
-https://doi.org/10.1145/3531146.3533118
 - Zagzebski (2019) Linda Zagzebski. 2019. Toward a Theory of Understanding. In Varieties of Understanding. Oxford University Press, 123–136. https://doi.org/10.1093/oso/9780190860974.003.0007
-- Zhang et al.
-(2023) Angie Zhang, Olympia Walker, Kaci Nguyen, Jiajun Dai, Anqing Chen, and Min Kyung Lee.
-2023.
-Deliberating with AI: Improving Decision-Making for the Future through Participatory AI Design and Stakeholder Deliberation.
-Proceedings of the ACM on Human-Computer Interaction 7, CSCW1 (April 2023), 1–32.
-https://doi.org/10.1145/3579601
 - Züger and Asghari (2023) Theresa Züger and Hadi Asghari.
@@
 https://doi.org/10.1007/s00146-022-01480-5
-Appendix
+Supplementary material
+Better Together? The Role of Explanations in Supporting Novices in Individual
+and Collective Deliberations about AI
 The supplementary material provides additional information on the study and use case.
@@
 Section B lists the self-report and interview questions that were asked during the first and second self-reports.
+Section C depicts all explanations that participants could receive in the study.
+Lastly, Section D provides the complete code book, listing themes and codes generated from the qualitative analysis.
 Appendix A Additional information on the employment prediction algorithm
-In this section, we give additional information on the features used by the employment prediction algorithm, the mock newspaper article introducing the system to participants, and the study task participants solved.999More detailed information on the deployment context of the system has been omitted to adhere to the anonymization policy but will be re-inserted for the final version.
+In this section, we give additional information on the features used by the employment prediction algorithm, the mock newspaper article introducing the system to participants, and the study task participants solved.
 A.1. Features used by the employment prediction algorithm
@@
 How do you feel about the decision the group made?
+Appendix C Collection of explanations
+This section depicts all explanations of the four categories data (Figure 12), system details (Figure 13), usage (Figure 14), and context (Figure 15).
+All explanation categories are split into three levels of detail; the background is colored differently for each level to facilitate distinction.
+Explanations of the base level were provided automatically; all others could be requested during the explanations phase.
+A detailed description of the explanations is given in Section 3.
+We insert the explanation overview again as Figure 11 for orientation.
+Appendix D Code book
+LABEL:fig:code_book lists the themes and codes developed from the interview data in the qualitative analysis.
+The table is split into three main sections: Deliberation, understanding, and experiences and opinions.
+The left column lists theme groups and the right columns list single themes as coded in the data.
+Deliberation | ||
+Deliberation - groups | Appreciation of group setting | Forming opinions on deployment |
+Deployment decision changed | Little discussion | |
+Difficult to make deployment decision | No influence from group discussion | |
+Discussing diverging views | Strategic decision-making | |
+Discussion triggers | Unanimous deployment decision | |
+Does not know what to say | Voting against own interests | |
+Following decisions of others | Weighing pros and cons | |
+Arguments - groups | Adverse cognitive effects | System can be misused |
+Conditions for deployment | What’s the benefit? | |
+Gaming the system | What’s the intention? | |
+Influence of Human Factors | Who can I trust? | |
+Integrating system into working processes | Who’s in control? | |
+Scrutinizing the system | Who’s responsible? | |
+Deliberation - individuals | AI harms trust in institution | Deployment despite criticism |
+AI will harm society | Difficult to make deployment decision | |
+Attitude remained unchanged | Weighing pros and cons | |
+Decision more uncertain than before | Would have liked a group setting | |
+Arguments - individuals | Conditions for deployment | Scrutinizing the system |
+Decision-makers are not the right people | Supporting decision subjects | |
+Is it contestable? | System inherits institutional dysfunctions | |
+Influence of human factors | Unions should take a role | |
+Integration is the deciding factor | What are the consequences? | |
+Judgment deviates from algorithm | What’s the benefit? | |
+Lack of transparency | Who’s in control? | |
+Necessary to inform individuals | Who’s responsible? | |
+Projecting hopes on the AI | ||
+Understanding | ||
+Explanations - groups | Cumbersome information uptake | Incomplete coverage of information |
+Differing information needs | Interest beyond time limit | |
+Exchanging explanation sheets | Locating information | |
+Explanation design flaws | Order of processing information | |
+Explanation structure does not make sense | Overwhelmed by information | |
+Explanation structure works | Piecing together information | |
+Explanations help understanding | Relying on intuition over information | |
+Explanations matched with participants | Requesting explanations in bulk | |
+Gap between explanation and application | Suggestions for explanation design | |
+Understanding - groups | Abandoning understanding | Outsourcing understanding |
+Debating task solutions | Participants work individually | |
+Discussing interpretations | Querying and explaining | |
+Dividing understanding burden | Reaching working understanding | |
+Impeded understanding | Sharing information with group | |
+Is system already in use? | Suggestions for explanations in groups | |
+Explanations - individuals | Cumbersome information uptake | Explanations require previous knowledge |
+Difficult to locate information | Focus on context | |
+Every category is important | Focus on system details | |
+Explanation design flaws | Focus on usage | |
+Explanation design suggestions | Going into detail | |
+Explanations adjusted mental model | Interest beyond time limit | |
+Explanations help understanding | Order of processing information | |
+Explanations influenced decision | Overwhelmed by information | |
+Explanations matched participant | Relying on intuition over information | |
+Explanations need to relate personally | Requesting explanations in bulk | |
+Explanations not suited to decision subjects | Skips category | |
+Understanding - individuals | Calculates employment chance | Reaching a working understanding |
+Faults in algorithmic design | Understanding vs information gain | |
+Impeded understanding | Understanding requires example | |
+Experiences and opinions | ||
+Experiences - all | AI is part of digitization | No idea of AI |
+Comparing lived experiences | Overburdened Human-in-the-Loop | |
+Decision subjects have no voice | Using AI at work | |
+Deficiencies in institution | Using AI without knowing it | |
+Digital Humanism as institutional practice | Workplace wants to integrate AI | |
+Is not affected by system | ||
+Opinions - all | AI aversion | Deploying institution has bad reputation |
+AI can assist in decisions | Disagrees with policy choices | |
+AI cannot replace humans | Discrimination with and without AI | |
+AI decisions must be revisable | Fears algorithmic imprint | |
+AI increases objectivity | Formalization is inevitable | |
+AI is inevitable | Good intentions, badly executed | |
+AI just appeared | Need for AI-Human collaboration | |
+AI misrepresents reality | No opinion | |
+AI openness | Peaked interest | |
+AI replaces humans | Public narratives | |
+AI will not improve work |